### v1

# Gaia Data Release 3: spectroscopic binary-star orbital solutions

Abstract
Context. The Gaia satellite constitutes one of the cornerstone missions of ESA. This is an astrometric space experiment that is measuring positions, proper motions as well as parallaxes for a huge number of stars. It is also performing photometric and spectrophotometric observations. Gaia also operates a medium-dispersion spectrometer, the RVS, that provides spectra and thus radial velocity (RV) time-series.
Aims. The paper is centred on the analysis of the RV time-series. We try to fit orbital models and trend models and restrict ourselves to the objects of spectral types F-G-K brighter than magnitude 12 presenting only one single spectrum (SB1).
Methods. Suitable time-series are processed and analysed object per object, providing orbital or trend solutions. The results of the various fits are further filtered internally on the basis of several quality measures in order to discard spurious solutions. The objects only having a spectroscopic solution are classified in one of the three classes SB1 (eccentric model), SB1C (circular model) or TrendSB1 (mere trend model).
Results. We detail the methods used and describe the derived parameters and results. After a description of the models considered, and of the related quality tests of the fit, we detail the internal filtering that is intended to reject bad solutions. We also present a full validation of the pipeline. A description of the current content of the catalogue is provided.
Conclusions. We present the SB1, SB1C and the TrendSB1 spectroscopic solutions contained in the SB-subcatalogue which is part of the DR3 catalogue. We delivered some 181327 orbital solutions in the class SB1, 202 in the class SB1C and 56808 in the associated class TrendSB1. This is a first release and the delivered SB-subcatalogue could be tuned and refined. However the majority of the entries are correct; this data set constitutes by far the largest set of spectroscopic orbital solutions ever constituted.
Key Words.:
Techniques: spectroscopy; Stars: Techniques: radial velocities; Stars: binaries; Catalogues: Survey; Gaia1 Introduction
The Gaia satellite constitutes one of the cornerstone missions of the European Space Agency. This is an astrometric space experiment intended to represent a new breakthrough compared to the previous pioneering one named Hipparcos (Perryman 2009).
The Gaia satellite has two telescopes, looking at 1065 of each other, and three associated scientific instruments, continuously scanning the celestial sphere thanks to its six-hour spinning on itself (around an axis perpendicular to the two telescope lines of sight), to the precession of the spinning axis around the direction of a fictitious, nominal Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2) with a 63-day period and to the one-year orbit of the satellite around the Sun accompanying the Earth in its revolution (Gaia Collaboration et al. 2016b). Multiple visits are necessary to reach a convenient S/N ratio, but also to gather information on binarity (or higher order multiplicity) and variability.
The Gaia satellite is measuring the positions and transverse proper motions for some 1.8 billion stars, as well as precise parallaxes and thus distances for a very large number of objects. The detectors are working in a Time Delay Integration mode (TDI). The satellite was launched on December 19th, 2013 and is scanning the sky since the end of the commissioning period in July 2014. The detailed description of the project and of the payload can be found in Gaia Collaboration et al. (2016b) that was published along with the first Data Release (DR1). The description of DR1 and of the related data products is provided in Gaia Collaboration et al. (2016a) whereas the DR2 was detailed in Gaia Collaboration et al. (2018a). The present paper is to be associated with the DR3 (Gaia Collaboration et al. 2023d). Besides astrometry performed with the Astrometric Field (AF) instrument (Lindegren et al. 2018, 2021), the satellite also produces photometry from the same instrument as well as spectrophotometry provided by another channel: the BP/RP spectrophotometer purveyor of very low resolution spectra as well as of photometric colours (Riello et al. 2018, 2021; Evans et al. 2018). The two above-mentioned channels are supposed to observe stellar objects down to about magnitude 20th in band.
Besides astrometry and spectrophotometry, the satellite also incorporates a third channel: a medium resolution spectrometer to measure the radial velocities (RVs) of stellar objects, named the Radial Velocity Spectrometer (RVS). This channel and the related instrument are described in great details in Cropper et al. (2018) which also includes a nice review of the genesis of the RVS instrument. The RVS brings the third component of the space velocity (to join with the transverse proper motion on the sky). It also provides for the brightest stars information on the nature of the object (, , chemical abundances, chromospheric activity, and so on). Stars brighter than magnitude = 2.76 produce saturated pixels and are not considered (for the definition of the magnitudes, see Sartoretti et al. 2023).
The scanning law for RVS is intended to typically produce 40 transits over the nominal five-year mission. The focal plane of the RVS is covered by 12 CCDs (4 rows of 3 in a row). The RVS instrument acquires, in TDI mode, spectra in the range 845-872 nm, including the well-known Ca ii triplet (in late-type stars). With a resolving power of about (see section 8.2 and 11 of Cropper et al. 2018), it obtains three spectra per transit, each corresponding to a maximum exposure time of 4.42 s (they are analysed individually and concomitantly). These spectra are spread over about 1100 pixels (Cropper et al. 2018), but the part used for RV determination is restricted to the range 846-870 nm (Sartoretti et al. 2023). Very bright objects benefit from 2D spectra whose across scan component is 10 pixels.
DR2 was the first release concerning the spectrometer channel and some 280 million spectra were treated and resulted in the publication of median (over transits) RVs for over 7 million objects (brighter than mag. 12, being single-line spectra and corresponding to assumed constant RVs) in the range of effective temperatures K. This catalogue of single-line, non-exotic stars has already been widely used to study the structure of the Galaxy (Gaia Collaboration et al. 2018b; Antoja et al. 2018; Kawata et al. 2018; Ramos et al. 2018; Monari et al. 2019; Koppelman et al. 2019; Marchetti et al. 2019; Trick et al. 2019; López-Corredoira & Sylos Labini 2019; López-Corredoira et al. 2020; Khoperskov et al. 2020; Nitschai et al. 2020).
The data coming out of the RVS are treated by ground-based spectroscopic processing. The spectroscopic processing is ensured by the Coordination Unit 6 (CU6) of the Data Processing and Analysis Consortium (DPAC; Gaia Collaboration et al. 2016b). The first description of this processing was given in Sartoretti et al. (2018) for DR2 and the properties and validation of the RV set were described in Katz et al. (2019). The stars exhibiting a double-line spectrum (also named composite spectrum) were rejected from the process (as described in Sartoretti et al. 2018; Damerdji et al. 2024). The spectroscopic processing pipeline is made of a series of modules, each taking into account a series of tasks. The first modules are working per transit and per trending epochs. The flowchart is illustrated in section 2.3 of Sartoretti et al. (2018). The Single Transit Analysis (STA) is the workflow that actually performs the measurements of the RV per transit. Three submodules are measuring the RVs of all the selected convenient stars (single spectrum ones). They are based on the idea to derive the RV by cross-correlating the observed spectrum with a synthetic spectrum built in generate_template (Sartoretti et al. 2018). The interest to make use of several methods is explained in the early investigation of the problem, whose conclusions are detailed in David et al. (2014) (see also section 7 of Sartoretti et al. 2018). The STA measures the RVs per object and per transit. Thereafter, the Multiple Transit Analysis (MTA) workflow performs some statistical analysis tests based on the RVs from all objects at all transits. It computes per star a probability of constancy that is used to spot out variable objects. It also computes the median of the RVs over all transits for all the bright objects (Sartoretti et al. 2018). The constancy probability is not used to define the list of assumed constant objects that are considered to compute their median RV. The updated version of the spectroscopic chain for DR3 is described in Sartoretti et al. (2022); Katz et al. (2023). In DR3, MTA further computes mean spectra for assumed constant objects (Seabroke et al., in preparation).
In the context of DR3, the RV values corresponding to variable objects are transmitted to the CU4 (object processing) and in particular to the Non-Single Star (NSS) part of the object processing that analyses these data in order to derive possible orbital solutions. Three channels running in parallel deal with the astrometric data (Halbwachs et al. 2023; Holl et al. 2023b), the photometric data (Siopis et al. 2024; Pourbaix et al. 2022, see section 7.6 by Siopis & Sadowski) and the spectroscopic data (the present paper, and Damerdji et al. 2024). The detailed description of the CU4 spectroscopic processing chain for SB1 (spectroscopic binaries with only one spectrum visible; hereafter labelled NSS-SB1) orbital solutions or trend ones is the subject of the present paper whereas Damerdji et al. (2024) describes the SB2 channel. These DR3 products are preliminary results; this is the first time that orbital solutions issued from Gaia are delivered. The results of the three channels can be further processed by combining their independent outputs. This is the task of the combiner (Gaia Collaboration et al. 2023a; Pourbaix et al. 2022, see section 7.7 by Gavras & Arenou). The scientific conception of the SB1 chain described here was elaborated by Yassine Damerdji, Eric Gosset and Thierry Morel; most of the codes were benchmarked in MATLAB and finally written in Java by YD with some help from TM; the validation tests (including the critical building of the ground-based data catalogues used for comparison) were mainly performed by TM. The whole work was made in the highly favourable and boosting context of the NSS team.
Section 2 describes the nature of the data that are entering the process, whereas Section 3 describes the simple mathematical formalism used and lists the models that are considered. Section 4 explicits the importance of the period search and the related difficulties. Sect. 5 briefly summarises the main trials we did to finally arrive at the method described in Sect. 6 that presents the chain as adopted for the DR3 (NSS-SB1 main processing) whereas Sect. 7 introduces the notion of post-processing. Section 8 illustrates the results. Section 9 concerns the validation of the results of the chain. Additional considerations in Sect. 10 explain the details of the post-filtering and provide additional information in the light of the conclusions from the previous Section. The catalogue description, the caveats and the precautions necessary in using the results of the NSS-SB1 processing are given in Section 11. Section 12 concludes the paper. The content of the present paper is more detailed and contains material that supplants the information given in the Gaia archives (see sections 7.4 and 7.5 in Pourbaix et al. 2022)111https://gea.esac.esa.int/archive/documentation/GDR3/ at the epoch of the release.
2 Description of the input data
The DR3 data correspond to the first 34 months of the nominal mission (25 July 2014 - 28 May 2017, almost 1038 d, 2.84 y). The instrumental spectra obtained at each RVS transit are converted to proper physical spectra (reduced) by the pipeline developed by the CU6 (spectroscopic processing), see Sartoretti et al. (2022); Katz et al. (2023). Per transit, one spectrum is acquired by each of the three CCDs in a row of the RVS (Cropper et al. 2018). The RVs are measured individually on the three spectra. Three methods are used to derive the radial velocities (Cropper et al. 2018), each of them providing three velocities (one per CCD) and a combined RV value (over the three CCDs). The three combined RVs issued from the three methods are transmitted to the Integrator that provides a single combined value (per transit). Only this single combined value is considered here but the other intermediate RVs are used as a sanity check to validate the final combined value (see Sects. 6.4.8.2 and 6.4.8.5 of Sartoretti et al. 2022). The RVs are measured by cross-correlating the observed reduced and calibrated spectra with a theoretical/synthetic spectrum computed on the basis of various stellar atmosphere models. These so-called templates are identified through a set of three atmospheric parameters (effective temperature, surface gravity, metallicity). The selection of the template used to measure the RVs is based on these three parameters only; the alpha element content is not considered as an independent parameter. The assignation of the set of parameters to the object treated is performed on the basis of three databases. They are ordered here in the order of preference. The first possible origin is an internal compilation of ground-based catalogues. The second one is from an early run of the Gaia General Stellar Parametrizer from Spectroscopy (GSP-Spec, Recio-Blanco et al. 2023). The third one is from an early run of the Gaia General Stellar Parametrizer from Photometry (GSP-Phot, Andrae et al. 2023). Finally, if none of these sources provides the necessary parameters, the latter are estimated as previously done for DR2 (DetermineAP: section 6.5 of Sartoretti et al. 2018). This process DetermineAP cross-correlates the observed spectrum with each template from a limited subset of 28 selected ones, and this for every transit. The template giving per transit the highest correlation is retained, whatever is the velocity. The final adopted template is the one that is the major contributor over all the transits. More details can be found in section 3.6.1 of Katz et al. (2023). The repartition among the various origins for the data set treated here is the following: Ground-based catalogues (14.4 %), GSP-Spec (33.4 %), GSP-Phot (45.6 %) and DetermineAP (6.6 %). The RV measurements from transit to transit are performed with the same template. Consequently, a template mismatch is expected to essentially introduce a systematic RV shift (for single-line objects).
The measurements of the RVs are performed under the responsibility of the STA Development Unit. The list of good RVs corresponding to a particular star are then analysed by the MTA in order to detect and separate variable RV sets from constant RV sets. For each star, the median of the RVs is computed and appears in the main catalogue. The stars with variable RVs are further considered by the global pipeline and are forwarded to the CU4 (Object processing: NSS). For DR3, the adopted cutting threshold corresponds to a probability of variability of 0.99 (rv_chisq_pvalue . The composite spectra and SB2 objects are separately treated already at the level of the CU6 processing (Damerdji et al. 2024) for the determination of both RVs and are then dispatched to CU4/NSS in order to search for the corresponding orbital solutions (Damerdji et al. 2024).
The Non-Single Star (NSS) processing is intended to perform the reduction of the corresponding data from the point of view of multiple star studies. In the present paper, we will consider the sole spectroscopic channel for the SB1 objects.
The input data are made of a list of RVs (and their 1-uncertainties) that are corrected to make reference to the barycentre of the solar system. It should be pointed out that for DR3 the rms error on the barycentric correction ( km s-1) is not propagated to the uncertainty on the individual RVs. The RVs are accompanied by a time of observations expressed in barycentric Julian days (BJD). The number of points of the time-series corresponds to the number of transits for which the RV has been properly measured with success. The data could present several weaknesses such that outliers and wrong RVs due e.g. to template mismatch errors (for the latter case, most probably the same systematic for all values within a time-series). These problems and additional selection rules are further documented in Katz et al. (2023) and will not be further detailed here. It should however be noticed that some objects have been rejected at the level of the input data filtering in CU6 because they correspond to a poorly performing Astrometric Global Iterative Solution (AGIS) solution or to a large error on the along-scan field angle (see its definition in Sartoretti et al. 2018). The threshold has been chosen at mas, equivalent to about 29 km s-1. Other filtering applied to the input CU6 data is described in section 4 of Katz et al. (2023). In the course of the CU6 STA/MTA validation, a bias on output RVs was noticed for stars fainter than (being 0 km s-1 at magnitude 11 and reaching 75 m s-1 at magnitude 12, Katz et al. 2023). The correction has not been applied for the catalogue, nor for the transmission of the RVs to the NSS processing. The effect is quite small and should only induce per object a global shift of all the RVs with as a consequence a small bias on the systemic velocities (see Sect. 3).
The current astrometric solution for DR3 has been released as part of the EDR3 and contains, among others, positions for a large number of objects. The astrometric solution (AGIS) is based on a treatment assuming the objects are single, unresolved stars and thus having an image characterised by the instrumental line-spread function (for the one dimensional observations in the astrometric channel) or point-spread function (for the two-dimensional observations), see the description in Lindegren et al. (2021). Some objects do not behave as expected according to this assumption. The pathological behaviour could be due to anomalies, but a large contribution to this problem originates in unresolved or poorly resolved double (or multiple) objects. The separation threshold is considered to be around 04-05 (see Fabricius et al. 2021; Holl et al. 2023a). The double objects not following the behaviour for single stars are associated with a bad adjustment to the global astrometric solution and are thus accompanied by a large renormalised unit weighted error (ruwe). Lindegren et al. (2021) also associate to these objects an excess noise (impacting the position) that is delivered in the catalogue under the name astrometric_excess_noise and which corresponds to the additional dispersion necessary to accept the concerned object in the global solution. Objects with a ruwe larger than 1.4 are certainly problematic, but smaller values already have an impact on the RVS calibration. Since the RVS is a slitless spectrometer, the calibration in wavelength of the spectra is requesting an epoch position of the object that is accurate enough. Therefore, for the objects that are deviating from the single star behaviour, we could expect that the wavelength calibration is not correct, and thus the measured transit RVs are somewhat biased and this creates artefacts in the time-series. No correction of this effect is offered for DR3. As a first approach, we preferred not to discard these objects on the basis of the ruwe. In order to take into account this problem, it was decided to inflate the uncertainty on the RV value by introducing (quadratically) an additional component linked to the astrometry. This term is derived as being (km s-1) = 0.146 astrometric_excess_noise (mas) where the factor 0.146 is related to the dispersion of the spectrometer (transformation of mas into km s-1). Typically, the values of astrometric_excess_noise are below 10 mas, and thus is below 1.46 km s-1.
Some selections are also made on the objects at the moment of their entry in the NSS-SB1 processing. In the case of DR3, the kind of stars is restricted to the range of effective temperatures 3875 K to 8125 K (mainly ((M))-K-G-F-(A) stars). No emission-line object has been selected, and in principle only stars without spectral peculiarities are entering the game. The objects with a rv_renormalised_gof larger than 4 were selected (Katz et al. 2023) to enter the NSS-SB1 processing because they were considered as clearly non-constant. The measurements of the individual (per transit) RVs are restricted to the range –1000 to +1000 km s-1. A global view of the selections applied beforehand is available in Table 18 (Appendix H). This measurement of the individual RVs on individual spectra is only possible for bright objects. The distribution in magnitude of the treated objects is provided in Fig. 15 in the validation Sect. 9.2.1. A threshold on has been selected at 12 mag. Fainter objects are considered to produce individual spectra for which the S/N ratio forbids to derive a sound radial velocity. This strongly reduces the amount of selected objects. The stellar RVs are not processed by the NSS spectroscopic channel if the number of transits (data points) is less than 10 (see details in Sect. 6.3.1).
The nature of the measurement itself (cross-correlation with templates) implies that it is not possible to discriminate between a change in RV due to a global shift of the line (as expected for SB1 objects) and a shift due to a line-profile variation (as expected e.g. for some intrinsic variables). Although possibilities to get rid of this problem exist, they are not implemented for DR3 and this certainly constitutes a limitation of the present orbital-solution pipeline. In particular, fake SB1 could persist in the sample corresponding to never-deblending SB2.
When the SB2 chain (Damerdji et al. 2024) analyses an SB2 time-series and is not able to derive the corresponding SB2 orbital solution, the object can be redirected to the SB1 chain. An SB1 time-series is built with the RVs characterised by the smallest uncertainties. The object is then considered as an SB1 and is forwarded to the present chain. This is a second possibility of entering the SB1 channel. These objects are flagged (for more details, see Damerdji et al. 2024, and also Appendix I).
3 The underlying orbital models
Four kinds of orbital solutions are considered here: they are labelled SB1, SB1C, TrendSB1 and StochasticSB1. They are described here below. The single-line RVs treated here and consequently the deduced motion is intended to refer to the sole object that is visible in the spectra. The model will be fitted through a classical least-square procedure and the selected objective function is the -function except when otherwise stated. The uncertainties on the parameters are calculated in the classical way on the basis of the diagonal elements of the inverse of the curvature matrix (i.e. the variance-covariance matrix, see Bevington & Robinson 2003). Despite the fact that we use the Greek symbol for all (i.e. concerning input data as well as output parameters) uncertainties, it should be pointed out that all of them are of course estimators. They are always expressed as 1 values.
3.1 Orbital models of type SB1
The general eccentric Keplerian model is, for the spectroscopic channel, expressed in terms of the Campbell coefficients (see the appendices in Halbwachs et al. 2023; Binnendijk 1960, as the seminal reference). The model RV is given by
where is the true anomaly, which is deduced from the eccentric anomaly by
which in turn is expressed as a function of the mean anomaly
where is the time of passage at periastron. The is the argument of periastron measured in the sense of the orbit with its origin at the ascending node. In this model, the eccentricity enters as a strongly non-linear parameter. This is also the case for . The parameters to be determined are , , , , , (respectively, the period, the centre-of-mass velocity, the semi-amplitude, the eccentricity, the argument of periastron and the time of passage at periastron). The first step of the computation is devoted to the determination of the period. Some significance tests are made on the very existence of the period. If the period is not significant, the model is fitting noise and is considered as being invalid. This is a statistical decision. Owing to the insufficient span of time covered for the DR3, no apsidal motion is considered; is considered constant in time.
3.2 Orbital models of type SB1C
Another particular kind of orbit is the circular one, i.e. . In this case, the model is much more simple and is expressed by the following equation
where here is the time of maximum velocity and the is fixed at zero corresponding to the maximum velocity. The free parameters to determine are restricted to , , , . The model is linear except, of course, for the determination of the period. Concerning the , the model can be linearised by using the following equivalent equation
We then compute
and
where this refers to a cyclic variable and is assigned the value closest to . The significance level of the period is here also determined.
3.3 Models of type TrendSB1
It can happen that the RV evolution as a function of time just exhibits a trend. This could be a transient behaviour (e.g. outburst) just observed by chance or correspond to a small piece of slowly accelerated motion (e.g. of a stellar atmosphere). It could also correspond to an orbital motion with a period well longer than the span of observing time covered by the satellite. This span of time is 34 months for the DR3. However, it could be slightly shorter for some objects because it combines with the gaps in the scanning law. The present modelisation is restricted to polynomials of degrees one to four and presenting only one extremum on the interval. The degree of the polynomial is chosen on the basis of an F-test for nested models. In practice, it turned out that only the one- and two-degree polynomials were necessary. No higher degree fit was required. The formalism used for the fit is the same as for Hipparcos (ESA 1997). The TrendSB1 solution is compared to the orbital solutions using a Fisher Snedecor F-test on the comparison of the respective s. It should be made clear that this comparison could be an approximation.
3.4 Models of type StochasticSB1
When the fit of a TrendSB1 solution is not possible nor the fit of an SB1 or SB1C orbital solution corresponding to a significant period, the solution is classified as StochasticSB1. The median RV is considered as constant, but with an extra dispersion compared to the adopted RV uncertainties. This extra dispersion is estimated. Properly speaking, this simple model is not an orbital one. Some badly sampled orbits and high-order multiple systems could fall into this category, perhaps momentarily. Some of them will certainly benefit from an increase of the number of transits in future releases. Short-period intrinsic variables (e.g. Scuti stars) may also receive such solutions.
3.5 The downstream combination of solutions
The outputs of the NSS spectroscopic pipeline are delivered to the combiner that operates downstream. The task of the combiner is to inspect the output of the astrometric, photometric and spectroscopic NSS channels in order to detect possible solutions corresponding to the same object and that could be combined. This will be described in Sect. 7.2.
4 Determination of the periodicity
The first step of the analysis is the search for a periodic behaviour of the RVs. The analysis is based on a periodogram (actually a frequencygram), which has the nature of the Fourier Power Spectrum. Since the time sampling of the Gaia satellite is particularly complex, the classical Fourier formalism is not applicable. To well understand the limitation of the method, it is necessary to have some knowledge of the impact of the sampling and of a few characteristic timescales.
4.1 Effect of the Gaia sampling
In order to explain the effect of the Gaia sampling, we have to start from a single, idealistic, case and progressively add some complexity. We start by considering the spinning of the satellite on itself that makes that a star is measured every 0.25 d (this is our first basic idealistic case since this simple view is never occurring in practice). The existence of the sample periodicity concerns both fields of view individually. Therefore, 0.25 d is the typical, smallest regular and periodic timescale in the sampling. The 0.25 d step should generate in the time-series some aliasing with a typical 2 d-1 Nyquist frequency. This means that a progenitor frequency will be duplicated at , , , and so on (in d-1 units). The duplication is also extended to the negative frequencies. This duplication represents a pure aliasing, since the sampling is in this hypothetical case regular. Therefore, the aliasing is pure and the Nyquist frequency exists. In such a case, intermediate and large periods could have aliases towards periods of 0.25 to 0.5 d (and vice versa).
However, the existence of two fields of view separated by 1065 on the sky renders the situation more complex by adding a regularity that is not properly speaking periodic. Thus, the aliasing is no more pure, we will call it pseudo-aliasing. The semi-amplitude spectral window is a tool that facilitates the analysis of such phenomenon. It has been introduced by Deeming (1975, 1976). In the case of an equidistant sampling, the spectral window is the well-known Dirac comb (or Shah function, see Bloomfield 1976). The general spectral window represents an extension of this Dirac comb to the general case of odd sampling. The property that the generalised Fourier Spectrum of the sampled data is the result of the convolution of the underlying Fourier Spectrum of the signal by the spectral window remains true for almost any kind of sampling (Deeming 1975, 1976). If the satellite was limited to a rotation on itself, the semi-amplitude spectral window corresponding to the combined two fields of view would have the aspect given in the upper panel of Fig. 1. Contrary to the classical case of pure aliasing, the aliasing is here not pure since all the peaks have different heights. Besides the main peak at , the next first peak is at d-1, but it is mildly high. The peak having a height comparable to the peak is the one at d-1 which could represent some almost pure aliasing. This figure represents an idealistic case. First, it is assumed that the data points in each field of view are perfectly regularly distributed. This is not the case because several phenomena induce a small perturbation of the time of the observations. The variations of the Basic Angle as well as micro-clanks and meteoritic-hits are perturbing the regularity (Gaia Collaboration et al. 2016b). We do not further consider these effects in our explanation. The satellite is not only spinning on itself, but also follows a more complicated motion. First, the rotation axis is precessing along an axis directed towards a fictitious nominal position of the Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2). In addition, the satellite located around the Lagrangian point L2 is orbiting around the Sun with the Earth annual motion (change in ecliptic longitude of the Sun, Gaia Collaboration et al. 2016b). The major consequence is that, for a given object, the majority of the turns of the spinning satellite will not correspond to a transit of the object and thus no observation will be acquired for this object. Most of the observations are not done and medium and large gaps occur during which the field of view is not directed towards the star. This effect is dependent on the position of the object on the sky (Gaia Collaboration et al. 2016b). Therefore, the sky sampling is markedly non-uniform (see section 5.2 of Gaia Collaboration et al. 2016b).
Practically, this generates in the spectral window a power leakage from the main peak at d-1 towards other frequencies. This effect is illustrated in the other panels of Fig. 1. We produce typical time-series of Gaia RVS with the GOST tool222https://gaia.esac.esa.int/gost/. Fig. 1 second panel represents a typical spectral window of a time-series containing 26 observations, the third panel 19 and the bottom one 14 observations (against the observations for the first panel). This small number of observations induces an aliasing but the aliasing typical of the Gaia sampling is not leaking power towards a very few other frequencies but towards numerous frequencies. Besides the main peak, there is a leakage towards a forest of frequencies. The contrast between the parasitic peaks and the main one is decreasing with the number of data points (following a for the semi-amplitude). Typically, pseudo-aliasing can occur at any . If existing, the value of is most probably located between and let say 3 d-1. The structure due to the motion considered in the top panel of Fig. 1 almost totally disappeared. In the power spectrum of the signal, the combination of noise and of aliasing is going to generate parasitic peaks at frequencies that mainly depend on the position of the object on the sky. Consequently, the time sampling is complex, and the aliasing is dependent on the star being observed.
Looking at the third panel of Fig. 1, one can notice marked aliasing at and at . Therefore, in a few cases, noise could conspire to shift the largest peak position from the progenitor peak by this amount, thus producing pseudo-aliases of the progenitor that could dominate. However, in the present case, predominant aliasing seems to be restricted to small , i.e. to small shifts of the progenitor peak. Other values of will be associated (or not) to other regions of the sky. Large shift aliasing is clearly possible but rather for values . This should have no impact in the framework of spectroscopic binaries.
Here above, we have shown three examples illustrating the strong leakage characterising the Gaia sampling (for RVS). An alternative view would be to compute some kind of typical spectral window. We obtain such a result using simulations. We generated 65536 positions randomly and uniformly distributed over all the sky. For each position, we extracted from the GOST software the series of observational epochs typical of this particular position and corresponding to the Nominal Scanning Law (NSL) for RVS. On purpose, we discarded the first month of scanning which was characterised by the Ecliptic Pole Scanning Law (EPSL). For the latter, the precession of the satellite was stopped and the scanning law is more relevant to the scanning due to the spinning of the satellite except for a slow rotation due to the orbital motion around the Sun. We thus neglect the EPSL to concentrate on the effect of the NSL. For each simulation, we derive the corresponding amplitude spectral window (for the RVS instrument). Figure 2 illustrates (in red) the median of the amplitude spectral windows. Despite the fact that we restricted our simulations to the NSL, the aliasing due to the spinning of the satellite is clearly visible. However, even the peaks at d-1 (and at d-1) are rather low compared to the peak at d-1, at most 30% (in amplitude). We also plot in Fig. 2 (in black) 32 randomly selected amplitude spectral windows. Clearly, the dispersion of the various realisations of the amplitude spectral window is rather large. Also, near the possible alias frequencies, the amplitude spectral window could present values nearing zero. This confirms the fact that in most cases, the spinning pseudo-aliasing will have little, if any, impact on the global aliasing. Although a few positions on the sky could still exhibit the effect of the spinning aliasing. Of course, if, over the 34 months, the EPSL scanning is dominating the definition of the times in the analysed time-series, the effect will increase exhibiting a stronger spinning aliasing. We must insist here on the fact that no signal is analysed in the present section; the dispersion of the amplitude spectral windows is originating from the randomness of the selection of the sky positions and from the variety of nature of the aliasing as a function of the latter. For any time-series, the observing time could be used to pre-calculate the particular spectral window independently of the signal.
4.2 Search for the periodicity
When a periodicity is present in a time-series, the periodogram exhibits a peak at the corresponding frequency (the progenitor), but other peaks are also present at the location of the pseudo-aliases of the progenitor. Contrary to the case of even sampling, the actual sampling generates pseudo-aliases that are in principle lower than the progenitor peak. However, the combination of the aliasing with background noise or other phenomena (blending) induces the fact that the choice of the true progenitor among the aliased peaks is never fully secure.
The total span of time of the DR3 time-series is at most of 34 months and thus the largest periodicity accessible from the point of view of the Fourier corresponds roughly to a frequency of 1/1000 d-1. This value could actually be larger taking into account the gaps generated by the actual sampling law. The constraints should thus be expressed as a function of the actual length of the time-series and thus as a function of . It has been decided that it would be adequate to limit the range of investigation for periodicities to since beyond the possible orbit is certainly not well determined. However, from the practical point of view, we decided to limit the range for period search to although no orbit are computed for periods located in this extension (see Sect. 6.3 for more details). The adopted frequency step is actually where is the actual total time-span of the individual time-series. This represents an oversampling of a factor one hundred that gives directly access to the amplitude of the signal as being the height of the peak. From the previous section discussion, we decided that the adopted largest frequency to be explored is 4 d-1. This choice is rather atypical since it is chosen to correspond to the spinning period.
If the star indeed varies in a periodic way, the largest peak in the periodogram is an estimate of the possible period but definitively not a proof of its existence. One should consider that a periodicity is proven if a minimum of three identical cycles are observed. Although the long periods are indicative, they are associated to large errors due to the small number of cycles included in the computation.
The method used is the one of Heck-Manfroid-Mersch (hereafter HMM: Heck et al. 1985; Gosset et al. 2001), which advantageously replaces the classical Fourier method (see e.g. Bloomfield 1976) in the case of odd sampling. In principle, the highest peak is assumed to be the progenitor one and to correspond to the orbital period. This last point is particularly true for the case of RV orbital motion, since most of the power is gathered in the fundamental frequency; the power of the peaks corresponding to high harmonics is gradually decreasing in the case of the Keplerian motion. This is due to the fact that RV curves always present one simple cycle during the orbital period. On the contrary, this is usually not the case for light curves of eclipsing binaries and ellipsoidal variables where most of the power is gathered in twice the orbital frequency, corresponding to the average depth of the minima, whereas the power at the orbital frequency is dominated by the difference between the primary and secondary minima and thus is much weaker. In their paper, HMM introduced the statistic which is a power related statistic (see Eq. A.11 in Gosset et al. 2001). At a predetermined frequency and under the null-hypothesis of Gaussian white-noise, it has the nature of the ratio of two independent statistics and it behaves as a Snedecor F statistic with 2 and degrees of freedom. It is reminiscent of a ratio of two quadratic forms, similarly to the use of nested models (here a pure cosine model versus a constant term). Since the global power is proportional to the variance of the white-noise process, it is mandatory to normalise the power by a function of this variance. This is precisely the purpose of the statistic.
The periodogram is explored and the frequencies corresponding to the various peaks that are above a certain threshold are further retained. This threshold is defined as the limit above which the highest peak can be considered as significant at a certain probability level. This is further described in the next Section. We keep in a list up to one hundred candidate frequencies. This is more than the amount of independent ones in the classical Fourier. If too many peaks are listed, only the 100 highest ones are kept. For each of these selected candidate frequencies, a fit is made of an eccentric Keplerian orbital model. At this stage, there is a possibility to add one additional candidate frequency coming for example from CU7 (e.g. on the basis of a photometric curve). In DR3 in particular, this possibility is only used for the subset of solutions of the type EclipsingSpectro combined with SB1 solutions (see section 7.6 by Siopis & Sadowski in Pourbaix et al. 2022; Siopis et al. 2024).
4.3 Significance level of the period
Thus, to search for the period, a Fourier method is used. In power spectra, the peaks in the periodogram are quite numerous. At a fixed frequency and under the null-hypothesis of white-noise with known variance, the height of the peaks is distributed as a decreasing exponential ( with a degree of freedom of two). A test must be made to decide if a candidate frequency is significant by computing the significance level which is the probability under the null-hypothesis of a white-noise process of unknown variance to observe a peak at least as high at a predetermined frequency: this is the above-mentioned F distribution. Since the algorithm is considering the highest peak over all frequencies, the adopted significance level is the probability under the null-hypothesis of white-noise to observe the largest peak at least as high. The actual statistic is thus
If the significance level is very close to zero, the peak could not be considered as generated by the noise and thus a high suspicion is permitted that the outstanding peak is due to a deterministic periodic signal. In the case of even sampling of the time-series, such a significance could be easily computed. However, in the case of irregular sampling and particularly in the case of the Gaia sampling, the situation is much more complex. Unfortunately, no rigorous analytical formula actually exists for these precise circumstances. Consequently, to test for the significance of the period, the procedure relies on the use of Monte-Carlo simulations performed in the conditions of the Gaia RVS sampling. In principle, this should be performed individually for each object characterised by their own sampling, but this is not affordable for such a large number of sources.
The first dependency of the statistical distribution is on the number of data points in the time-series. The secondary dependency concerns the actual distribution in time of the data points. The dispersion of the distribution decreases with increasing. Therefore, we first attempted to perform such simulations for each of the possibilities on this number . Keeping this number constant, it is easy to notice through the simulations that the statistical distributions of the highest peak at fixed remain quite similar whatever is the actual distribution in time of the data points (i.e. over the various simulations). This is particularly true for the interesting regime where the significance level (SL) is lower than 0.2, but also when is getting large (see Sect. 4.4). Therefore, various Monte-Carlo simulations were performed and the value of the height of the highest peak associated to various fixed significance levels has been computed. From these simulations, a look-up table is built that is used by the algorithm.
The probabilities (conf_spectro_period) given in the data model (DM) are . Despite the given name, this is thus the probability of rejecting the null-hypothesis of white-noise and is by no means the probability to have a periodicity. Independently, it should be made clear that the frequently encountered notion of significance in the database has nothing to do with the presently defined significance levels. The significance rather corresponds to the value of the parameter over its uncertainty. Such a significance is also used in the spectroscopic channel in the particular case of the semi-amplitude where it is defined as .
4.4 Build-up of the look-up table
The computation of the empirical function in our software can either be performed on-line (on an individual source basis) or read from a look-up table. The on-line computation, being necessarily more accurate but requiring an important amount of computational time (more than 10 minutes per source), was no longer considered by the processing centre after it threw time-out exceptions during pre-operation tests. Therefore, for DR3, we adopted the method of the look-up table which is approximate since it is based on typical average properties of the actual samplings (with data points) and not on the actual individual sampling.
To build the look-up tables off-line, we gathered the observation dates from final RV curves of RVS sources used as standards, kindly provided by CU6, and we grouped them on the basis of their number of valid RV measurements. We preferred to rely on concrete observation dates rather than on the predicted NSL dates provided by the GOST tool in this exercise. Indeed, there are significant discrepancies between the NSL prediction and the final actual observation dates. The reason is that many spectroscopic observations fail to provide a valid RV measurement at the end of the CU6 spectroscopic processing. This implies a sparser sampling.
For each source, we computed 10000 HMM periodograms of Gaussian white-noise realisations on the very same date sampling as that of the considered source, by adopting the same frequency step and limits as in the operation pipeline. For each periodogram, we retained the largest value of the , the so-called statistic. Then we merged all the values and ranked them by increasing order and build up an Empirical Cumulative Distribution Function (ECDF) with the 10000 values of the statistic. At fixed, each realisation of the sampling provides a particular ECDF function. The chosen length 10000 of the ECDF allows us to assign false alarm probabilities with a precision better than 0.5 % at 95 % of confidence.
The look-up table contains the set of median ECDFs for every given amount of valid transits. The real distribution of the number of observations per source is reflected in the number of ECDFs accounted for in each median with this scheme (i.e. sources with 17 valid observations, which are the most common in DR3, have a median ECDF covering the greatest number of individual ECDFs).
Figure 3 shows the simulated ECDFs for = 10, 17 and 30 and their corresponding median. As time sampling approaches the similarity of the samplings for large values, the ECDF dispersion decreases and their median approaches the regular time sampling CDF. The values in ECDFs also decrease when increases. Indeed, fitting a cosine function to a large sample of pure noise will not significantly improve its variance, unlike in small samples where perfect fits may occur, especially when high frequencies are explored. In addition, the signal-to-noise contrast increases with .
When calculating the periodogram of an input time-series with the length , its corresponding adopted ECDF (on the basis of the value) is read from the look-up table. The table has a comprehensive coverage of the number of transits, without gaps from = 10 to 72, but there are some entries for large values that are not present. In this case, the algorithm opts for the closest lowest input.
5 The calculation of the orbital solutions
Our main task is to derive simple orbital motion based on Keplerian models. The general model (see Eq.1) is highly non-linear and usually necessitates an a priori approximate knowledge of the true values of the parameters to fit. This is not possible in the framework of a pipeline treating so many objects. To define the proper method to use, we investigated numerous possibilities; they are evoked in Appendix A. Despite the many trials performed, it turned out that approaches based on classical Levenberg-Marquardt iterations were much more robust and precise, if appropriate precautions are taken. As explained in Appendix A, the alternative idea to perform simulations were rejected for excessive processing time reasons.
The Keplerian fitting, even at fixed period, is time-consuming if it is to be performed over a vast set of periods: as for example in a genuine Keplerian periodogram (see e.g. Cumming 2004; Zechmeister & Kürster 2009, their sections 2.3 and 5, respectively).
Therefore, we decided to strew subtleties over our procedure in order to render it much more robust. The first step is to search for periodicities using a periodogram (HMM method) as described in Sect. 4. This corresponds to a fit of sine/cosine, which is a linear process (at fixed frequency). This is particularly convenient: even if the RV curve is highly eccentric, and thus the power of the signal is spread over several harmonics, the power related to the period of the orbit is still dominating the noiseless periodogram. If the number of cycles is sufficient, the periodogram will always present increased power at the correct period. However, the detection of a periodicity does not imply that the motion is Keplerian. Therefore, after the establishment of a list of candidate periods, it remains to fit the Keplerian model for each of them and to deduce the orbital solution and corresponding parameters that better fit the data. We first perform a Keplerian fit that is linearised at the maximum (but not entirely) and represents a preliminary fit, not definitive but robust. The actual least-square fit is performed with the true anomaly as the independent variable.
The Fourier periodograms have the reputation to be inefficient at addressing the case of eccentric RV curves (for , see e.g. Endl et al. 2002; Cumming 2004). We tend to think that this is due to the fact that, for eccentric orbits by opposition to the circular case, not all the data points contain the same amount of information. In fact, for very eccentric orbits, the data points related to observations around the maximum separation are more determining, in particular for the semi-amplitude. If such data points are missing, have not been acquired, this will have major impact at least on the determination of the value. Consequently, this will have a huge importance on the detectability of the signal: more abruptly, we cannot detect something that is not observed. Actually, this is not a problem of ability of the Fourier periodogram to address eccentric orbits, this is a problem of fair sampling of the RV curve. The correct way to interpret things is that if a period is detected, we are authorised to have some confidence in its existence. Therefore, the nature of the Fourier periodogram improves the reliability and robustness of the algorithm at the expense of a difficulty to have a proper estimation of the completeness. This is a choice we made.
We finally decided to adopt the method fully described here below, which is based on fully analytical and classical numerical developments. After several tests, the adopted method appeared to be faster, more stable, trustworthy and robust.
6 The main NSS-SB1 processing chain
6.1 Global structure of the chain
Five Java packages operate in sequential order in our pipeline. These are InputChecker, PeriodFinder, SB1Refiner, SB1CRefiner and TrendSolution. The package InputChecker reads the input time-series in the NSS data model and copy them into arrays of primitives that are shared with the entire pipeline. This reduces the number of modifications required after every data model update. It also performs sanity checks, data cleaning and some other modifications. Finally, it passes the rearranged data to the four remaining processing packages.
The second package (PeriodFinder) is working in two steps. Its first step is to establish a list of candidate periods that are significant. Its second task is to test each of the significant candidate periods by fitting the corresponding Keplerian model to the data, and to select the best period on the basis of the comparison of the various models. The best of these models represents a first approximation of the orbit. The third package (SB1Refiner) has the charge to start from the selected approximate solution and to refine it. As part of this computation, it derives the variance-covariance matrix of the fit. The fourth package (SB1CRefiner) addresses the case of circular orbits when the eccentric solution generates a badly configured variance-covariance matrix, and tries to adopt the best compromise between a circular solution and a singular eccentric one. After these computations, a decision is to be taken to adopt the most adapted periodic orbital solution. The fifth package (TrendSolution) deals with the RV time-series exhibiting a trend as a function of time (see Sect. 3.3). This means time-series that can be fitted by polynomials. These are acceleration models. After the execution of all these packages, a choice is made to select the solution that presents the best fit: SB1, SB1C, TrendSB1 or StochasticSB1.
The five packages are fully described in the respective five Sects. 6.3.1, 6.3.2, 6.3.3, 6.3.4 and 6.3.5. The flow chart of the full NSS-SB1 processing is described in Fig. 4 along with the relations between the five packages mentioned above. It is interesting to notice that after the upper violet lozenge that contains a test on (that will be defined in Sect. 6.3.1), there is a bifurcation towards two sub-chains that receive the data. One sub-chain is treating the objects exhibiting a periodicity that could conduct to an orbital solution. The other sub-chain treats the cases where no periodicity is involved (only trends). From the purely logical point of view, they are working in parallel although from the practical point of view, they are executed in sequence. During the processing of the full NSS-SB1 chain, some flags are raised to retain more information about the path followed by the time-series, as well as to point out a few characteristics that could be useful for further treatment. In Sects. 6.3.1, 6.3.2, 6.3.3, 6.3.4 and 6.3.5, the flags are identified by their number and a full description of their nature is given in Sect. 6.5.
6.2 Generalities about the fits
After the InputChecker intervention, the four other NSS-SB1 processing packages are performing a set of linear or non-linear fits, involving linear algebra operations on moderate rank matrices (rank of matrices never exceeds 6). A home-made linear algebra toolbox, based on the Eigenvalue and Cholesky decompositions, has been developed for that purpose.
Although the former is known to be time-consuming compared to QR or Cholesky algorithms, it gives the best estimates of the condition number of involved algebraic systems, and thus ensures a better robustness. This is why it is applied all over the pipeline except for trend analysis, for which we adopted the Cholesky decomposition instead.
Moreover, thanks to the eigenvectors, singular algebraic systems can still be solved by the Moore-Penrose pseudo-inverse technique (Press et al. 2007, Sect. 2.6.2). This is extremely useful because it solves the discontinuity problem between circular and eccentric models. Indeed, when eccentricity approaches zero, and parameters become progressively redundant (Sterne 1941), and thus the eccentric model becomes overstated, too powerful and thus prone to overfitting. The boundary between the two regimes, circular and eccentric, can not be set by a fixed value of the eccentricity since it is governed by the signal-to-noise ratio of the input time-series, and thus depends on many parameters such as the sampling, the amplitude of the signal and the involved noise (Lucy & Sweeney 1971).
Eigenvalue decompositions of matrices with rank smaller than 5 are performed analytically to improve robustness and reduce the processing time. Cardan-Tartaglia and Ferrari methods are applied to compute eigenvalues of rank 3 and 4 matrices, respectively.
The Cholesky decomposition is very suitable for the trend analysis in which we successively increment the polynomial degree. Since the polynomial order is nested into the one, we just need to update the order curvature matrix decomposition to the new dimension instead of completely recomputing it.
Our home-made linear algebra toolbox defines the same singularity threshold for both decompositions. It is chosen equal to 10 2.22 the square of the input matrix size. The value 2.22 is the lowest non-zero double precision number in current processors. The threshold corresponds respectively either to the ratio between the extreme eigenvalues (in the eponymous decomposition method), or to the square of the ratio of the extreme diagonal elements of the Cholesky matrix.
6.3 Processing steps
6.3.1 Ingestion of the data
All stars that present RV variations according to the following criteria rv_chisq_pvalue0.01 and rv_normalised_gof4 are processed by the pipeline starting with the InputChecker described here. Other details about these input data were already provided in Sect. 2.
NSS input data are requested from the Main Data Base (MDB) and copied into StarObject Java objects during data ingestion. A StarObject contains the single star astrometric solution (the reference epoch, the mean position, the proper motions, the parallax, ruwe and astrometric_excess_noise), the atmospheric parameters as measured by CU8 (Coordination Unit Astrophysical Parameters, Andrae et al. (2023); Recio-Blanco et al. (2023)), and the light curve classification along with the orbital period if the corresponding source is suspected to be an eclipsing binary by CU7 (Coordination Unit Variability Analysis, Mowlavi et al. (2023)). In addition to these mean data, a StarObject contains an array of Transit Java objects, where each Transit contains the relevant epoch data such as the source position, the integrated flux in the , and bands, the RV (along with its uncertainty) as well as some related quality flags. Observation dates at the centre of the astrometric field are given in BJD. With the satellite spin, an observation at the centre of the RVS CCDs is performed roughly 75 s after the observation at the centre of the astrometric field: this delay has been neglected. The observation date, the RV and its uncertainty are extracted from each Transit and copied into arrays of primitives for limiting code modifications required after every data model update. Those that are not finite numbers are filtered out. Observation dates and error bars must be positive numbers, while RVs must fall within the range of km s-1. The length of the time-series is expected to be larger than or equal to 10 and if it is not the case a flag is raised (flag nr.8, see Sect. 6.5) and the star is rejected from further processing.
After these sanity checks, we start by cleaning up non-valuable entries. A Transit is discarded when its RV is detected by CU6/MTA as an outlier, either on its own or together with the rotational velocity or the integrated flux in the band or both. In addition, we also discard the Transit if it exceptionally corresponds to a composite spectrum (see Damerdji et al. 2024). The resulting time-series after cleaning contains data points distributed over a genuine individual time span (the difference between the two dates of observations at the extreme of the interval).
Finally, we perform a few modifications. We correct uncertainties for sources with large ruwe, as described in Sect 2. The reference epoch J2016.0 (BJD 2457389.0) of the astrometric solutions, is subtracted from the observation dates for reducing the correlations between the derived solution parameters. A cleaned time-series of RVs, still containing the required number of 10 observations, is shared with the whole software for the rest of processing, whilst, in the opposite case, it is rejected from the pipeline, and the corresponding flag is raised (flag nr.10, see Sect. 6.5 for flag descriptions). The rejection from the pipeline is done by outputting an empty solution containing the flag and no parameter values. Such an empty solution is for internal use and statistical analysis. It is not further considered for the final output of the chain.
6.3.2 Inventory of the periods and approximate model
The package PeriodFinder is first establishing a list of candidate frequencies. The time-series of length is analysed thanks to an HMM periodogram, as explained in Sect. 4.2. The package reads the ECDF corresponding to (see Sect. 4.4) and establishes a threshold above which the peak is considered significant against the null-hypothesis of white-noise. All the peaks in the periodogram that are above the threshold value have their position in frequency entering the list. This is clearly an approximation since the test is constructed on the rigorous concept that we are looking at the highest peak only (see Sect. 4.3). No fully sound counter-indication to this assumption has been identified yet. Up to 100 frequencies/periods are accepted in the list. No additional smaller peak is considered above this amount. If the object under analysis is an Eclipsing Binary, we also add to the list the value of the period issued from the photometry. We test the significance of any available external period (i.e. in DR3 from CU7) before adding it to the set of candidate periods. For DR3 the periods entering the list were selected on the basis of an adopted threshold probability value conf_spectro_period larger than 0.95 ( smaller than 0.05). If the list of significant candidate periods is empty, no full orbital solution is possible. The relevant flag is raised (flag nr.13).
Since the aim of the pipeline is to compute orbital solutions, it is worth trying to limit the possibilities to Keplerian motion. For each of the frequencies in the list, a Keplerian model is fitted. The pipeline is not performing the fit in phase (defined by the candidate frequency under test) but in true anomaly. In this domain, the RV curve to fit is a cosine. This of course necessitates the knowledge of and . Since the eccentricity is a strongly non-linear parameter, the pipeline proceeds as follows. Starting from a two-parameter grid of and , we minimise, using a Levenberg-Marquardt procedure, the quadratic sum of the differences between the observed RVs and the model ones (). The vector of model RVs (coming from an expansion of Eq. 1) is given by
where and . From these and , the Campbell parameters and are easily extracted. In Equation (10), , , and appear in a linear way. They can be optimised through an ordinary linear least-square regression and this for every set of trial . This system of three linear and three non-linear parameters, can be efficiently solved with the separable variable technique (Golub & Pereyra 2003). Indeed, we optimise three non-linear parameters only: , and (1/). Consequently, the dimensional space of variables is reduced, leading to more robust and rapid convergences. The pseudo-inverse implemented in the linear algebra toolbox smooths singularities near zero eccentricities. The rank of the involved linear systems is smaller than or equal to three, allowing them to be solved using more precise analytical computations. We give in Appendix C a summary of the involved computations. The minimisation is limited to 1000 iterations for each of the 8 considered starting points, and this for every candidate frequency . The solution is taken into account for further comparison, even when the maximum number of iterations is reached. The relative convergence of the 8 different descending tracks are analysed to estimate the robustness of the convergence (the arrival point should be the same in a perfect world). Ultimately, the best solution is selected for the considered candidate frequency. This algorithm is quite robust and provides adequate fits for orbital curves up to eccentricities of 0.995, well beyond what is necessary for the majority of the objects. The procedure is executed for every candidate frequency and gives a of the fit that is translated into an statistic. The smallest value identifies the best fit and the best set of parameters, including the period. The algorithm is efficient at finding the best minimum, but the fit is performed in two steps which is clearly an approximation concerning the cross-covariance terms between families of parameters (linear versus non-linear). Therefore, we need to refine the approximate model.
The use of a fit in the true anomaly space ensures that no analytical solution could present a bad behaviour in the phase space. Zones in the phase diagram where no data are present, the phase gaps, could exist in some pathological cases. This problem is tackled in Sect. 7.1.
6.3.3 Final fit of the eccentric orbital solution
After having identified the best approximate orbital solution, the global solution refiner comes into play (SB1Refiner). It is a Levenberg-Marquardt minimisation based on first order derivatives as given by Schlesinger (1910). It returns the eccentric model as well as its variance-covariance matrix. The algorithm thus performs, at the best minimum (the approximate model), a classical least-square fit over all the parameters with all the non-diagonal terms being taken into account. This has the advantage to provide sound (but classical) estimations of the uncertainties on the parameters. The variance-covariance matrix may be singular due to an over-determination of the model. In that case, a flag (flag nr.15) is written before activating the following package. If the matrix is not singular, the SB1 solution is forwarded to the final decision module. If the iterative process does not converge after 1000 iterations, a flag is raised (flag nr.14). If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.18).
The eccentricity is a parameter that suffers from a bias related to the noise propagation. The RV noise could enforce a fitted value for that is much larger than the true one. This effect is sometimes compensated by enforcing for the solution when the eccentricity is not significant. This is common practice in the literature to apply a test and to decide to go circular (i.e. to decrease the number of free parameters) for low-amplitudes and noisy cases. This has the pitfall to enforce the circular solution for objects whose eccentric solution is only marginally good (non-significant eccentricity). This produces a deficit of non-zero but small eccentricity solutions. The present algorithm does not do that. It never decides that an orbit could be forced circular. It is allowed to possibly (but not necessarily) go circular when the eccentric model cannot lead to a well-behaved non-singular variance-covariance matrix.
6.3.4 Fit of the circular orbital solution
The module described here is the SB1CRefiner. The best approached period obtained in the PeriodFinder package is refined and fitted along with the linear parameters , and to fit a circular model to data (see equations in Sect. 3.2). When a circular orbit is selected, corresponds to the maximum of velocity by definition. A flag is written when the derived variance-covariance matrix of the circular solution is singular; this is flag nr.16. If the iterative process does not converge after 1000 iterations, a flag is raised (flag nr.21). If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.19).
In some sense, the less complex circular model is nested into the eccentric model through the eccentricity. Indeed, if we put into the eccentric model, we obtain the equations of the circular model (except for the trivial change of the definition of ). It is thus tempting to opt for a classical statistical test for nested models to choose between the two. This is actually what is done. However, this is not perfectly rigorous since, although the circular model is linear, the eccentric one is not. Therefore, the number of degrees of freedom of the is not rigorously defined. In any case, as a first approach, we decided to use the classical nested model statistic F (also described in Lucy & Sweeney 1971, i.e. Lucy’s test). We thus test here the eccentric orbit model (as computed in SB1Refiner but corresponding to a possibly unreliable solution having necessarily at least one of the three flags nr.14, nr.15 or nr.18 raised) against the circular orbit model taken as null-hypothesis. The eccentric model is thus retained when the Snedecor-F test significance level is lower than (or equal to) 0.05. This decision is taken regardless of the quality of the variance-covariance matrix of the circular model. The SB1CRefiner thus produces SB1 and SB1C type candidate solutions. When Lucy’s test has been applied, flag nr.22 is raised.
6.3.5 Fit of trends
Not all the observed RV curves correspond to a true periodicity nor to a pseudo-one. One widespread alternative possibility is the existence of a mere trend: the data as a function of time can be represented by a polynomial function of (see Sect. 3.3). The last computing package TrendSolution applies polynomials to data to search for acceleration models (trends) rather than periodic orbits. We start by applying a zero degree polynomial (corresponding to a noisy constancy), then we successively increase the order at each step. The lower degree polynomial is thus nested in the new one. The choice between the lower and the higher degree polynomial is done through a classical Snedecor-F test on the ratio of the improvement of the to the value of the for the higher level model, since the polynomials given above are indeed nested models. A Snedecor-F test for nested models is thus applied. The higher degree polynomial is accepted by the Snedecor-F test for a confidence larger than 0.99865 (3- probability). We stop iterating at the 4th order, or when the higher-degree model exhibits more than one extremum over the time baseline. This last criterion has been adopted in order to avoid any possible overlaps with the orbital models with periods between and (see below). Most of the objects entering the trend category are probably long-term periodic curves but with a timescale much longer than the present time span of the mission (). The code is built to accept polynomials of degrees 1, 2, 3, and 4 but in DR3, only degrees 1 and 2 were encountered (at the selected probability threshold). A flag is written when the derived variance-covariance matrix of the trend solution is singular; this is flag nr.17. If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.20). A regrettable mistake occurred in the code and, for the last operational run, the flags raised were not the correct nr.17 and nr.20 but instead flags nr.15 and nr.18, respectively. This problem has been corrected for DR4.
The coefficients of the polynomial are included in the database but the work is performed in an Hipparcos-like formalism (see section 2.3.3, ESA, 1997). The coefficients saved in the database are related to the derivatives of the velocity curve and are expressed with respect to the DR3 reference epoch (J2016.0, equivalent to BJD 2457389.0).
We thus decided to express the trends as a function of the various derivatives. The simple mathematical model is
where is the time (in days) relative to the reference epoch. We thus fit to the time-series the polynomial
with potentially the fixed to zero. In the case of the first degree polynomial ( = 0), we adopt the following rigorous formulae
The and parameters are not correlated if . This is approximated by subtracting to the BJD dates the reference epoch (J2016.0) as a proxy of the middle of the current mission. Practically, not all actual time-series are centred on this value since some transits might be missing (e.g. more numerous before the mid-epoch or alternatively after it). This is in any case still an approximation since it is implicitly assumed that the distribution of the sample times are homogeneously apportioned. This is certainly not true, particularly if one includes EPSL data.
In the case of the second degree polynomial ( fitted), we note that the fitted polynomial can be expressed as
where is the total duration of the mission treated in the framework of DR3. We thus adopt the approximate formulae
The term shifting by has been introduced such that the contribution by the second derivative is tending to be neutral over the interval . This of course induces a shift on . The first derivative is considered as averaged over the full interval. This property is here also true for an homogeneous distribution of the observing times. A strong disparity of the observing times could have introduced some biases.
Initially, the threshold was chosen at a probability level of 0.95 (following the Snedecor-F test of nested models) and polynomials of first, second, and third order (with only one extremum) were found. Following validation, we decided to be more conservative and we instead adopted a new threshold (0.99865); this induced a downwards cascade among the different degree categories and restricted the palette of solutions to first and second order polynomials. But the decision occurred too late and no more operational run took place. Therefore, we kept formally adopting the threshold of 0.99865 and decided to redistribute the various solutions and to blacklist the objects that, in case of a rerun, would have been discarded as trend solutions. This certainly induced the loss of a small number of objects in other categories.
The acceleration identification as performed in DR3 can be affected by false detection, revealing the subtle effect of an unfortunate sampling on the polynomial fit. In the upcoming data releases, changes to the algorithm will be made to ensure that trend models are first detected in the lowest frequency region of the periodogram, as introduced in Boulkaboul et al. (2022).
6.3.6 Selection of the best solution
Therefore, the SB1Refiner and the SB1CRefiner packages provide an SB1 or SB1C candidate solution. They could in some cases provide no solution if no period is found. The TrendSolution package delivers in parallel polynomial fits and thus candidate trend solutions (of degree 0, 1 and 2). It is now necessary to decide which model is performing the best interpretation of the data. The adopted orbital, circular or eccentric, candidate model is ultimately compared to the less complex best trend candidate (i.e. the null hypothesis is that the model is a trend). Acceptance of the orbital model is granted when the Snedecor-F test confidence is higher than 0.95. If this is the case, the final delivered solution is the selected SB1 or SB1C. If this is not the case, the trend solution is adopted. If the degree of the fit is 1 or 2, the object is entering the delivered output as a TrendSB1 solution. If the degree is zero, we have to deal with a constant RV with excess noise compared to the expected uncertainty. In this last case, the solution is output as a StochasticSB1. Thus, when no significant periodic variability nor trend are possible models, the algorithm associates the star to a stochastic solution. These are objects that have very little chance of being associated to a mere orbital or trend motion. As explained elsewhere, some badly sampled orbits and high degree multiple systems could fall into this category, perhaps momentarily. Some of them will certainly benefit from an increase in the number of transits. These StochasticSB1 solutions are not provided in the DR3 catalogue. The correct choice between an orbital solution and the trend one is particularly difficult because the independent variable is time for the trends and let say phase for the well covered orbits. The time entering in a non-linear way in the Keplerian equations and the first selection of the period being based on a HMM periodogram, the statistical Snedecor-F test is not fully correct.
Actually, the procedure described above has been applied when orbital solutions presented a period lower than ; it concerns the regime where the orbital motion is expected, when selected, to be convincingly robust. For longer periods, the folding with the period does not permit to test the perfect overlap homogeneously at all phases. We found it preferable to adopt an alternative procedure. If the orbital period is comprised between and , we decided to enforce the choice of the orbital solution. For periods larger than , i.e. outside the zone of possible convincing orbital motion, we decided to enforce the choice of the TrendSB1 solution. This last point explains why we decided to explore the periodogram in the range between and .
6.4 Quality of the solution
For each object processed, along with the various resulting model-parameters, the code also computes some statistics and various entities that might help in estimating the quality of the solution. The first statistic is the one given by
where is the classical statistic of the least-square fit and where is the related number of degrees of freedom. This statistic was first introduced by Wilson & Hilferty (1931). Since the eccentric fit is non-linear in nature, we performed various simulations to test the estimation of the number of degrees of freedom to use. Quite surprisingly, it turned out that, for the Keplerian motion, the number of degrees of freedom is very close to the number of data points minus the number of parameters (similarly to the linear case).
Besides the , the code also computes the which is the largest gap in the phase diagram where no data points are present. It also computes the solution efficiency whose definition can be found in Eichhorn & Xu (1990), and is corresponding to a measure of the correlation between the computed parameters. A total absence of correlation, and thus a diagonal variance-covariance matrix, will give an efficiency of 1. Values of the efficiency tending to 0 indicate strong correlation. Practically, an efficiency identically equal to 0 is not possible since this is associated to a perfectly singular matrix to invert.
In addition, we also compute the solution significance defined for spectroscopic data as . Finally, we should add the significance level of the periodicity as described in Sect. 4.3 and expressed as a confidence ().
6.5 Flagging
Beyond the determination of the various parameters, the code also produces various flags that allow the user to have a good understanding of the results and to select some solutions that could be better or more secure than others. The flags generated by our NSS-SB1 processing are incorporated into a more global set of flags issued from the global NSS processing. We will restrict the discussion here to the set of flags directly related to the SB1 processing chain, with a few information linked to flags issued from the SB2 processing that is fully described in Damerdji et al. (2024). All the NSS flags are included in a unique long variable, whose every bit acts as an individual logical flag. All the flags involved in the NSS spectroscopic processing are enumerated in Appendix I. As an example, the SB1Refiner processes the time-series of an object and produces a singular matrix. The flag nr.15 is thus raised, giving to the global flag an additive value of 32768. The object is then processed by the SB1CRefiner and, let say, it concluded to a singular matrix associated to the circular orbit. The flag nr.16 is thus raised (associated to a value of 65536). Consequently, the object is associated to both flags nr.15 and nr.16 raised, giving together a value of 32768+65536=98304. The object could appear in the catalogue as a TrendSB1 with the two above-mentioned flags raised. Although the majority of the flags are important for the decision tree of the algorithm, some of them are purely indicative. This is the case of the flag nr.24 that draws the attention on the fact that the attributed eccentric orbit appears, in the eccentricity versus log diagram, outside the zone where the circularisation of the orbit could have partly taken place. The borderline used is inspired from the one defined by Halbwachs et al. (2005) but modified by us333Using the formalism of Halbwachs et al. (2005), we estimated a value of on the basis of a fairly crude analysis of the full SB9 catalogue (Pourbaix et al. 2004). This estimate has been deduced without proper treatment of the dispersion of due to noise and is considered as a conservative border. The purpose is to flag extremely pathological cases to trigger the cautiousness of the general user.. Flag nr.25 just indicates that the selected adopted period was compatible with one period suggested by the photometry. Flag nr.22 informs that the Lucy test has been applied. Flags nr.11 and 12 concern SB2 objects that turned out to be impossible to treat as SB2 either because the amount of RVs was too small for the secondary (but not for the primary) or because the period found by the SB1 chain was not coherent with the one of the SB2 chain (see Damerdji et al. 2024). Flag nr.26 is raised by the SB2 processing to indicate that the pair of RVs of the SB2 measurements were not anticorrelated (more details are given in Damerdji et al. 2024).
7 The post-processing
The post-processing described here further cleans and distributes the results of the NSS-SB1 main processing chain. The detailed flowchart is provided in Fig. 5 and additional explanations are given in this section.
7.1 Internal filtering
From summer 2020 to spring 2021, the NSS pipeline processed the Gaia data in the framework of several validation runs. At least seven of them were concerning the software linked to the present paper. This processing helped in finding remaining failures of the algorithms and in correcting them. This also helped in fine-tuning the decision tree of the pipeline by adjusting some statistical thresholds. The operational runs took place in summer 2021.
In order to maintain the quality level of the various solutions, we defined an internal filtering of these solutions that will allow keeping the very best and most secure ones. The aim is also to spot out and reject solutions that are problematic. Most of the filtering is done via the definition of thresholds. Very often, the selection is made in a statistical way and by using probability thresholds based on the distribution of some statistical variables. As always, the selection near the borders is never perfect. We detail here the internal filtering for SB1 solutions.
Since the individual (per transit) RVs are searched for in the range between –1000 km s-1 and +1000 km s-1, we found wise to test that all solutions have 1000 km s-1. This is the case but this sane test is not highly restrictive.
On the opposite, one of these important statistics is the . The statistic represents a classical goodness of fit criterion that is directly related to the objective function of the fit and to the classical statistic. The statistic is expected to have under the null-hypothesis a Gaussian distribution. Typically, above = 5, the solution is not satisfactory and presents extra-dispersion of the residuals, be it due to the data or to the model. In order to be conservative, we decided to put the threshold such that .
In the course of the NSS-SB1 processing, various flags are raised, tracking the details of the decision tree and of the problems encountered. As explained in Sect. 6.5, some combinations of flags imply that the solution is not acceptable or is presumably inaccurate or too hazardous. The corresponding selection was of course considered. Since the internal filtering is supposed to protect the results from spurious solutions and that, in the present purely spectroscopic SB-subcatalogue, no confirmation comes from other channels, we decided to be highly cautious for this step. For the SB1 solutions, we discarded objects having at least one of the three flags nr.14, nr.15, nr.18 raised. As a consequence, this suppresses all the SB1 solutions issued by the SB1CRefiner.
We further found that a conservative attitude is to restrict the palette of the solutions to those yielding a 250 km s-1 which eliminates a few cases of surprisingly large semi-amplitudes that must be further validated (a work beyond the scope of the present paper).
We also used a threshold on the solution efficiency. Any solution efficiency below 0.1 is rejected because the corresponding fit is considered as spurious (Eichhorn & Xu 1990). The solution efficiency is part of the catalogue, allowing the user to adopt a more restrictive attitude.
When the supposedly periodic RV data are folded in phase, it happens that a zone in the phase diagram is not populated by any data points. If this region is too wide, the corresponding solution could be pathological with respect to the sampling. Therefore, we choose to discard the solution with a too large where this value represents the largest interval in phase with no data points. We reject solutions leading to . The value of the gap for remaining solutions is not propagated to the SB-subcatalogue but will be accessible in DR4 with the release of the epoch RVs. The threshold value itself has been optimised on the basis of the various validation runs (see Sect. 9 and in particular Sect. 9.2). Due to a small bug in our software, a small part of the objects presenting have not been properly rejected in DR3.
The mere existence of the periodicity in the data is a strict test of the validity of the orbital solution. It is reasonable to consider values for the probability conf_spectro_period from 0.95 to 0.995 or even stricter. During the validation, it appeared a strong population with short-period orbital solutions sometimes associated to relatively large eccentricities. This is somewhat in contradiction with the expected circularisation phenomenon (or these are not properly speaking orbital solutions). Therefore, we decided again to be conservative and to put a threshold that is period dependent. We hope to improve this procedure for DR4. For periods above 10 days, we selected a lower threshold of 0.95 on conf_spectro_period (significance level 0.05). For periods less than one day, we adopted 0.995 as the lower limit. In-between, the lower threshold on conf_spectro_period is given by . The solutions associated to a value less than 0.95 are filtered out. The solutions associated to a value above 0.995 are kept. Solutions in-between are kept but a flag (nr.13, i.e. 8192, see Appendix I) is raised if they are below the defined threshold.
Another natural statistic is the spectroscopic significance defined as . This ratio should not be confused with the probabilistic quantity canonically named significance level. A high significance is necessary to adopt a solution. However, one should not forget that is itself an estimator and that statistical fluctuations could give a value of that is exceedingly small, boosting the significance to very large values. The opposite case is also true. We decided to keep the significance as a possible sorting criterion for the future data applications. Therefore, we only rejected all the solutions with a significance lower than 5. This threshold has been fixed for the internal filtering, but it could be advisable to the general user to select a value of the order of 20 to 50 depending on the details of the scientific application. An example of this way of fine-tuning the selection is illustrated in Gaia Collaboration et al. (2023a).
Finally, some solutions corresponding to values of the argument of the periastron are accompanied by a very large value of their estimated uncertainty. Conservatively, solutions characterised by a 1- uncertainty larger than 2 (in radians) are considered as problematic, since this actually means that the is properly undefined. We preferred to discard these objects. The filtering could have been more stringent but we preferred to rely on user’s discernment. As expected, the majority of large uncertainties on correspond to very small eccentricities () characterised by rather large uncertainties. However, some have eccentricities that can be considered as significant.
An internal filtering of the kind performed here above was also applied to the SB1C type of solutions (except of course for the last one on ). For the SB1C solutions, we discarded objects with either flags nr.16 or nr.19 raised.
An internal filtering was similarly defined for the TrendSB1 solutions. This last filtering was on , on the bad flags and on the equivalent of the above-mentioned gap in phase. Concerning bad flags for TrendSB1 solutions, we should have discarded objects with either flag nr.17 or flag nr.20 raised. However, these flags were never activated. We decided to filter out on the flags nr.15 and nr.18, although there could be some unfortunate interactions with the SB1 channel. We rejected solutions presenting a gap (in time) in the time-series larger than 0.3.
7.2 Combined solutions
After the spectroscopic processing described here above, the solutions are delivered to the combiner. The combiner and combinations are described elsewhere (see section 7.7 by Gavras & Arenou in Pourbaix et al. 2022) and are not extensively addressed here. This combiner explores opportunities to unveil solutions common to other channels. In the framework of DR3, two possibilities were considered.
Either the NSS combines objects from the spectroscopic channel with those issued from the photometric channel and present in the class EclipsingBinary, i.e. presenting eclipses in photometry (see section 7.6 by Siopis & Sadowski in Pourbaix et al. 2022; Siopis et al. 2024; Gaia Collaboration et al. 2023a). The main criterion for the combination is the existence of a common value for the period. If the tentative combination is successful, a global fit over the photometric and the spectroscopic data is performed. In such a case, the new (improved) parameters supersede the old one and are introduced in the catalogue in the new class EclipsingSpectro (see section 7.7 by Gavras & Arenou in Pourbaix et al. 2022). Most of the contributors to this class are issued from our set of SB1 solutions, although some other types of solutions could also be considered (SB1C, SB2, SB2C).
Alternatively, NSS can try to find objects for which the results of the astrometric channel (Halbwachs et al. 2023; Gaia Collaboration et al. 2023a) can be combined with objects having a solution issued from the spectroscopic pipeline. The relevant combined solutions and the related objects are then redirected towards the new class labelled AstroSpectroSB1 (see section 7.7 by Gavras & Arenou in Pourbaix et al. 2022). New fits are here also performed and the old values of the parameters are also superseded. We recall that, in this case, the fit in DR3 is done in the framework of the formalism of Thieles-Innes (Halbwachs et al. 2023; Gaia Collaboration et al. 2023a). This formalism implies that in DR3 AstroSpectroSB1, the combined solutions include two different values (one astrometric and one spectroscopic) for the argument of periastron. The combination with astrometry accepts several types of possible spectroscopic input (SB1, SB1C, TrendSB1, StochasticSB1).
The objects corresponding to a successfully combined spectroscopic solution are removed from the spectroscopic output. Even after the combination process, it might be that an object appears in two classes. In the framework of DR3, these are the sole allowed/activated possibilities. The criteria to consider a combination opportunity are described in details in section 7.7 of Gaia archive documentation444https://gea.esac.esa.int/archive/documentation/GDR3/ (Pourbaix et al. 2022).
If, for an object, the combination with the spectroscopic output turned out to be impossible, the sole spectroscopic solution (and the object) enters the newly created class that is eponymous of the spectroscopic solution type, either SB1, SB1C, TrendSB1 or StochasticSB1. Contrary to the first three, the StochasticSB1 class is not published in DR3. The classes EclipsingSpectro and AstroSpectroSB1 are not further considered in the present paper.
In order to avoid losing interesting objects and in a purely prototypical context, the combination was performed before the internal filtering. It could have been more secure to perform it after the internal filtering. Also, the combination of the astrometric solution with the StochasticSB1 one should certainly be a matter of concern. The StochasticSB1 solutions by essence present no actual periodicity and are not expected to be further combined (in principle). On the opposite, solutions that are purely spectroscopic and are ultimately populating the classes SB1, SB1C and TrendSB1 were necessitating increased caution and were submitted to the internal filtering.
7.3 Post-filtering
On the basis of the use of results from other origins, it is possible to further refine the set of solutions that is coming out of the chain. This is made by blacklisting spurious solutions. For a good understanding of the reader, this post-filtering is explained in details below in Sect.10.1 after the validation sections.
8 Results
DR3 is the first release providing orbital solutions for the RV time-series. As such, the present results must be considered preliminary even if the performances are already good.
8.1 Tallies
Some 33,812,183 objects have validated (non NULL) RV values in DR3. Among them some 995315 objects obey the input criteria. We recall them: RVs always between –1000 and +1000 km, brighter than 12, effective temperature between 3875 and 8125 K, and rv_normalised_gof above 4. Only the objects suspected of variability enter the orbital solution chain. Thus, an additional cut-off of rv_chisq_pvalue at 0.01 (kept below 0.01) further decreases the sample. A little more than 866 kilo-objects (866403) entered the present main NSS-SB1 processing. If we consider the SB1+SB2 processing, the total amount is 876644. The repartition of the results among the different types of solutions introduced in Sect. 3 is given in Table 1. The missing objects correspond to a lack of possible solutions. An important contribution comes from the objects having smaller than 10.
After this preliminary main NSS-SB1 processing, the internal filtering is applied, as explained in Sect. 7.1. The corresponding numbers are given in Table 2 for the SB1-type solutions, in Table 3 for the SB1C-type solutions and finally, in Table 4 for the TrendSB1-type solutions. We give the effect of each criterion on the tallies. The internal filtering has been actually performed in the same sequence as given in the table. These tallies are however only indicative because they can be dependent on the order of the adopted filtering. Only the number of entering objects and the number of remaining objects are not sensitive to the adopted order. As explained in Sect. 2, some objects can enter the SB1 processing coming from the SB2 processing (see also the details in Damerdji et al. 2024). Table 5 deals with the very few objects coming from the SB2 chain. Due to the filtering on flags nr.11 and nr.12, these objects do not pervade the final SB-subcatalogue. No class is associated to these objects.
We have now to estimate the number of solutions from our pipeline that were removed because they were combined and are now appearing in the EclipsingSpectro and AstroSpectroSB1 classes. Since we are only interested here on the disappearance of some solutions from our output due to the combination process, we can formally consider the tallies as if the combination were made after the internal filtering. The results of the combination mechanism for the 198885 SB1-type solutions are detailed in Table 5. The column Single corresponds to the solutions appearing in one class only whereas the column Duplicated refers to the objects appearing in two classes. All the 198885 objects are classified, except 2201 that will not be published due to a rejection during the process. Among these 2201 objects, 2107 correspond to a rejection at the level of the post-filtering which is detailed below. Some 94 other objects were rejected for other reasons. The majority of these 94 unpublished objects is corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB-subcatalogue. The Orbital and Others classes are given for the sake of completeness but they are not further considered either in the present work nor, for some, in DR3. Only the first three classes of the Table 5 are of importance for the present work. We will in addition restrict the present analysis of the SB1-type solutions to the SB1 class. To illustrate the approach, we will further consider the case of EclipsingSpectro. According to Gaia Collaboration et al. (2023a), some 155 objects are in this class. Over these 155 objects, 1 was originally attributed a solution of type SB2C, one SB1C, 2 SB2 and 151 SB1. Among these 151 SB1 solutions coming out of our main processing, 79 were internally filtered out. Therefore, some 72 solutions were dispatched to EclipsingSpectro and did not reach the SB1 class.
The solutions of type SB1C were cross-correlated along the same lines as the SB1 ones. No object was classified in another category by the NSS combiner. One object was rejected at the level of the post-filtering. Therefore, all the 202 (i.e. 203–1) SB1C solutions entered the eponymous class SB1C.
Finally, the TrendSB1-type solutions were compared with the astrometric channel results and 891 turned out to be redirected into the AstroSpectroSB1 class. They thus are not present in the TrendSB1 class, leading to an amount of 56808 objects. The details are explicited in Table 7. The majority of the 21 unpublished objects is corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB-subcatalogue. In this table, we also counted the TrendSB1 objects that presented or not a significant period at the level of the main processing (flag nr.13 raised means no period found).
| Solution types | From | Redirected from |
| SB1 chain | SB2 chain | |
| SB1 | 367328 | 107 |
| SB1C | 3533 | 2 |
| TrendSB1 | 168930 | - |
| StochasticSB1 | 322420 | - |
| Total | 862211 | 109 |
| Internal filtering step | Number of objects |
| Entering objects | 367328 |
| 71227 rejected | |
| Bad flags | 2840 rejected |
| 250 km s-1 | 89 rejected |
| Bad efficiency | 27749 rejected |
| Phase gap 0.3 | 25656 rejected |
| Non-significant period | 24251 rejected |
| Significance 5 | 15829 rejected |
| Too large | 802 rejected |
| Remaining objects | 198885 |
| Internal filtering step | Number of objects |
| Entering objects | 3533 |
| 1920 rejected | |
| Bad flags | 61 rejected |
| 250 km s-1 | 23 rejected |
| Bad efficiency | 1 rejected |
| Phase gap 0.3 | 588 rejected |
| Non-significant period | 735 rejected |
| Significance 5 | 2 rejected |
| Remaining objects | 203 |
| Internal filtering step | Number of objects |
| Entering objects | 168930 |
| 87223 rejected | |
| Bad flags | 1381 rejected |
| Time gap 0.3 | 12103 rejected |
| Passage to 0.99865 | 10503 rejected |
| Remaining objects | 57720 |
| Internal filtering step | Number of objects | |
| SB1FromSB2 | SB1CFromSB2 | |
| Entering objects | 107 | 2 |
| 64 rejected | 1 rejected | |
| Bad flags | 3 rejected | 0 rejected |
| 250 km s-1 | 0 rejected | 0 rejected |
| Bad efficiency | 3 rejected | 0 rejected |
| Non-significant period | 3 rejected | 0 rejected |
| Significance 5 | 4 rejected | 0 rejected |
| Too large | 0 rejected | – |
| Remaining objects | 30 | 1 |
| Class | Single | Duplicated | All |
| SB1 | 175725 | 5602 | 181327 |
| AstroSpectroSB1 | 15227 | 58 | 15285 |
| EclipsingSpectro | 71 | 1 | 72 |
| Orbital | 55 | 5291 | 5346 |
| Others | 17 | 371 | 388 |
| Total (all classes) | 191095 | 11323 | 202418 |
| Total (first three classes) | 191023 | 5661 | 196684 |
| Not published | - | - | 2201 |
| Total (distributed) | - | - | 198885 |
| Class | Count | Period | No |
| Period | |||
| FirstDegreeTrendSB1 | 24083 | 8391 | 15692 |
| SecondDegreeTrendSB1 | 32725 | 16989 | 15736 |
| Total TrendSB1 | 56808 | 25380 | 31428 |
| AstroSpectroSB1 | 891 | - | - |
| Not published | 21 | - | - |
| Total | 57720 |
After the combination process, a further step of filtering took place. This one was called post-filtering, and the motivation and practical details are described below in Sect. 10.1. The post-filtering was done on the basis of data coming from outside the RVS channel and is basically a process of blacklisting aiming at improving the purity of the catalogue. However, it should be made clear that it is partly constrained by the fact that the spectroscopic solutions are further used by the combiner, and thus we should be careful not to eliminate good solutions that could be useful downstream in the process. In the present section, we restrict ourselves to the effect of the applied post-filtering on the countings of the objects among the various classes. Table 6 details the effect on the number of objects considered in the important classes. The amount of objects before the post-filtering and after it are mentioned. The column after post-filtering gives the countings that are corresponding to the final DR3 catalogue.
The spectroscopic channel presented here had a strong importance in the building of the AstroSpectroSB1 class and a last information of interest must be given here. Namely, the origin of the various spectroscopic solutions that were combined with other types of data in order to build up the AstroSpectroSB1 class. The detailed origins are explicited in Table 9 for the four kinds of solutions. We dwell again on the fact that none SB1C-type solutions were combined. We also notice that some combination has been performed on the basis of the StochasticSB1 solutions.
A few pie charts illustrating these tallies are available in Appendix D.
| Countings | Discarded | Catalogue | |
| Class | before | at | after |
| post-filter. | post-filter. | post-filter. | |
| SB1 | 183434 | 2107 | 181327 |
| SB1C | 203 | 1 | 202 |
| EclipsingSpectro | 72 | 0 | 72 |
| AstroSpectroSB1 | 36082 | 2615 | 33467 |
| TrendSB1 | 56808 | 0 | 56808 |
| Original solution-types | Before | After |
| filtering | filtering | |
| SB1 | 16980 | 15285 |
| SB1C | 0 | 0 |
| TrendSB1 | 6104 | 891 |
| StochasticSB1 | 10383 | 10383 |
| Total | 33467 | 26559 |
8.2 Some illustrative results
In the present section, we want to illustrate various solutions pointing out the efficiency of the code. Here below, we adopt the convention to express the uncertainties as error bars of .
8.2.1 Good results
In the present section, we introduce a few examples of the solution fitted to selected RV time-series. The selection is somewhat arbitrary and is essentially motivated by the illustration of the efficiency of the chain. All these illustrating good solutions are part of the DR3 spectroscopic orbit catalogue (SB-subcatalogue). Figure 54 illustrates various orbital solutions that are part of the SB1 class. A small subset of these SB1 solutions is also provided in Fig. 6. The displayed phase diagrams cover a palette of values for the period associated to the SB1 solution. It is a first positive point that the pipeline is able to treat such a large domain of periodicity. At mid to large period, we also present a few examples of highly eccentric orbits, up to . It is a matter of fact that a good fit to a very eccentric orbit necessitates a large number of data points in the time-series, particularly near the periastron. An homogeneous distribution in phase is another necessity.
Since the vast majority of the analysed time-series covers a span of time around 800-1000 d, the solutions with short periods are extremely convincing since the large number of cycles folded in the phase diagram do not mitigate the goodness of fit and this can be associated to a high coherency of the periodicity. This conclusion holds when more than three cycles are covered (i.e. , or less). For larger periods, the coherency of the variations is not fully guaranteed by the quality of the goodness of fit. At larger values (), we are no more able to guarantee the coherency of the variation, since only one cycle is observed and this could correspond to a transitory event, not to a periodic variation. In any case, a tentative fit is still more indicative than no fit at all. However, the SB1-type solutions with such large up to (the adopted upper value allowed) should be considered as not secure as periodic solutions and the reported period is at the very best an indicative timescale. Indeed, for given periods between and 1.5 , the periods are not well determined and could even be potentially biased, due to the non-linear nature of the fit interacting with a possible bad phase coverage. This topic is further addressed in Sect. 9.4.1 and in Sect. 9.4.2. As illustrated below, the latter bias could induce biases on other parameters (see e.g. the case of the eccentricity in Fig. 64). The objects Gaia DR3 6422103351757549056, Gaia DR3 2509001872917567360, and Gaia DR3 5818903954144888320 are certainly concerned by this possible pitfall.
It is interesting to notice that several objects in the sample correspond to a quasi-circular solution: the eccentricity is very small, even perhaps not significant. A perfect illustration (see Fig. 6) is the case of Gaia DR3 1955225454244655872, whose eccentricity is with a 1- error of 0.021. This object most probably corresponds to a circular orbit, but the adopted solution is not an SB1C but an SB1. This SB1-type solution is associated to a solution efficiency of 0.16 and is thus on the verge to be processed with an SB1C model. This is due to the fact that we adopted the philosophy of not enforcing a circular solution even if the eccentricity turned out not to be significantly different from zero. We only adopt a circular solution when the structure matrix is such that it could not be properly inverted in the classical way and thus a full SB1-type solution is not possible. We are consequently constrained to decrease the number of free parameters. In fact, Gaia DR3 1955225454244655872 benefited from a large number of observations (good transits) and this large number of data points allowed us to fit an SB1 solution even for a very small apparent value of the eccentricity. Another example of truly circular orbit could be e.g. Gaia DR3 515682154702230272. Starting from these examples, we encourage the user of the catalogue to understand that the classes SB1 and SB1C differ by the fitted model and NOT by the true nature of the object. For example, someone interested in objects in circular orbits, i.e. to study only genuine circular orbit should certainly consider both SB1 and SB1C classes. If a user wants to transform an SB1-type solutions into an SB1C one, they can certainly use the period, the centre-of-mass velocity and the semi-amplitude as given by the SB1-type solution. However, it should not be overlooked that the argument of periastron is thus modified, since the for SB1C solutions corresponds by definition to the time of maximum velocity. Therefore, the must be accordingly corrected for this change of . Otherwise, the SB1-type solution is applicable to objects in true circular orbits.
Another interesting and enlightening case is the one of Gaia DR3 4983214300285468288. As can be seen from Fig. 6, the fit of the SB1-type solution is very good and no individual data point deviates significantly from the model. However, an expert eye (not the pipeline) will detect that around phase 0.65, the observed RVs are systematically grouped below the model and then above the modelled curve. Both groups of data points contain some 10 points and the offset could thus be significant (at least marginally) although the grouping is observed a posteriori. In any case, if this anomaly at phase 0.65 is real (and this will become clear in future releases), it is hardly explained in the framework of the Keplerian motion. The time corresponds to a phase where the primary is approaching the maximum velocity and thus the possible lines of the secondary are close to the maximum separation and should not influence the determination of the primary RVs (primary meaning here the star that dominates the combined spectrum). We are also (most probably) far from a possible eclipse, and this could not be due to a Rossiter-McLaughlin effect. The reported anomaly, if confirmed, will necessitate exotic effects to be explained in the context of a Keplerian model. It is worrying that several intrinsic variables (pulsating stars at least) could exhibit such a phenomenon at approximately this location in their RV curve. It is beyond the scope of the present paper to make a decision for this particular object but we would like to take this opportunity to highlight the following statement. It is not because a beautiful fit can be performed with a Keplerian model by the pipeline that the true nature of the object is the one of a binary. The membership to the SB1 or SB1C classes of solutions does not automatically imply the binarity of the object. Additional clues are necessary to converge to a sound classification. The binary nature of SB2 objects is much more convincing (although problems could also exist, see Damerdji et al. 2024). Concerning SB1 or SB1C objects, a conclusion could be drawn if, for example, the orbit is confirmed by the photometric or the astrometric channels (classes EclipsingSpectro or AstroSpectroSB1). The combined solutions are not part of the classes treated in the present paper. For the isolated spectroscopic orbital solutions reported here, we recommend at the very least, to build up a subcatalogue constituting some kind of gold sample, or to combine with additional data (e.g. positions in the HR diagram) in order to increase the veracity of the proposed classification.
The object Gaia DR3 1956836204422469888 is also found on the list of Hipparcos-Gaia proper motion anomaly of Kervella et al. (2019) and the list of Brandt (2021). All the objects included in Fig. 6 and in Fig. 54 were not previously known to have periodic RV variations.
In Fig. 57, we exhibit a few examples of orbital solutions of the class SB1C. A still smaller subset is provided in Fig. 7. A general finding is the fact that most of the concerned time-series are characterised by a rather limited number of data points. They also correspond to objects characterised by . This is related to the decision we took to fit an SB1C-type solution only when the SB1 failed: we must be constrained to limit the amount of free parameters. We recall here that the convention has been adopted that, for SB1C-type solutions, the represents the time of passage at maximum velocity. We note from Fig. 57 that we have circular solutions for a few large period objects. Here also, the majority of the objects were not previously known to present a periodic behaviour in their RV.
The object Gaia DR3 3384500857079232512 has been observed with Kepler and a period d (or twice that value) has been detected (Armstrong et al. 2015). This does not seem to be coherent with our results (d). The solution of this discrepancy will certainly be automatically found in future releases thanks to the important increase of the number of data points in the time-series under analysis. This conclusion is supported by the simulations reported in Sect. 9.1.
Another very interesting case is the one of Gaia DR3 1871154871054688384, also known as V2197 Cyg (NSVS 5761314). From ground-based observations, this star is a well-known SB2 star. From photometric data, the system seems to be an eclipsing system on a circular orbit with an inclination of about 80 °. It is in near contact and has a period of 0.4657489 d (=0.0000001 d). This is in good agreement with the Gaia spectroscopic period d (=0.0000034 d). The difference amounts to 2.3 . It is the first time ever that a spectroscopic period has been determined for this object. From the work of Nelson & Robb (2017), the primary star is of type F2-4 V and emits about 80 to 90 % of the visible light of the system. With 7 spectra taken at mid-resolution, these authors derive a RV curve leading to (km s-1 and (km s-1. These values compare very well with the (km s-1 and (km s-1 as derived from our pipeline. This illustrates the fact that the orbit of a relatively bright primary in an SB2 system observed as an SB1 one (here within the SB1C class) can be trusted. It is particularly interesting to look at the orbital solution in the phase diagram for this object (Fig. 7). From phase 0.5 to 1.0, the measured RVs are rather accurate: this corresponds to the primary being in front. On the opposite, from phase 0.0 to 0.5, the dispersion of the data points is larger; the unseen secondary is in front. The dispersion around the model is particularly important at phase 0.2-0.25. However, Nelson & Robb (2017) predict at these phases the presence of a Rossiter-McLaughlin effect and at phase 0.22, their model predicts a measured velocity of 40-50 km s-1, again in good agreement with the Gaia observations. However, the agreement is not so good for the point at a slightly later phase. In any case, this could be the first time that this effect is detected by Gaia. This will be further tested in future releases. We would like to recall here that the measurement of RVs by cross-correlation with templates is not the best method to address this effect.
Finally, some illustrative examples of single-line trend solutions are proposed in Fig. 58 with the small subset provided in Fig. 8. The various objects are ordered randomly except that the second-degree solutions occupy the last panels. None of these objects were previously known to exhibit RV variations. Gaia DR3 4979506918874792576 ( HD 4130) is also present in the lists of Kervella et al. (2019) and of Brandt (2021).
Gaia DR3 6070331129187853824 and Gaia DR3 5059000922370217728 have been observed a few times in the framework of the RAVE survey (Kunder et al. 2017); the first object of the two was also observed in the framework of the GALAH survey (Buder et al. 2018). No strong RV discrepancy with our orbital solutions is observed.
The velocities of the RV variable object Gaia DR3 4660348139680343424 (SV* HV5893) could be surprisingly large (see Fig. 8). Actually, this star is a known M supergiant in the Large Magellanic Cloud. It has been observed in photometry in the framework of the MACHO survey (Fraser et al. 2008) exhibiting a period of about 12.8 d. This is not directly detectable in Gaia RVs, but this is not either incompatible with the trend claimed here. Prevot et al. (1985) reported for this star a velocity of 310.4 km s-1 on HJD 2445662.77. The large difference in time between the two observed passages at this velocity is about 11975 d and suggests that the variation could be periodic with a scale of several thousands of days.
The solutions shown in Figs. 54, 57 and 58 are just illustrative and much more objects with solutions of similar quality are present in the SB-subcatalogue. This subcatalogue also contains solutions of slightly less good quality as well as solutions of intermediate quality. In the Appendix E, we provide a few examples of SB1-type solutions selected from the SB-subcatalogue at random but with around 0 (Fig. 59). This is to have fits of similar quality but for which the number of data points is lower (10 to 45) than in Fig. 6. Although the quality of these solutions is sometimes less evident, only very few could be easily questionable. Since we decided to push the software to its limit, solutions of more limited quality are equally present. For that reason, we encourage the general user to further filter the data. This is possible thanks to the quality factors that we deliver for each object (see Sect. 6.4). For a general statistical use, this should generate no problem, provided the user adopts the filtering to its particular application. We encourage the user to be very careful about the results on individual objects that could not benefit from the statistics. Some further information and advice can be found below in Sect. 11.2.
8.2.2 Problematic cases: multi-variability
In the framework of DR3, the SB1 pipeline is intended to search for orbital solutions. It is however designed to treat the cases which present only one type of variability (either SB-type with a single period or pure trends). If one object presents two types of variability, most of the time the algorithm could fail to find an adequate solution. We present here a typical case of badly handled data. Figure 9 shows the time-series of a star observed by Gaia RVS ( d). The data are unpublished and the solution is not part of the DR3 catalogue. The star is thus kept anonymous. The pipeline detected for this object a linear trend of about 1.15 km s-1 y-1. The solution corresponds to . This is easily understood because the RV curve exhibits extra-variations with a kind of oscillation with a period a little over 300 d. The variation is larger than the error bars and is thus significant. It necessarily induces a large value for coming from the fit of the trend, and the corresponding solution has thus been rejected from the DR3 catalogue. This object actually presents two superposed, different, types of variability. A trend that could be associated to a long period binary and, in addition, a roughly sinusoidal variation that could be due to an SB1 system in a triple system, or to an intrinsic RV variation or to something else. If, by chance, the pipeline would have first detected and reported an SB1 solution, the existing trend would have also implied a large associated to the SB1 fit and the solution would also have been rejected. This is clearly a weak point of the DR3 chain.
For the coming releases, the problem will be corrected and the pipeline will be able to process properly this kind of multiple variability by simultaneously fitting two types of solutions. This will open the door to the treatment of objects like triple or quadruple SB1 stars, as well as of intrinsic variables in a binary system.
8.2.3 Problematic cases: fake SB1 and spectrocentre
We present here the illustrative case of HD 223323. This star ( = 6.968) is a well-known binary object of type SB2 that has been well studied from the ground; it has a period of 1175.1 d (Griffin 2007). The two components are rather similar, and only a merged spectral type in the range F2-4 IV-V is reported (see references in Griffin 2007). An accurate orbital solution has been published by the latter author, and the elements of interest for the present discussion are given in Table 7. Figure 10 illustrates the fitted orbital model published by Griffin (2007). The ground observations at phase 0.98 (almost at larger separation, see figure 5 of Griffin 2007) suggest that the lines of the two stars are almost deblended.
| Parameter | Griffin | Gaia |
| (d) | 1175.11.3 | 977.9104.3 |
| (km s-1)∗*∗∗*∗The systemic velocity is expressed in the Griffin RV reference system (see also Sect. 9.4.1) | –9.560.06 | –9.940.10 |
| (km s-1) | 16.380.11 | 1.188 |
| (km s-1) | 16.400.16 | … |
| 1.001 | … | |
| e | 0.6040.003 | 0.5860.150 |
| (degree) | 77.80.60 | 56.011.0 |
| (solar mass) | 0.2720.006 | 9.0710-5 |
| (solar mass) | 0.2730.008 | … |
| (km s-1) | 0.53 | 1.61 |
The star has been observed with Gaia and is present in the DR3 catalogue under the name Gaia DR3 2852594583674129152. A total of 24 measurements () were collected by the RVS over a span of time of d. Two of the transits exhibit a double-line spectrum, and RVs have been derived as explained in Damerdji et al. (2024). For the remaining 22 transits, the star has been measured as a single-line object. Owing to the small percentage of time when the star appeared as double-lined, the pipeline decided to discard the two corresponding transits (see the treatment described by Damerdji et al. 2024). Therefore, the object entered the SB1 chain with data points. An SB1 fitted model was presented in the catalogue with a period 977.9 104.3 d. This value is in principle only indicative, since it is larger than the total span of time by a factor 1.11. In any case, the Gaia period is in good agreement with the precise Griffin period since it is only at 1.89 . The value of the period will automatically improve in the coming releases due to the increase of the . The other parameters of interest derived by Gaia are available in Table 7. The measured Gaia RVs are plotted in Fig. 10. The red dots represent the RVs derived by the single-line channel (they are put in the Griffin reference system by adding 1 km s-1, see Sect. 9.4.1 below). The other data points represent the primary (in blue) and the secondary (in cyan) RVs as derived in CU6 (Damerdji et al. 2024). They were also corrected to the Griffin reference zero point. It is remarkable that the Gaia solution yields a very small value for the semi-amplitude. The interpretation is quite straightforward. Except for the phases near the extrema in RVs, the resolution of the RVS and the predominance of the broad Calcium lines of the triplet makes that the lines never deblend at the other phases. After the rejection of the two double-line transits, the star becomes a fake SB1 which is actually a never-deblending SB2. There is a tendency for the position of the blended lines to moderately follow the behaviour of the primary variation. According to (Griffin 2007), the lines of the secondary are slightly broader and the lines of the primary slightly deeper. This thus implies that the primary lines are moderately dominating the blend. The secondary line depths are roughly 90 % of those of the primary one (we will call this ratio B = 0.9). The effect of the width of the lines of the secondary is attenuated since the lines are rather strongly blended. Tatum (1968) presented a computation of the position of the minimum of the blend of two Gaussian profiles (which is unique in the case of a strong blend) as a function of the widths and of the depths of the two profiles. Following their work, using the same notation as in their Eq. 5, and deciding to neglect the difference in broadness of the lines, we derived a position of the minimum which is given by
According to the latter author, adopts a value of 0 if the minimum is at the velocity of the primary and 1 at the secondary position. In the present case, we derive = 0.45 which means that the minimum is indeed attracted by the primary. As an approximation, we consider that the position of the minimum and the one of the blend are similar. In the same approach, we consider that the position of the blend can be approximated by a weighted sum of the velocities of the primary and of the secondary. It follows
consistently giving more weight to the primary velocity. Starting from the modelled RVs by Griffin, we computed the evolution of the position of the blend as a function of phase. This is illustrated in Fig. 10 by the green line. Despite the fact that this line is in no circumstances a fit, the agreement with the red dots is very good. This demonstrates the interpretation of the Gaia results as a fit of the blend. A more rigorous approach would be much heavier and is not in the scope of the present analysis which is intended to remain purely illustrative. By analogy with the notion of photocentre, we could introduce here the concept of spectrocentre which would correspond to the measurement of the RV of the blend (in the spectrum) constituted by the presence of two sets of undeblended lines. This could thus be termed a spectrocentre effect888It should however be pointed out that this effect is mainly dependent on the characteristics of both sets of lines (with a different contribution from line to line) and thus on the Astrophysical Parameters characterising the two stars..
It is interesting to notice that, in the case of HD 223323, the most affected parameter is essentially the semi-amplitude. On the opposite, it is clear from Table 7 that the other parameters are not strongly perturbed. The eccentricity in particular is well determined, as well as the argument of periastron. This is a consequence of the slight domination of the primary on the secondary effect.
The case of the never-deblending SB2 stars is crucial. If the two object spectra are similar (except for the Doppler shift), the semi-amplitude associated to the blend will near zero. Therefore, the corresponding object will be characterised by a very-small and also be associated to an abnormally small mass function. An extreme case with two identical stars will lead to a fake RV constant star. On the opposite, if the similarity of the two systems of lines for the two objects does not hold, the resulting semi-amplitude could of course be much larger and it is to be expected that such a case will be harder to detect. At the other extreme, if the secondary spectrum is negligible, the solution will tend towards a true SB1. The detection of the fake SB1 never-deblending SB2 is important, and nothing global has been implemented for DR3. We identify two avenues to solve the problem for future releases. The first one is to look at the evolution with phase of the broadness of the lines (even epoch by epoch). This could be done directly on the spectra. For DR3, an a posteriori look at the epoch vbroad values (Frémat et al. 2023) failed to produce results. An alternative would be to systematically reconstruct the mean spectrum at rest (according to the SB1 solution) and to detect line-profile variability through inspecting the dispersion of the reconstructed mean spectrum in the neighbourhood of the main spectroscopic lines. Due to the general architecture of the Gaia pipeline, these possibilities were impossible to implement in the DR3 pipeline. Similarly, we did not investigate in DR3 the possibility to use the information from astrometry to alleviate the problem.
8.2.4 Problematic cases: possible spuriously variable RVs
Some stars are associated with an orbital SB1-type solution that looks very convincing, and is identified as good according to the various quality tests. These stars thus enter the orbital solution catalogue. However, ground-based observations do not confirm these orbital variations. An illustrative case is the one of Gaia DR3 702159502868472832. Its RV time-series suggests a significant period of 1.0844000.000058 d with a probability of 0.997 and a resulting folded RV curve that is given in Fig. 11. With a significance of 15.41 and an , nothing allows us to doubt of the nature of this object. However, besides the Gaia RV curve and the fit, we plot also in the figure the few folded RVs that can be found in the literature from the APOGEE survey DR17 (Abdurro’uf et al. 2022). The two data sets are clearly not compatible. This suggests that some of the spectroscopic data from Gaia could not only present spurious RVs that are variable, but also that a period can be found from these measurements. This also suggests that the effect could be instrumental. The origin is currently not securely identified. A small number of objects enter this category, but the related sparsity could be due to the small percentage of objects having ground-based observations. This is supported by the validation performed in Sect. 9.4 below. Therefore, it is primordial to discover the origin in order to improve the robustness and purity of the future Data Release catalogues. We listed a few such doubtful cases that were detected here: they are Gaia DR3 66572718940795136, Gaia DR3 286652890498864768, Gaia DR3 487550398091020160, Gaia DR3 164423034699251968, Gaia DR3 577600125523574400, and the illustrative case Gaia DR3 702159502868472832. All of them have been compared to APOGEE DR17 data.
Another similar case has been detected by Bashi et al. (2022) but this time issued from comparison with the LAMOST RV survey (Cui et al. 2012, 999https://dr6.lamost.org). The object is Gaia DR3 3376949338201658112 and a significant periodicity has here also been identified ( d) and the corresponding RV curve is shown in Fig. 12. Again the RV variations are convincing as well as the fit. The object is one of the very few ones that have been observed at several epochs in the framework of the LAMOST survey. All the corresponding RVs are contained between the two green lines, further underlining the incompatibility of the ground-based data with the Gaia ones.
The fact that two ground-based surveys disagree with the results from Gaia underlines the fact that some instrumental effects are most probably present in the Gaia data. It is particularly intriguing that the RV time-series exhibit what could appear as a convincing periodic variability characterised by timescale not known to be problematic. Clearly, the problem is linked to the production of the RV time-series rather than to the behaviour of the algorithm of orbital parameter determination, except perhaps for the period.
9 Validation of the results
The analysis of the huge data set generated by Gaia and the establishment of the related catalogue necessitates some consolidation mechanism. This mechanism can be made of several approaches: they are described in the following subsections. The aim is evidently to test and fine-tune the pipeline itself in order to address its efficiency but also to discern the unavoidable biases. The understanding of the pipeline behaviour is directly related to the quality of the SB-subcatalogue and looking at its characteristics from all angles is a necessity. This validation process can be made on the basis of the internal consistency of the solutions but also on the basis of a comparison between the parameters delivered from the Gaia data with those obtained from the literature. This last approach is limited to the dramatically smaller amount of objects that have been the subject of a previous study. The validation paves the way for the determination of the global selection function.
9.1 Simulations
As a very first step, it is necessary to study the ability of the pipeline to recover the period. The study described here is based on the generation of synthetic RV curves of eccentric systems (with a random ), and of circular systems. Both sets of RV curves are generated with a random phase. We considered a number of data points of 10, 17, 30 and 80. The times of observations are mimicking the Gaia time series and are drawn in the same manner as in Sect. 4.4. The adopted signal-to-noise ratio is considered for values 5, 10 and 50. Fig. 13 reveals the recovery rate (in percent) for various true periods. The period is assumed to be recovered if its value does not differ from the input period by more than .
For equal to or larger than 10, the recovery rate is mostly better than 80 percent except when the number of data points is too low, see e.g. . This suggests that a too low number is a problem for the algorithm. It is also clear that an of 5 is equally detrimental. In Table 11, we give the corresponding value for the equivalent Significance given by
as derived from Eq.12 of Lucy & Sweeney (1971).
| 5 | 10 | 50 | |
| 5.6 | 11.2 | 55.9 | |
| 7.3 | 14.6 | 72.9 | |
| 9.7 | 19.4 | 96.8 | |
| 15.8 | 31.6 | 159.1 |
As an alternative, a cut-off at a Significance of about 10 is certainly a good attitude. It is interesting to note that the recovery rate is slightly better for circular orbits (see right panel of Fig. 13).
9.2 Internal verification of the results
9.2.1 SB1 and SB1C classes
The analysis presented here below is essentially based on the output of the spectroscopic pipeline after the internal filtering. The SB1 data set contains 183434 objects. The basic reason for this choice is that the principal motivation is to first scrutinise the behaviour of the pipeline. The post-filtering is based on information from other channels and could thus be more complex to interpret, although it is necessary in order to constitute the DR3 SB-subcatalogue. In any case, for DR3, the differences between the analysed data set and the data set after post-filtering that contains 181327 objects, are almost negligible from a statistical point of view. The resulting data coming out of the pipeline were systematically inspected in order to verify the conformity to the expectation and the expected statistical distribution of the fitted parameters. Basically we are mainly treating the SB1 class excluding the SB1 solutions that were combined downstream with the solutions from other channels. The SB1C class contains too few objects to allow us to perform a fully sound statistical analysis. However, these solutions correspond to a very simple model that should not bear any particular problem.
The distribution of the number of good transits (i.e. data points) for the SB1-type solutions processed by the pipeline is illustrated in Fig. 14. The objects having less than 10 transits were not further considered as explained above. This cut-off below 10 transits should not discard too many objects. Typically, we have slightly over 20 data points in our various time-series. The largest value is 228 (not shown here) but these objects are exceptional and are corresponding to the EPSL observations. As a first test, we examine the distribution of the magnitude of all solutions in the SB1 class. Figure 15 exhibits the histogram of for all the 183434 objects. We immediately notice that no object is present between and 5.5. This is an artefact of the selection based on the rv_renormalized_gof at the level of the input data. This quantity is defined when and thus below this threshold, the value for these objects contains NaN even if these bright objects should be kept. This is a weakness of DR3 and it will be corrected in future releases. On the other hand, it is also clear that signs of incompleteness are starting to appear above magnitude 11.5. A minimum of objects are lost between magnitude 11.75 and 12.0. This is most probably an effect of the increased importance that noise could have for the fainter objects, thus leading to badly determined solutions. However, an incompleteness also appears in the global RV output of the CU6 pipeline (see fig. 4 of Katz et al. 2023) for objects fainter than (corresponding for a solar-like star to ). Therefore, the difficulties to measure RVs on noisy spectra, the selection of variable objects in noisy time-series and the subsequent fit of orbital solutions conspire to induce this incompleteness.
The observed distribution of the quality-of-fitting statistic deduced from the chi-square, the goodness of fit , is given for the filtered data set of solutions (see Fig. 16). The recommended and applied cut-off at is clearly visible. Otherwise, the original decay of the histogram is going up to values of several tens. Large values of the statistic are to be associated with bad, unacceptable solutions. This is certainly so for larger than 5. The cases between 3 and 5 could be further discussed. Conservatively, we only kept values below 3. This induces the rejection of some acceptable solutions, and consequently of the related objects. Along with the observed distribution, we also plot a Gaussian around the mean zero and with a dispersion (in magenta). This curve represents the expected behaviour of under the null-hypothesis of a perfect Keplerian model with additive Gaussian noise (corresponding to the uncertainties associated to the individual RVs). Another Gaussian curve with a (in blue) is also shown. This latter better follows the histogram on its left tail. The side of the negative values of corresponds to solutions for which the realisation of the noise conspires to produce better fitting. On the other hand, the right tail is associated to the opposite effect where the solutions are noisier although still acceptable. It is interesting to notice that, in this region, the histogram reveals a remarkably too large number of objects. This is true even comparing to the blue Gaussian curve; an excess is thus present. They most probably correspond to solutions for objects following a Keplerian orbit (thus generating a rather good fit) but affected by additional variations (pulsations, activity, spots, …), not due to the observational noise, but at a sufficiently low level such that the dominating effect is still the orbital variation. Among these objects, multiple systems and intrinsic variables could be the main contributors to this effect. A more sophisticated study, beyond the scope of the present paper, is necessary. However, one should not be misled by the normalisation of the comparison Gaussian curves. Actually, if we normalised with respect to the area, there would appear a deficit (roughly 25%) of objects with solutions characterised by an broadly located around zero (essentially between -1. and +1.). These solutions are displaced towards the right tail. The effect is present both on the individual (with a selected value of , the number of degrees of freedom) and on the global statistics. The transformation of the values into the global statistic is not the cause of the deviation.
Coming back to the anomaly of the left tail, it could be due to a general over-estimation of the individual uncertainties on the measured RVs. This would displace the whole distribution by 15% to the left but the above-mentioned effect on the right tail would prevent any confirmation based on this right tail. This explanation would induce an effect independent of the number of data points.
Around , there is a deficit of solutions with large significance (roughly above 40). Consequently, any attempt to correct the shift to the left of the histogram by filtering out low-significance solutions, implies the appearance of a shift of the maximum of the peak in the histogram towards positive values of . This is not the solution either. It is interesting to notice that, if we limit the population of the solutions to those with and thus to a number of degrees of freedom , the tail is only shifted by a factor 1.07 (7% of excess). Therefore, the observed shift seems to be more a matter of the amount of data points. The distribution of solution efficiencies is peaking at a value of 0.44. Indeed, if alternatively we select solutions with between 4 and 9, we obtain the distribution exhibited in Fig. 17, further strengthening our conclusion. For low , the efficiencies are peaking at values lower than 0.3. The behaviour of the left tail as a function of could be due to the apparition for small amount of data points of overfitting or of effects due to the intrinsic non-linearities of the model that are inducing correlations among the parameters due to particularly bad distributions in phase of the observations and/or effects due to the fact that the individual uncertainties on RVs are estimators with a random distribution far from Gaussian.
We further inspected (not shown here) the distributions of the solutions as a function of two parameters, in particular versus rv_n_good_obs_primary or versus . Nothing particular is deduced from these plots. The reason is that the 2D distributions of the solutions are well interpreted as the product of the marginal 1D distributions.
The next step we adopted was to create the histogram of the argument of periastron distribution (see Fig. 18). The distribution is rather uniform, with a very slight tendency to have more objects at 90° (/2) and at 270° (3/2). This results from a very small bias because saw-tooth RV curves are easier to detect than shark-tooth RV curves, particularly at high eccentricity. The effect is practically negligible.
Another interesting plot is the distribution of versus the period. It is given in Fig. 19. Since is a cyclic variable of modulo (see Sect. 3), an infinity of possible values are available. The data points being located around , the value presenting the best precision and a minimum correlation is necessarily situated between and . It is clear that the graph is essentially as expected with a rather homogeneous density of data points slowly decreasing towards the largest periods. For periods around 900 to 1100 days, there is a slight tendency to have a larger density of solutions near the values . A part of the solutions located there correspond to low significance and rather large eccentricities. This is dominantly an artefact of the decision we took to enforce an SB1 solution in this region of periods against the possibility to consider trends. Indeed, let us consider a time-series of a noisy linear trend. The enforced SB1 solution will deliver a period roughly about (only one ‘cycle’ but this depends on the exact sampling). The trend folded in phase will very often mimic an eccentric curve with a jump between the last data point of the time-series and the first one. Therefore, this jump, after folding, will be associated to the passage at periastron and, consequently, the returned will be attributed a value around . This conclusion is further supported by the 2D distributions of solutions (i.e. versus and ) shown in Fig. 20. One clearly sees that the excess of solutions is effectively due to the range of periods enforced to an SB1 solution (prominent zone in yellow in the bottom right panel, not present on the upper left and right ones). It is also clear that the effect is absent for quasi-circular orbits (which are fully correct) and is associated to larger eccentricities. In addition, one clearly sees that the solutions at the basis of the excess are confined to values of around and (very same zone in yellow). This indicates that the trends enforced into an SB1 solution generate a RV curve that is predominantly of the saw-tooth type, further confirming our conclusions. The excess partly contributes to the excess at specific noticed in Fig. 18. This problem will be corrected for DR4; it will also occur at another period since will be larger. Some other effects might also be present. Note that not all the solutions in this region are problematic.
In the next figure, Fig. 21, we plot the histogram of the distribution of the RVs of the centres of mass. This histogram is very similar to the one built for the assumed constant stars for EDR3 (see figure 4 of Seabroke et al. 2021) and for DR3 (see figure 11 of Katz et al. 2023), which is a strong indication of the quality of these centre of mass RVs. The latter curve from DR3 is overplotted in Fig. 21 to facilitate the detailed comparison. The histogram for the constant stars is slightly more dispersed, but the agreement is actually good. The remaining discrepancy is a matter of sampling different stellar populations. Indeed, if we compare with a population restricted to stars cooler than 8125 K and with brighter than 12, the agreement becomes almost perfect.
The distribution of the derived semi-amplitudes is given in Fig. 22. The major part of the distribution is decreasing for increasing . The decay law seems to be essentially linear on this logarithmic scale. Taking into account small number statistics (as seen above 150 km s-1), no break is visible in the distribution function and we will retain that the distribution is rather smooth, which indicates a very good quality of the results. No anomaly is detected.
In the case of noiseless data, the difference between the two extrema of the individual RV curves is expected to be a good estimator of twice the semi-amplitude. We show in Fig. 23 a plot of the rv_amplitude_robust (Sartoretti et al. 2022) versus the ’s. We conclude that the expected relation is certainly present. The distribution below the reference line (in blue) is essentially explained by the noise on the two values. The dispersion over the reference line is markedly larger. This is due to the fact that the rv_amplitude_robust has a larger dispersion, being the difference of the two extreme values (even after some cleaning). This means that it is more sensitive to persisting outliers or even slight extreme values (or noise) whereas the semi-amplitude is the result of a fit on all the data points. It appears that the discrepant data are more numerous at large . This very small population of solutions (in red) at large is associated to solutions with a large ; the corresponding objects are characterised by large values of vbroad (Frémat et al. 2023). Based on a similar picture (not shown here) for the objects in the SB1C class, the conclusions are quite identical.
Figure 24 presents the distribution of the uncertainty on the semi-amplitude as a function of the uncertainty on the RV of the centre of mass . The ratio of the two is expected to be the square root of 2 (Lucy & Sweeney 1971) which is what is observed despite a rather large dispersion. No particular anomaly is present on the uncertainties of the considered parameters.
The critical point when fitting the orbital SB1 models is the determination of the period (see above). Fig. 25 displays the histogram of the period (in log). Several features are visible. The log-distribution is rather flat, with a marked excess at very large periods. This excess can be due to the fact (see above) that the periods of the same order of magnitude as the span of time of the mission are rather indicative and certainly not a perfect estimator of the true period. This uncertainty implies the concentration of several solutions near these long periods. This excess is very sensitive to the threshold separating SB1 and TrendSB1 solutions. Another explanation is the fact that the orbital solutions are enforced in this region. On the other hand, the deficit of solutions around the period of 182.6 days is due to the yearly Earth orbit that the Gaia satellite follows by staying at L2. The satellite avoiding the direction of the Sun, this produces regular gaps with a timescale of half a year. These gaps presented by the sampling are associated to large gaps in phase leading to the removal of the solutions because they correspond to badly determined periods and orbits. Fig. 27 is showing a diagram of the value of the largest gap in the solution as a function of the period. Exceptionally, here we use the data set of 367328 objects constituting the set of SB1 solutions before the application of the internal filtering. It is clear from this figure that some of the solutions at periods near 182.6 d generate phase diagrams that are associated to the presence of vast gaps. This situation seems also to be the case, but to a lesser extent, for the period around one year. The phenomenon is perhaps also present for periods of 91.3 d. A detailed look at this Fig. 27 suggests that a conservative choice for the threshold on the gap of 0.3 removed all the potentially spurious or at least unsafe solutions. The decrease of density of solutions for lower periods, as seen in Fig. 25, is an artefact of the application of an ad hoc selection criterion (see the paragraph on the internal filtering of the period in Sect. 7.1) that prevents the proliferation of the small periodicities. This effect should be further investigated for the next Data Releases.
In Fig. 26, we present a diagram showing the efficiency of the solutions as a function of the adopted period. We highlight here the long-period objects. For periods between 200 and 1000 d, the efficiency is essentially distributed between 0.2 and 0.8, independently of the period with a maximum of occurrence around 0.44 (as already reported above). The values of solution efficiency above 0.8 are basically absent, indicating that all the solutions present correlations among the final parameters. These values are part of the catalogue. There are very few solutions with efficiencies between 0.1 and 0.2, and all these solutions could be considered as doubtful, although we preferred to maintain them in the catalogue and to let the user the possibility to reject them or not. Above periods of 1000 d, the typical efficiencies are lessening and a population below 0.2 is appearing. This is a direct consequence of the fact that the associated period is larger than . It is evident that these solutions should not be blindly trusted, but in any case the reported periods can be considered as indicative. Again, the low value of the efficiency unveils existing correlations among the final parameters due to bad phase coverage. As mentioned above, for variabilities at still larger timescales, the trend solution is taking over. However, addressing solutions slightly above would necessitate a too large degree of the fitted trend polynomial, introducing a lack of robustness of the fit. The orbital solution is thus more informative despite the possible biases in the derived parameters. For future releases, the range of concerned periods will change accordingly with the increasing value of .
Figure 28 presents the histogram of the spectroscopic significance . The significance is distributed as a decreasing power law vanishing at 800. The maximum accepted value for being 250 km s-1, the maximum significance value corresponds to a typical uncertainty of about 0.3 km s-1 which is coherent with expectation. The distribution shows no inflection locus that could help to set a selection threshold. In this context, we decided to introduce a threshold at 5 but not more restrictive ones. Below a value of 5, the corresponding solutions are clearly not trustworthy. We leave the possibility to the catalogue users to define a threshold that could be appropriate for their own work.
Fig. 29 exhibits the distribution of the standard deviations on the eccentricity. The typical uncertainty on the eccentricity is of 0.03-0.10 . It is necessary to draw the attention to the fact that with such an uncertainty a derived eccentricity of 0.3-0.4 is still compatible with a circular orbit (this is within three keeping in mind that the distribution of eccentricities is not Gaussian). RVs are well-known to be a poor source of determination of eccentricities in contrast to the astrometric orbital solutions. Errors larger than 0.35-0.4 (there are a few) could be associated to insecure adjustments and we encourage the catalogue user to be careful with these questionable solutions.
As an example, Fig. 30 shows a plot of the eccentricities as a function of the corresponding period. The plot of all the objects (in red) includes short period solutions with some eccentricity up to 0.4. As mentioned just above, these values of up to 0.3-0.4 are not particularly anomalous. However, the general aspect of the diagram is not the expected one in terms of the theory of circularisation of the orbits (close binaries are not expected to remain eccentric for long). If we restrict the data set to values with a significance larger than 30, we obtain the distribution in blue where the expected effects of the circularisation start to be well visible. A restriction to the significance values above 80 generates the distribution in orange. Another study of this topic can be found in Gaia Collaboration et al. (2023a). Evidently, the impact of the selection on the basis of the significance is a delicate topic and we advise the greatest caution for the selection of the threshold to be applied. This remark can also be extended to other quality factors than just the significance.
Finally, Fig. 31 represents the plot of the derived SB1 mass function as a function of the adopted period. One clearly sees that the majority of the data points are between 0.001 and 0.5 M☉. Objects outside this range are very few, in particular on the side of the high mass function. From the statistical point of view, the diagram is rather close to the expected one. Data points at an unexpected location could be due to problems in the derivation of the period or of the eccentricity. In particular, one should be very careful while searching for abnormal mass function values or massive black holes. In addition, the overdensities tracing vertical lines at days and at one year (much broader) are due to objects with a varying apparent position inducing systematic effects on the RVs associated to the variation of the scan-angle. These are objects presenting small amplitudes at these particular periodicities (bad astrometry inducing a wrong wavelength calibration). These objects could easily be filtered out but they were not before the application of the combiner with the astrometric channel in order not to miss interesting objects. The effect in astrometry is described in Lindegren (2019). A more detailed analysis of the scan-angle dependent signals is given in Holl et al. (2023a).
9.2.2 FirstDegreeTrendSB1 and SecondDegreeTrendSB1 classes
Concerning trend solutions, we show in Fig. 32 the distribution of the goodness of fit for all the TrendSB1 solutions. It appears that the expected distribution that should be centred on zero, is actually rather centred on 1. This points out to an anomaly of this distribution, tending to suggest that the fits are not perfect although still good. The reason of this shift is for the moment unknown; to understand it is very important because we should study if the cut-off at 3 is still justified in the context of the observed shift of the distribution. The characteristics of the distribution is similar for the two subtypes of solutions.
The trend is certainly not a purely physical model, since it is basically induced by the shortness of the time span of the observations compared to the true cycle of the star considered. Statistical decisions are adopted to select a model (orbital, linear trend, high-order trend) but it seems unlikely that the adopted thresholds could induce the shift of the distribution in such a global manner. Another explanation for the shift could be the presence of some intrinsic variability additional to the trend signal, as already evoked for orbital solutions. Again, it is very unlikely that this could affect such a large percentage of the objects. Thus, the last possibility and most likely hypothesis is that the individual uncertainties associated to RVs are underestimated. The trend models are certainly a good way to investigate this possibility. Indeed, the linear trend is a very simple model which is linear in the parameters and the fit should be robust and unambiguous at least compared to the non-linear orbital model whose behaviour could be more complex (see Sect. 9.2.1). The overfitting is here not a concern, and the number of degrees of freedom of the of the fit is precisely established. We thus further investigated the global distribution (over all the objects) of the observed minus calculated residuals normalised by the associated individual uncertainties. To be strictly conservative, we only considered objects with an astrometric ruwe less than 1.4 for which no additional contribution to the uncertainties was applied; only the values delivered by CU6 are involved. The distribution of the normalised residuals follows very closely a Gaussian probability density function, centred on zero, but with a standard deviation larger than : actually we deduced .08. Therefore, an increase by 8 percent of the uncertainties could solve the reported anomaly of the distribution except for a small effect that could be explained by some additional variability. This statement needs further confirmation.
Fig. 33 illustrates the distribution of the first derivative normalised by its uncertainty. The distribution is very similar for the two classes and presents an almost perfect symmetry, as expected. The deficit of solutions around the zero derivative is due to the fact that these solutions are not included because they correspond to constant objects. A lot of the solutions are well above . A few well-defined derivatives are further away up to .
Finally, we try to validate the values of the derived . They are plotted in Fig. 34 with in ordinates the mean velocity computed by MTA. We observe that the two parameters are very well correlated, providing further support to the values we are delivering.
9.3 Validation against CU7 photometric periods
Some intrinsic variables can exhibit RV variations that could be mistaken as the effect of an orbital motion. As explained in Sect. 10.1.1, a sample of securely identified Cepheids (333, Ripepi et al. 2023) and RR Lyrae (147, Clementini et al. 2023) have also an SB1 solution. They should be eventually excluded from our SB-subcatalogue but they also offer the unique opportunity to have an independent test of the spectroscopic periods by comparing them to the photometric ones. Indeed, we compared our spectroscopic periods to those determined from DR3 photometric data for the 147 RR Lyrae and the 333 Cepheids. The difference is expressed as [P(spectroscopy)–P(photometry)]/, where is the square root of the sum in quadrature of both uncertainties. We find a recovery success rate to within 3 of 95% for the RR Lyrae’s. It is not only lower for the Cepheids (85%), but the outliers are also much more severely discrepant (Fig. 35). We find that the solution efficiency is better for the RR Lyrae’s than for the Cepheids.
9.4 Validation against other data sets
We present below a comparison between the DR3 orbital parameters and those from external catalogues, either relying on spectroscopic (Griffin, SB9, and APOGEE) or photometric (WISE and ASAS-SN) data. The most distinctive feature between the various validation samples is the widely different range of orbital periods that span about five orders of magnitude (see Fig. 36). The WISE and ASAS-SN catalogues are made up of candidate short-period binaries that are useful for assessing the reliability of our orbital solutions in this regime. We consider exactly the same DR3 data set as discussed in Sect. 9.2. In the following, except for ASAS-SN, literature results without an uncertainty given are ignored. The difference for a given orbital parameter is expressed as [value(NSS)–value(reference)]/, where is the square root of the sum in quadrature of both uncertainties.
9.4.1 Comparison against Griffin’s results
We first considered the orbits determined in a series of papers (not only in The Observatory) by Roger Griffin and collaborators. The orbital elements were taken from the 2 March 2021 version of the SB9 catalogue101010https://sb9.astro.ulb.ac.be/ (original paper: Pourbaix et al. 2004), which is made up of 4021 systems. The association to the relevant DR3 ID was simply tracked from the DR2 alias available in SIMBAD111111https://simbad.u-strasbg.fr/simbad/ at that time. Since for validation we needed to be confident in our reference catalogue, we remained conservative at this level, avoiding adverse and insecure cross-matching. This procedure can now certainly be much improved. To avoid any ambiguities in this cross-match, or a possible bias in centre-of-mass velocity, , visual/multiple systems were excluded. They were identified as having a component field in SB9 not empty or to have a SB9 ID entry in the 23 June 2020 version of the Multiple Star Catalog (MSC)121212http://www.ctio.noirlab.edu/~atokovin/stars/. This validation catalogue contains the orbits of 414 SB1’s and 101 SB2’s.
The breakdown of DR3 spectroscopic solution types is as follows for Griffin’s SB1 catalogue:
SB1: 83 systems
SB1C: 0 system
TrendSB1: 10 systems
SB2: 1 system
SB2C: 0 system
Without any DR3 spectroscopic deterministic solutions: 320 systems.
The period of the ten systems with a trend solution is much longer than the time span of the RVS observations, which is as expected. Five and one systems among the SB2 validation catalogue have received a SB1 and TrendSB1 solution, respectively.
Figures 60, 61, 62, and 63 show the comparison between the DR3 and Griffin’s SB1 orbital parameters, as a function of various DR3 parameters. The deviations in expressed in units are shown in Fig. 37, as a function of the reference period. Systems for which the latter is difficult to find because of the limited length of the RVS observations (i.e. = time span observations ) are indicated with a different colour. Systems for which this period is not recovered to within 3 are also flagged.
Some deviating points can be noticed in Figs. 60, 61, 62, and 63, especially when the period is not recovered. The incidence of outliers is quantified below. In some cases, the agreement is satisfactory in absolute terms and the deviation is only due to (very) small uncertainties. There is clear evidence for a systematic offset, although the discrepancy only amounts to km s-1 on average. It might be due to the fact that the RV values adopted by Griffin for his four reference stars back in the late sixties (Griffin 1969) appear to be larger by roughly the same amount than those in the recent literature, including Soubiran et al. (2018) for Lyr (Fig. 38).
Table 12 summarises the recovery rate of the SB1 parameters to within 3. Cases where the time span of the RVS observations is longer/shorter than the reference period are listed separately. As may be expected, the performance of the pipeline greatly improves when the observations cover more than one orbital cycle.
| 96.4 (55) | 12.0 (50) | 94.1 (51) | 96.1 (51) | |
| 88.2 (17) | 17.6 (17) | 58.8 (17) | 35.3 (17) | |
| All | 94.4 (72) | 13.4 (67) | 85.3 (68) | 80.9 (68) |
9.4.2 Comparison against SB9 catalogue
As a next step, the analysis above was repeated for the whole SB9. Although the sample is now considerably larger, it is much less homogeneous and contains solutions of generally lower quality. The most recent orbit was selected when a given system had multiple entries in the catalogue. As described above, visual/multiple systems were also excluded. Six SB2 orbits were replaced by those determined by Lester (2020). This validation catalogue contains the orbits of 1750 SB1’s and 647 SB2’s.
The breakdown of DR3 spectroscopic solution types is as follows for SB1’s in the SB9 catalogue:
SB1: 226 systems
SB1C: 0 system
TrendSB1: 45 systems
SB2: 3 systems
SB2C: 0 system
Without any DR3 spectroscopic deterministic solutions: 1476 systems.
All but one of the 45 systems with a trend solution have a very long period. Eighteen and two systems among the SB2 validation catalogue have received a SB1 and TrendSB1 solution, respectively.
Figures 64, 65, 66, and 67 show the comparison between the orbital parameters. The deviations in expressed in units are shown in Fig. 39, as a function of the reference period.
The results are consistent with those obtained for Griffin’s sample. Some features are expected. For instance, the difficulty to find the correct period for (very) wide, low-amplitude systems: the orbital periods are generally systematically underestimated (Fig. 67). Once again, the large discrepancies may be misleading in some cases. For instance, some short-period systems are outliers in because the uncertainties are very small (10-4 d). The same conclusion holds for SB2’s (see Damerdji et al. 2024). The slight systematic discrepancy and large dispersion for is likely due to the fact that a significant fraction of the orbits are actually taken from Griffin (see discussion about offset in Sect. 9.4.1). The deviations show some evidence for a bimodal distribution (Fig. 65), and it indeed appears that the smaller peak is largely dominated by results from Griffin. The RV bias of up to 75 m s-1 affecting the faint end (Sect. 2) is negligible in this respect. Furthermore, there are very few objects with 11 mag (Fig. 65).
Table 13 summarises the recovery rate of the SB1 parameters to within 3. Once again, not surprisingly, the incidence of outliers decreases when the time span of the observations exceeds one orbital cycle.
| 89.2 (167) | 53.9 (152) | 90.4 (156) | 92.7 (150) | |
| 81.1 (37) | 43.2 (37) | 64.9 (37) | 37.8 (37) | |
| All | 87.7 (204) | 51.9 (189) | 85.4 (193) | 81.8 (187) |
9.4.3 Comparison against APOGEE
Orbital parameters for 20,000 SB1’s have been released as part of APO Galactic Evolution Experiment (APOGEE)131313http://www.sdss3.org/surveys/apogee.php DR16 (Price-Whelan et al. 2020a, hereafter PW20). Our validation sample is their ’Gold Sample’ that contains 1032 systems with high-quality, unimodal posterior samplings.
The breakdown of DR3 spectroscopic solution types is as follows for the SB1’s in PW20:
SB1: 126 systems
SB1C: 0 system
TrendSB1: 4 systems
SB2: 3 systems
SB2C: 0 system
Without any DR3 spectroscopic deterministic solutions: 899 systems.
All the systems with a trend solution have a period much longer than .
Figures 68, 69, 70, and 71 show the comparison between the orbital parameters. The deviations in expressed in units are shown in Fig. 40, as a function of the reference period. For the period, two distinct situations are encountered: either there is a reasonable correspondence or, in about 25% of the cases, the values are dramatically discrepant. For those, systematic differences are not observed. The solutions in PW20 for the majority of these binaries rely on much fewer RVs (less than ten). It is likely that the paucity of measurements prevents a robust determination of the orbital period. We caution that these systems are off scale in Figs. 40 and 71.
Table 14 summarises the recovery rate of the SB1 parameters to within 3. The lower incidence of outliers when a full orbital cycle is not covered may be attributed to small number statistics.
| 87.6 (121) | 73.6 (121) | 85.1 (121) | 69.4 (121) | |
| 100.0 (5) | 100.0 (5) | 100.0 (5) | 80.0 (5) | |
| All | 88.1 (126) | 74.6 (126) | 85.7 (126) | 69.8 (126) |
9.4.4 Comparison against WISE
A total of 56,000 short-period ( 10 d) binary systems observed by the Wide-field Infrared Survey Explorer (WISE)141414https://www.jpl.nasa.gov/missions/wide-field-infrared-survey-explorer-wise have been discussed by Petrosky et al. (2021). Most of them are (near-)contact EBs. However, as will be shown below, a non-negligible amount are classical pulsators, especially RR Lyrae stars. The infrared photometric measurements are made through the W1 filter centred at about 3.4 m. A subset of 49,465 systems can be unambiguously associated to a DR3 source151515http://cdn.gea.esac.esa.int/Gaia/gedr3/cross_match/. This sample allows us to test the reliability of our orbital periods for very close binaries, although other orbital parameters are not available.
The breakdown of DR3 spectroscopic solution types is as follows for the SB1’s in Petrosky et al. (2021):
SB1: 704 systems
SB1C: 3 systems
TrendSB1: 12 systems
SB2: 69 systems
SB2C: 23 systems
Without any DR3 spectroscopic deterministic solutions: 48,654 systems.
We find a close correspondence in most cases with either the main WISE period or its second harmonic (Fig. 41). As discussed by Petrosky et al. (2021), the orbital period they found for the binaries is usually half the true value because of the nearly symmetrical nature of the WISE light curves. The WISE period is recovered to within 1% for 25.3% and 55.4% of the cases for the first and second harmonic, respectively. The overall recovery rate at this level is thus about 81%. It remains quite high for lower tolerances (e.g. 78% for 0.1%). The deviations lie in the vast majority of cases within five times the mutual uncertainties (Fig. 42).
Most of the WISE sources were assigned a classification in DR3 based on their light variations (Mowlavi et al. 2023; Ripepi et al. 2023; Clementini et al. 2023). As can be seen in Fig. 43, a (nearly) circular solution is found for the majority of the EBs, as may be expected for such close binaries. In contrast, the systems with high eccentricities are mostly classical pulsators (Cepheids or RR Lyrae) mistaken as binaries. As described in Sect. 10.1.1, those pulsating variables were rejected from the final DR3 catalogue during the post-filtering process.
9.4.5 Comparison against ASAS-SN
The physical parameters of 35,464 detached eclipsing binaries (DEBs) observed in the framework of the All-Sky Automated Survey for Supernovae (ASAS-SN; Kochanek et al. 2017)161616https://www.astronomy.ohio-state.edu/asassn/ have been determined by Rowan et al. (2022). A total of 35,307 systems have a Gaia DR3 ID. The data are based on the analysis of - and -band light curves. The orbital periods nicely bridge the gap between the WISE sample of close binaries and other more general catalogues, such as APOGEE or the SB9 (Fig. 36). Contrary to WISE, other orbital elements (e.g. eccentricity) are provided besides the period, although without any uncertainties. For , we assume a typical relative uncertainty of 10-6 (e.g. Holanda et al. 2018).
The breakdown of DR3 spectroscopic solution types is as follows for the binaries in Rowan et al. (2022):
SB1: 626 systems
SB1C: 3 systems
TrendSB1: 6 systems
SB2: 92 systems
SB2C: 35 systems
Without any DR3 spectroscopic deterministic solutions: 34,545 systems.
Most objects have 13 mag and are fainter than our magnitude cut-off.
The majority of the 626 objects are classified in DR3 as EBs based on their photometric behaviour (Mowlavi et al. 2023). Conversely, none are identified as classical pulsators (Clementini et al. 2023; Ripepi et al. 2023). For this sample with SB1-type solutions, the overall recovery rate of the photometric period to within 3 amounts to 83% (Fig. 44). This estimate includes one system where the ASAS-SN period is actually the harmonic (as discussed for WISE above) or for nine objects twice the true value because one of the two eclipses is too shallow to be detected (see figure 6 of Rowan et al. 2022 for an example). Visual inspection of the ASAS-SN light curves for the nine DEBs supports the latter explanation.
The relevance of filtering out short-period systems with a low solution significance was discussed in Sect. 9.2. We show in Fig. 45 the DR3 and ASAS-SN data in a - diagram for various cuts in solution significance (see also Fig. 30). The upper envelopes defined by Mazeh (2008) and Halbwachs et al. (2005, modified to at 1 d) are overplotted. The former is based on an empirical fit of the SB9 data, but the latter rests on theoretical arguments. While is recovered in 67% of the cases for / values up to 15, the proportion sharply increases for higher significances (success rate of 90% on average). The differences are generally small in absolute terms, which explains the close match in the period distributions shown in Fig. 45. However, an excess of seemingly eccentric systems with respect to Rowan et al. (2022) is readily visible. The eccentricity distributions become more and more similar when the significance cut-off increases until they become statistically indistinguishable above about 50. Rather than a poor determination of the period for close binaries, the issue thus seems to rather lie in solutions that are too eccentric when they are ill-constrained (see also Jayasinghe et al. 2023).
9.4.6 Summary
As a preamble to summarising the main conclusions of our analysis, it is worth emphasising that some of the validation data sets discussed above are of limited usefulness in the sense that they do not offer stringent tests of our results. In particular, the APOGEE orbits from PW20 are based on fewer epochs (about 14 on average concerning the 126 objects) than most NSS sources (see Fig. 14). As a consequence, their orbital solutions are not necessarily more robust and precise. We indeed find a poorer match for the periods with respect to the other validation catalogues, even at high significance. It might be no coincidence that, except for , the highest recovery rates for systems monitored over a sufficiently long baseline are found for the high-quality Griffin solutions (Table 12).
As discussed above, the success rate for the main orbital parameters (, , , and ) is usually well above 80% for the whole set of validation catalogues considered here (Griffin, SB9, APOGEE, WISE, and ASAS-SN). As expected, the performance increases for relatively short-period systems for which at least one orbital period is covered by the observations (about 2.5 years for the RVS spectroscopic processing data set considered here). In this case, the recovery rate commonly exceeds 85-90%. The WISE sample of (near-)contact binaries allows us to investigate the performance of our period search at high frequencies. We recover the WISE photometric periods at a level of about 80-85%. It increases to about 90% for the DEBs with wider orbits discussed by Rowan et al. (2022). One can also note the excellent success rate for the pulsation periods of the RR Lyrae stars that are below one day (Sect. 9.3). The good agreement on the period with both EB data sets advocates for a deeper study using these data. Their combination will certainly result in an extensive analysis that could reveal various kinds of objects such as EW binaries and/or ellipsoidal variables. Such a task is beyond the scope of the present paper. Our analysis globally suggests that the orbital periods below about 10 days are generally robust provided the solution significance is not too low. We find a roughly constant success rate at the level of about 90-95% once the significance exceeds 15-20. However, although they can often be regarded as compatible with zero within the uncertainties, it appears that the eccentricities may be overestimated even if the significance is larger.
We ascribe the poor correspondence between the DR3 and Griffin/SB9 ’s to a slight zero-point offset at the km s-1 level. There is indeed no evidence for a significant bias with respect to APOGEE (Fig. 69). For the other parameters, no obvious systematic discrepancies are observed.
In order to obtain an order-of-magnitude estimate of the period whenever possible, a choice is made to systematically favour an SB1 solution over a TrendSB1 one when the period is estimated to be in the range 0.5 to 1.5 . As a result, an SB1 solution could have been assigned despite the fact that the length of the observations is insufficient. However, although the proportion strongly depends on the validation data set, it concerns less than 15% of the cases. Conversely, overall about 79% of the TrendSB1 solutions are associated to binaries with periods significantly longer than the observations.
Our ability to recover the orbital parameters quite significantly hinges on the retrieval of the true period. When only counting the binaries for which this condition is satisfied in the reference catalogues discussed in Sects. 9.4.1 to 9.4.3, we find an increase of 5-10% in the recovery rate of , , and . For and , it typically reaches 95%. It is evident that other limitations, either related to the characteristics of the observations (e.g. number of RVS transits) or the very nature of the binary system (e.g. velocity semi-amplitude) can also potentially play a significant role in the quality of the solutions. Some clues about the sensitivity (or lack thereof) to the combination of some parameters can be gauged from plots such as Fig. 46. Besides suggesting a lack of dependency between the recovery rate of the period and some quantities (e.g. ruwe), it illustrates, as discussed above, the difficulty in successfully finding the period at low significance values. However, there may be more complex dependencies lurking in the data that cannot be easily and reliably quantified given the quite limited number of cross-matches with the validation catalogues (a few hundreds systems at most).
In this respect, the fact that only a relatively small fraction of the SB1’s in the validation catalogues have a spectroscopic deterministic DR3 solution is the consequence of the various filtering steps, both at CU4 and CU6 levels, described previously. In addition, we note that for the SB9 and Griffin’s sample, only a subset of the reference catalogue was used. The case of the SB9, and the steps that led to the relatively small overlap with respect to DR3 (see Sect. 9.4.2) are discussed in detail in Appendix G. Similar arguments apply to the other external catalogues and explain the limited number of benchmark binaries with a spectroscopic deterministic solution in DR3, although the proportion of systems lost at every filtering step may markedly vary. As an illustration, only the very brightest ASAS-SN eclipsing binaries get a DR3 solution. While the range is not an issue, we find that 90% have a (combined) magnitude fainter than 12 mag. Similar figures apply to the WISE sample. Regardless of these intrinsic limitations, it is naturally expected that the completeness level will substantially improve in DR4 thanks to the larger amount of RVs collected and improvements in data analysis.
We finally note that a few systems securely reported as double-lined in the literature have received an SB1-like solution (either eccentric or trend). The causes for this are multiple. For instance, the spectral signature of the secondary could be too weak to be detected by Gaia or the lines in the composite spectrum cannot be fully deblended at the moderate resolution of the RVS, as discussed in Sect. 8.2.3. The various cases which can be encountered are discussed in Damerdji et al. (2024).
10 Additional considerations
10.1 Post-filtering
The results of the sole NSS-SB1 processing (see Sect. 8) are providing a plethora of orbital solutions covering a large domain of the values for the orbital parameters. The pipeline nature of the processing necessarily implies that part of the solutions could be wrong or spurious (whatever could go wrong will go wrong at least once). Therefore, we scrutinised a subset of the results in order to detect surviving problems. Since the DPAC policy with respect to the final production of the DR3 catalogue is intended to favour purity of the list rather than completeness, it is necessary to blacklist solutions that are most certainly spurious. We call this procedure post-filtering and it intervenes between the internally filtered output of the pipeline and the publication of the catalogue. This post-filtering can be based on information coming from outside our internal validation. A clear flaw of this procedure is that this analysis will never identify all the pitfalls. Conversely, some objects could be blacklisted whereas the solution was correct. In the following, we address the identified problematic cases in the context of the classes associated to the sole spectroscopic processing.
10.1.1 Intrinsic variables
Since the NSS-SB1 processing is applied to time-series of variable RVs, it is expected to also treat some intrinsic variables that are exhibiting well-defined RV variations. This could be the case for example for Cepheids (actually the bright ones). The only firewall preventing the processing to deliver spurious orbital solutions for intrinsic variables is that the processing only considers an harmonic content of the RV curve that is Keplerian (or nearly so). Unfortunately, Cepheids RV curves, for example, are renowned for their harmonic content being quite similar to the Keplerian one. This has been pointed out by Imbert et al. (1985); Imbert (1987); Imbert et al. (1989). The problem generated by intrinsic variables has been pointed out in Sect. 9.4.4 and in Sect. 9.4.5 where some of them were detected in the binary catalogue, mainly at short periods and high eccentricities.
As part of DR3, Ripepi et al. (2023) published a complete analysis of the Gaia photometric curves for a little more than 15,000 Cepheids (of all classes), 799 having time-series of RVs. This list was cross-correlated with our SB1/SB1C list. They found 338 objects in common. Among them, the Cepheid nature is confirmed from the literature for 276 objects. After an eye inspection of photometric and RV curves, five of them were found doubtful Cepheids and have been rejected. Finally, 333 spurious SB1 solutions due to the Cepheid nature of the objects have been blacklisted from the SB-subcatalogue.
Along these lines, we also considered the case of RR Lyrae stars (of all classes too). Clementini et al. (2023) gathered a list of a little more than 270,000 such objects, among which some 1096 have time-series of RVs. The cross-match with our SB1 list yields 152 objects in common. Among these, 5 turned out to have a discrepancy between the period from our SB1 pipeline and the one derived by Clementini et al. (2023). Their RR Lyrae classification were thus considered as insecure. Therefore, we rejected (blacklisted), from the SB1 list, 147 objects considered as bona fide RR Lyrae. In summary, we rejected from our SB1 list some 480 sources as being intrinsic variables. Although a few pulsating stars (Cepheids, RR Lyrae) are rejected from our SB-subcatalogue, they are nevertheless studied as pulsators elsewhere (see e.g. Eyer et al. 2023).
Another class of intrinsic variable objects exhibiting large photometric amplitudes are the Long-Period Variables (LPVs). No similar secure list exists in DR3 for the LPVs. A tentative preliminary list exists that is based on the analysis of the photometry: the second Gaia catalogue of long-period variable candidates (Lebzelter et al. 2023). However, the related results were not entirely accessible at the time of the execution of our pipeline. In addition, from the later provisional analysis, it became clear that this preliminary list most probably also contains a lot of other objects and in particular ellipsoidal variables.
We will illustrate here our concerns by further studying the case of Gaia DR3 5829160851462523008. This object was not known to vary before Gaia. It appears in the catalogue of LPV candidates with a derived photometric period of 90.5 d. This object also appears in the SB-subcatalogue, produced in the present work, as a binary with a period of 178.886 d. Figure 47 exhibits both the RV curve and the photometric light curve of this object with the phase computed according to the RV period. From this figure, it is clear that the RV curve is unimodal, as it should be, whereas the light curve presents two cycles in good agreement with the period attributed by Lebzelter et al. (2023). If the two maxima are very similar, the two minima are significantly different. In addition, the minima happen at phases where the RV curve crosses the systemic velocity. All this strongly suggests that this star is an ellipsoidal variable. Therefore, at the time of DR3, it appeared hazardous to filter our solutions on the basis of the existing LPV catalogue. Objects like Gaia DR3 5829160851462523008 would have been unduly rejected and we cautiously decided to not reject LPVs on this basis. Certainly, a sound analysis necessitates the concomitant study of RV and photometry, which was not foreseen for DR3. This should be done in the subsequent FPR (see Gaia Collaboration et al. 2023c) where some 20 % of the LPV candidates are expected to be ellipsoidal variables. We encourage the user of the SB-subcatalogue to filter out true intrinsic variables thanks to these subsequent results.
Yet another source of contaminants are main-sequence pulsators. The main periods of Scuti stars are typically below 0.2 d (Gaia Collaboration et al. 2023b) and therefore lie outside the frequency range explored by our pipeline (even considering aliasing). However, Doradus stars are a more serious concern because they pulsate with periods in the range 0.3-1.4 d (Gaia Collaboration et al. 2023b) and exhibit RV variations of the order of a few km s-1 (e.g. Mathias et al. 2004).
The problem of the rejection of intrinsic variables is certainly a difficult one and much progress is still to be made on this topic. As mentioned above, the measurement of RVs by cross-correlating the spectra with template prevents to identify the origin of the RV variations.
10.1.2 Double-line transits
Since the RVs are originally measured by the STA processing on the basis of a per transit approach, there is initially no constraint from transit to transit. In particular, the same object could exhibit a single-line spectrum at a given time and a composite spectrum at another moment. As part of the spectroscopic processing, a value representing the percentage of transits leading to double-line detection is computed. If the star is constant or assumed non-variable, the median velocity is computed, a cut-off is then applied on this percentage in order to avoid the propagation of the effect of suspicious secondary RVs. If no transit exhibits a composite spectrum, the object presenting variations is processed as a single star. On the opposite, if a sufficient number of transits show a double-line spectrum, the object is processed as an SB2 (see Damerdji et al. 2024, for details). If the assumed constant object presents only a minority of double-line transits, the latter are ignored and the object is analysed on the basis of the sole single-line data. The presence of 10 % or more double-line transits is considered as a source of difficulties for the processing of single constant objects. It has been proposed a posteriori that the same rule (as the one applied to constant stars) should apply to the NSS-SB1 chain. This led to the rejection of 1475 SB1 objects and of one SB1C object.
10.1.3 Scan-angle dependent signals
The RVS being a slitless spectrometer, the knowledge of the exact position of a star is necessary to have a correct calibration of the wavelength for its observed spectrum (Sartoretti et al. 2018). The astrometric solution assumes that all the objects are point-like, and the objects presenting incompatibility with this hypothesis have potentially wrong RVs. They are thus associated to large values of the ruwe parameter. This point was already mentioned in Sect. 2. To these deviating objects is associated an excess noise compared to the global astrometric solution (labelled astrometric_excess_noise). One of the main categories of the non-point-like objects are close pairs (physical or not) unresolved or partially resolved. A blend of objects that presents some elongated image will induce the following effect. The measured position of the star will have a position dependent on how the elongated axis of the pair enters the astrometric CCD field. Thus, the resulting RVs are dependent on the orientation of the main axis of the elongated objects with respect to the scan. It depends on the scan angle in a rather smooth manner. This potentially could generate, from transit to transit, RV variations that are a simple function of the scan angle. In particular, solutions for RV-constant objects could be associated to solutions mimicking a spectroscopic orbit at a period of 62.97 d; these orbits are consequently spurious. During the precession cycle, the scan angles of the various transits are performing a complete round. The phenomenon manifests itself as an excess of solutions with small and thus small mass function (because the period is fixed). It is visible in Fig. 31 and is further illustrated in Fig. 48. Clearly, most of the solutions are spurious and are associated to large values of the astrometric excess noise. The suspicious span the approximate interval between 0.8 and 16 km s-1. Another illustration can be found in Gaia Collaboration et al. (2023a), where it is also shown that the phenomenon could be related to a particular behaviour of ipd_gof_harmonic_amplitude. A detailed investigation of this behaviour is to be found in Holl et al. (2023a).
Thus, the epoch RVs are affected by unmodelled astrometric shifts along the scan direction. In order to alleviate the problem, the objects in the class SB1 were first selected on the basis of an astrometric ruwe larger than 1.4 and a solution leading to a period within 3% of the precession period (62.97 d). In addition, we selected objects obeying the condition
where is the mean of the uncertainties on the individual epoch RVs over the full time-series. Fig. 49 illustrates the distribution of the astrometric_excess_noise as a function of the parameter for the selected sample. One clearly sees that all the values are lower than three times the astrometric_excess_noise suggesting that these orbital solutions are spurious and are due to a scan-angle effect. The semi-amplitude is too small relatively to the error induced by the astrometric solution. It has been decided to reject and blacklist the 87 objects presenting these characteristics.
A similar approach concerning the satellite revolution period around the Sun (i.e. within 10% of one year) led to the rejection of another 77 sources.
10.1.4 Further remarks
All in all, 2107 objects were blacklisted for the above-mentioned three reasons. The mere sum of the individual rejected objects amounts to 2119, but 12 objects appear in two lists. The post-filtering described here above does not concern classes that have been combined. It could happen that a successful combination leads to a removal of some objects from the purely spectroscopic classes. If later the global fit is not successful, the object will disappear from the catalogue and will not be re-injected in any purely spectroscopic classes. Thus, some countings could be difficult to perform.
Although the post-filtering has been applied with great caution, it also happens that some problems persist. One of these cases is the following. Some objects exhibit a good SB1 or SB1C solution with a well established significant orbital semi-amplitude for Gaia. However, these solutions could sometimes be doubtful because the same object observed in the framework of the APOGEE DR17 survey (Abdurro’uf et al. 2022) or LAMOST survey (Cui et al. 2012; Bashi et al. 2022) appears to be constant in RV as detailed in Sect. 8.2.4. This anomaly has not been fully explained but it could be related in some cases to the presence of a nearby neighbour (see below). These objects have not been rejected. We encourage the catalogue user to be very cautious when using the present SB-subcatalogue despite the fact that persisting problems should only represent a minority. All the remaining and identified problems should be corrected in the future DR4.
10.2 An explanation for Gaia DR3 3376949338201650112
In Sect. 8.2.4, we report the existence of SB1 solutions presenting a marked periodic variation, whereas ground-based data conclude to the constancy of the related object. Here below, we propose an explanation for this discrepancy by investigating in details the case of Gaia DR3 3376949338201650112. This example has the advantage to be enlightening and will allow us to correct the pipeline for future releases.
The astrometric solution for this object is characterised by a ruwe of 25.35, pinpointing to a problem with the fit. Referring to the IPD parameters (Image Parameter Determination, see e.g. Holl et al. 2023a), we remark that ipd_gof_harmonic_amplitude is equal to 0.164 which is certainly different from zero. This is a clue that the object is badly treated as a single object. The value of ipd_frac_multi_peak turns out to be 84 suggesting that the object is expected to be double or multiple on the plane of the sky: most probably a resolved or, more precisely, a partially resolved pair (according to the definition of Holl et al. 2023a). As explained by these authors, such pairs (or higher order) have a predisposition to suffer from scan-angle effects that induce a bias on the measured RVs that is function of the relative orientation of the scanning direction on the sky at the time of the observations. This is confirmed for Gaia DR3 3376949338201650112 by Fig. 50 where it becomes clear that the measured RVs are function of the scan angle. The parameter ipd_gof_harmonic_phase adopts a value of 103.4 ° that suggests that the line joining the two stars of the pair has a position angle on the sky of 13.4 °, after the application of a 90 ° correction as explained in section 3.1.2 of Holl et al. (2023a). The semi-amplitude of the fake orbital solution is about 16.6 km s-1 which translates according to equation 2 of Holl et al. (2023a) into a separation of the pair of 114 mas. The existence of such a pair is confirmed by a preliminary run of AGIS (AGIS 4.1) linked to the future DR4 that actually detects two objects Gaia DR4 3376949342501945344 and Gaia DR4 3376949342501945472 that are separated by some 200 mas and are exhibiting an orientation of 9 °, thereby supporting our interpretation.
In principle, the bias due to the scan-angle effect is intended to produce variability with a period either of 62.97 d or of 1 y. This property has been used in Sect. 10.1.3 to apply post-filtering. However, in the present case, the period is 0.657582 d which does not appear to be compatible and could shed some doubt on our interpretation. The spectral window associated to the DR3 RV time-series corresponding to this object is given in Fig. 51. Besides the main peak at d-1 as usual, it exhibits an important peak at d-1 that is exceeding a value of 0.8; this is an indication that isolated aliasing could be important for this particular time-series. An aliasing with d-1 explains the ambiguity between a period of 0.657582 d and one of one year. The combination of the aliasing phenomenon and of the scan-angle effect explains the existence of such fake SB1 objects in our SB-subcatalogue: such objects were not post-filtered, since the post-filtering was based on the 62.97 d or 1 y timescales. We conclude that it would be more appropriate to include a statistical test inspired by graphs like the one shown in Fig. 50 rather than on a criterion based on the period. However, a true periodicity with these timescales will imply a similar behaviour of the RV dependency on the scan angle even if no scan-angle effect is present.
The explanation given here does not apply to the other object cited in Sect. 8.2.4 (Gaia DR3 702159502868472832).
10.3 The case of astrometric binaries
In Sect. 10.1.3 and in Sect. 10.2, we discussed the effect of astrometric close pairs on the determination of RVs. A close pair fixed on the sky could often generate a scan-angle effect and lead to corrupted RV values but also could generate fake orbital motions. An additional degree of complexity is reached if one considers astrometric binaries. Due to their binarity, these close pairs are not fixed on the sky and the main axis of the pair is rotating due to the orbital motion. Therefore, there is interaction between the above-mentioned scan-angle effect and the astrometric orbital motion itself. The situation still increases in complexity if the astrometric orbital motion and the spectroscopic RV changes correspond to two aspects of the very same orbital motion and thus have the same period. In this latter case, the correction presents a timescale equal to the phenomenon we are studying. The correcting scheme is explained in section 3.3.1 of Holl et al. (2023a) and is also illustrated there with the (somewhat extreme) example of Gaia DR3 6631710607341412096. For details, we refer to this paper.
The effect addressed here was not anticipated and was not fully understood at the time of the final delivery of the NSS code for DR3 processing. As a consequence, some of our solutions are damaged by this phenomenon. In addition, those pertaining to the AstroSpectroSB1 class should be the most affected. For DR4, a preliminary run of the NSS astrometric channel has been included in order to compute a correction before the CU6 processing performing the RV measurements.
11 The SB-subcatalogue
11.1 Description of the output
The present first catalogue of pure spectroscopic solutions is part of the Gaia Data Release 3. The orbital solutions (in class SB1 or SB1C as indicated in the entry nss_solution_id) are part of an auxiliary table named gaiadr3.nss_two_body_orbit. We output in this table the classical Campbell orbital parameters introduced in Sect. 3.1 for the SB1 class (, , , , , ) along with their estimated uncertainties (1). The same is done for circular orbit parameters (, , , ) introduced in Sect. 3.2 and corresponding to the SB1C class. Beyond the orbital parameters themselves, this table also contains some indicators of the quality of the solution (see Sect. 6.4). The flags associated to the solution (see Sect. 6.5) are also provided. The boolean mask bit_index defines the parameters that are actually fitted, and corresponds for the SB1 class (respectively SB1C) to the value 127 (respectively 31). In addition, is also provided the upper triangle of the correlation matrix of the parameters. It is expressed as corr_vec that contains the cross-coefficients folded into a vector and corresponding to the ordering of the adopted list of parameters. The first term of the vector gives the element (0,1) of the matrix, the element gives the element (0,2) of the matrix and element the (1,2) one. Elements , , correspond respectively to the matrix elements (0,3), (1,3) and (2,3) and so on. This is thus a column-major storage scheme. The details of the output are described in Table 15 for the class SB1 and in Table 16 for the SB1C one. The table contains 181327 objects (with solutions) in the SB1 class and 202 in the SB1C one. The marked difference in the counting is due to our decision to consider eccentric solutions down to as small eccentricities as possible before going circular. This was described in Sect. 6.3.
We also released the trend solutions as discussed in Sect. 3.3 and in Sect. 6.3.5. The objects belonging to the TrendSB1 class are part of the auxiliary table gaiadr3.nss_non_linear_spectro. The subtype of trend solutions (linear or second-order trends) are indicated in nss_solution_id (respectively FirstDegreeTrendSB1 and SecondDegreeTrendSB1). They are characterised by a bit_index of 7 and 15, respectively. A few statistics about the quality of the fit are also given. The table contains 24083 objects with solutions of type FirstDegreeTrendSB1 and 32725 objects with SecondDegreeTrendSB1 ones, thus totalling 56808 objects with solutions in the class TrendSB1. The details of the content of the latter table are described in Table 17. We recall that the StochasticSB1 solutions are not delivered in DR3.
11.2 Weaknesses, caveats and recommendations
The present section gathers the caveats dispersed all over the paper and delivers a summary of the characteristics of the catalogue that could be beneficial to the general user. The RVs studied here have the quality and shortcomings of the measurements of individual (epoch) RVs for constant stars. The properties and precision of these RVs for SB1 studies should be of the same order. Therefore, we refer the reader to the work of Katz et al. (2023) concerning the characteristics and condition of utilisation of these data. Seabroke et al. (in preparation) is also a work to be consulted.
The information extracted from variations of the RVs could be delicate to use. Firstly, it must be clear that the way the measurements are performed (cross-correlation with a template) makes the implicit assumption that the line is supposed to vary without changing its shape. This is not necessarily true. Two textbook cases at least come to mind. Some stars, notably the intrinsically variable ones, could exhibit dominant variations of their line shapes. In such a case, the variations do not represent a Doppler shift of the object as a whole and thus are not representative of an SB1 (see Sect. 10.1.1). In principle, the fact we are fitting models where the harmonic content is typical of a Keplerian motion is certainly a protection against variable stars appearing in the SB1 class. However, some variable stars, like Cepheids are well-known to display RV changes that closely mimic the Keplerian motion (e.g. Imbert et al. 1989). As explained in Sect. 10.1.1, efforts have been made to post-filter the Cepheid and the RR Lyrae stars, but this post-filtering is certainly not complete for the time being. In addition, no such step was made for other intrinsic variables in DR3 (notably Long-Period Variables).
Another case of trickery is the one of true SB2 stars, where the two lines never fully deblend. They are appearing most probably in the SB1 channel. The RV variations are then due to the corresponding change of the full profile made by the blended lines. This has little effect on the period determination but the impact on the eccentricity and mainly on the semi-amplitude might be huge (Sect. 8.2.3). The related solutions are characterised by a too small value of and of the mass function. They could be considered as suspicious but we can not have any certitude that all of these types of solutions are necessarily wrong.
The algorithm used is intended to fit orbital solutions to the RV curves that it analyses. It is doing so but is not protected against some anomalies. The RV data set suffers from all the problems not attenuated by the CU6 processing and is not particularly protected against outliers, miss-classified double-line objects, intrinsic variables, emission-line objects and other problematic situations (e.g. badly normalised continuum, etc…). In addition, it is not (for the moment) able to properly deal with higher degree systems. It is possible that the code detects a binary motion, but this one could be quantitatively corrupted by the presence of additional components (see Sect. 8.2.2). In the future (DR4), we hope to include a cleaning (whitening) of the dominant signal followed by a new search for a second period or a trend. However, the presence of two signals renders the statistical analysis much more complicated.
Instrumental effects are also present. The scan-angle effect notably induces periodic RV variations for constant stars that do not fulfil the assumptions used in astrometry (objects are assumed to be isolated and thus single). Close astrometric pairs could thus mimic SB1 RV variations (see the work by Holl et al. 2023a), and are polluting the present sample (even if some post-filtering took place, see also Sect. 10.1.3). If, in addition, the close pair is rotating on the plane of the sky as is the case for astrometric binaries, the measured RV could be corrupted (see Sect. 10.3). This problem will be corrected for DR4.
Besides, it should be noted that all decisions are taken in a statistical way; this means that the output is statistically correct but that some incorrect individual fits can survive near the borders where the decisions are taken (e.g. between an orbital fit and a trend solution for example).
We recommend to the user of the catalogue to keep in mind the weaknesses and caveats described here above. We also strongly recommend, when possible, the use of the various quality factors accompanying each solution. If the catalogue is expected to be well-behaved from the statistical point of view, the situation could be more intricate for individual objects or particular set of objects. It remains that biases in the catalogue could be present from the point of view of stellar population studies. Such an analysis is beyond the scope of the present paper. For the time being, the genuine selection function of the catalogue is poorly known and any user should be aware of this and behave accordingly. The user is strongly encouraged to adapt the selection of solutions to their particular objective instead of using it blindly. They should combine the various delivered quality factors according to their aim. We advise a cautious approach. An example of interesting complex combinations is detailed in Bashi et al. (2022). To some extent, other information can be considered as quality factors (such as flags but, mainly, the number of points in the time-series ). They could also be included in the selection process.
The global weak point of the algorithms could be the determination of the period even if here above it is finally shown that the success rate is at the very least 80% and is more often located in the range 85-95%. There is evidently at a very low number of transits, a difficulty to correctly determine the period that forced us to neglect data with less than 10. However, at slightly larger , the situation is not so simple. From the simulations described in Sect. 9.1, we conclude that a value of smaller than 17 is somewhat problematic (although it depends on the signal-to-noise ratio) for the period recovery. We also showed from the validation (see Sect. 9.2), that an overfitting phenomenon characterises part of the solutions associated to an in the range 10-15. One should further add that, at small , the periodogram has a larger propensity to produce isolated aliasing peaks (generating ambiguity due to aliasing). While interpreting the simulations, it appears that it is unclear if the determining parameter is the or the significance. According to these simulations, significance lower than 10-15 can be associated with difficulties to have the proper recovery rate. This seems to be confirmed by the validation in Sect. 9.4, in particular in the corner plot of Fig. 46. It is also clear, from Fig. 30, that the circularisation phenomenon is not outstandingly visible for significance much lower than 30. Therefore, we consider that depending on the studies performed, the significance should be cut between 15 and 25, or more conservatively at 30. If 10 is still affordable in some cases, the solutions characterised by a significance between 5 (our cut-off) and 10 are most probably hazardous to accept.
The spectroscopic solutions could be merged with photometric or astrometric ones, and in these cases the resulting output is certainly very secure for the period. If the user of the catalogue only considers the sole results of the spectroscopy fits, it is advisable to restrict the sample to some kind of gold sample. This can be done by performing additional selections in the catalogue: for example, a larger number of transits, a higher significance, periods in a dedicated range (between ten days and , or allowing for going down to smaller periods but at the expense of a selection on the basis of a larger significance), period with small uncertainties, solutions with negligible phase gaps (although the gap values are not provided in the DR3 catalogue), etc…
In Appendix J, we present indicative but simple examples of extraction of a gold sample from the catalogue.
12 Conclusions
In the framework of the Gaia Data Release 3, we present here the first version of the Gaia catalogue of spectroscopic orbital solutions (SB-subcatalogue). We present 181327 objects with a solution of class SB1 and 202 objects with a solution of class SB1C. We also present objects exhibiting trends of variations of their RVs with time.
The present catalogue constitutes by far the largest collection of objects with a spectroscopic orbital solution, as can be evaluated by comparison with the notorious SB9 compilation171717https://sb9.astro.ulb.ac.be/. The improvement in number of concerned objects is more than one (almost two) order of magnitude. The present SB-subcatalogue has also the major advantage of being homogeneous from the point of view of both data acquisition (properties) and data processing, surpassing the characteristics of existing compilation. The only exception is the typical number of measurements that constitutes the individual time-series, which remains limited (at least for the present DR3 catalogue). The situation will undoubtedly improve for future data releases.
After a description of the data entering the NSS-SB1 processing, we describe in details the logic and structure of the related code. The general philosophy behind it is explicited along with the mathematical formalism used. We also describe, after the main processing, the filtering of the results that we performed to discard bad solutions and improve the catalogue. Examples illustrating the quality of the output are given as well as a few examples of detected problematic cases. A full validation of the output is performed both on the basis of internal consistency and on the basis of external existing catalogues from the literature. The characteristics of the catalogue are further analysed. The way the SB-subcatalogue is included in the Gaia Data Release 3 is also described. A section is included that scrutinises the weaknesses and provides caveats and recommendations for the scientific use of the catalogue.
Data availability
The epoch RVs constituting the time-series analysed here have not been published in the context of the Gaia Data Release 3. The orbital solutions and trends we derived are available since June 13, 2022, in the ESA Gaia DR3 archive accessible at https://gea.esac.esa.int/archive. The same results can also be found at CDS Strasbourg via the link https://cdsarc.cds.unistra.fr/viz-bin/Cat?I/357.
Acknowledgements.
The authors are particularly thankful to Alain Detal who maintained over the years the Liège computer park with great efficiency. EG is also thanking F. Mélen for her consequent help in the manipulation of various computer codes. The authors are acknowledging the numerous scientific discussions that took place with Christos Siopis. The participation in this endeavour has been made possible through an Impulsion Credit from the University of Liège to EG. This work presents results from the European Space Agency (ESA) space mission Gaia. Gaia data are being processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC is provided by national institutions, in particular the institutions participating in the Gaia MultiLateral Agreement (MLA). The Gaia mission website is https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia. The Gaia mission and data processing have, in Liège, been continuously supported by the Belgian Politique Scientifique Fédérale (BELSPO) through PRODEX contracts Gaia DPAC: Binaires, étoiles extrêmes et objets du système solaire (convention PEA 4200090289; years 2007-2013), Gaia DPAC Early Mission Belgian consolidation (Gaia-be) (convention PEA 4000110152: years 2014-1016), Gaia DPAC Early Mission Belgian consolidation (Gaia-be2) (convention PEA 4000119840: years 2017-1019; years 2020-2022). TM acknowledges financial further support from BELSPO for contract PLATO mission development. We acknowledge support from an ARC grant (13/17-06; Massive Stars: drivers of the evolution of the Universe) for concerted research actions financed by the French community of Belgium (Wallonia-Brussels Federation). We are also largely indebted to the Fonds de la Recherche Scientifique (F.R.S.-FNRS, Belgium) for various supports. YD acknowledges support to the Algerian Centre de Recherche en Astronomie, Astrophysique et Géophysique of Bouzareah Observatory. GS acknowledges support from BELSPO/PRODEX under various contracts and in particular convention PEA 40000119826. AS acknowledges support from the Italian Space Agency (ASI) under contract 2018-24-HH.0 ”The Italian participation to the Gaia Data Processing and Analysis Consortium (DPAC)” in collaboration with the Italian National Institute of Astrophysics.References
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35
- Andrae et al. (2023) Andrae, R., Fouesneau, M., Sordo, R., et al. 2023, A&A, 674, A27
- Antoja et al. (2018) Antoja, T., Helmi, A., Romero-Gómez, M., et al. 2018, Nature, 561, 360
- Armstrong et al. (2015) Armstrong, D. J., Kirk, J., Lam, K. W. F., et al. 2015, A&A, 579, A19
- Bashi et al. (2022) Bashi, D., Shahaf, S., Mazeh, T., et al. 2022, MNRAS, 517, 3888
- Bevington & Robinson (2003) Bevington, P. R. & Robinson, D. K. 2003, Data reduction and error analysis for the physical sciences, 3rd edition (MA: McGraw-Hill)
- Binnendijk (1960) Binnendijk, L. 1960, Properties of double stars; a survey of parallaxes and orbits (University of Pennsylvania Press, Philadelphia)
- Bloomfield (1976) Bloomfield, P. 1976, Fourier analysis of time series: an introduction (Princeton University, John Wiley and Sons, Inc., New York)
- Boulkaboul et al. (2022) Boulkaboul, A., Damerdji, Y., Morel, T., et al. 2022, MNRAS, 517, 1849
- Brandt (2021) Brandt, T. D. 2021, ApJS, 254, 42
- Buder et al. (2018) Buder, S., Asplund, M., Duong, L., et al. 2018, MNRAS, 478, 4513
- Clementini et al. (2023) Clementini, G., Ripepi, V., Garofalo, A., et al. 2023, A&A, 674, A18
- Cropper et al. (2018) Cropper, M., Katz, D., Sartoretti, P., et al. 2018, A&A, 616, A5
- Cui et al. (2012) Cui, X.-Q., Zhao, Y.-H., Chu, Y.-Q., et al. 2012, Research in Astronomy and Astrophysics, 12, 1197
- Cumming (2004) Cumming, A. 2004, MNRAS, 354, 1165
- Damerdji et al. (2024) Damerdji, Y., Gosset, E., & Morel, T., et al. 2024, A&A, in prep.
- David et al. (2014) David, M., Blomme, R., Frémat, Y., et al. 2014, A&A, 562, A97
- Deeming (1975) Deeming, T. J. 1975, Ap&SS, 36, 137
- Deeming (1976) Deeming, T. J. 1976, Ap&SS, 42, 257
- Delchambre (2010) Delchambre, L. 2010, Recherche d’une évaluation robuste de l’excentricité dans des systèmes binaires d’étoiles, Master Thesis Liège University (available at https://orbi.uliege.be/handle/2268/322803)
- Eichhorn & Xu (1990) Eichhorn, H. K. & Xu, Y.-L. 1990, ApJ, 358, 575
- Endl et al. (2002) Endl, M., Kürster, M., Els, S., et al. 2002, A&A, 392, 671
- ESA (1997) ESA, ed. 1997, ESA Special Publication, Vol. 1200, The HIPPARCOS and TYCHO catalogues. Astrometric and photometric star catalogues derived from the ESA HIPPARCOS Space Astrometry Mission
- ESA, (1997) ESA, 1997, VizieR Online Data Catalog, I/239
- Evans et al. (2018) Evans, D. W., Riello, M., De Angeli, F., et al. 2018, A&A, 616, A4
- Eyer et al. (2023) Eyer, L., Audard, M., Holl, B., et al. 2023, A&A, 674, A13
- Fabricius et al. (2021) Fabricius, C., Luri, X., Arenou, F., et al. 2021, A&A, 649, A5
- Ford & Gregory (2007) Ford, E. B. & Gregory, P. C. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 371, Statistical Challenges in Modern Astronomy IV, ed. G. J. Babu & E. D. Feigelson, 189
- Fraser et al. (2008) Fraser, O. J., Hawley, S. L., & Cook, K. H. 2008, AJ, 136, 1242
- Frémat et al. (2023) Frémat, Y., Royer, F., Marchal, O., et al. 2023, A&A, 674, A8
- Fukushima (1996) Fukushima, T. 1996, AJ, 112, 2858
- Gaia Collaboration et al. (2023a) Gaia Collaboration, Arenou, F., Babusiaux, C., et al. 2023a, A&A, 674, A34
- Gaia Collaboration et al. (2018a) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2018a, A&A, 616, A1
- Gaia Collaboration et al. (2016a) Gaia Collaboration, Brown, A. G. A., Vallenari, A., et al. 2016a, A&A, 595, A2
- Gaia Collaboration et al. (2023b) Gaia Collaboration, De Ridder, J., Ripepi, V., et al. 2023b, A&A, 674, A36
- Gaia Collaboration et al. (2018b) Gaia Collaboration, Katz, D., Antoja, T., et al. 2018b, A&A, 616, A11
- Gaia Collaboration et al. (2016b) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016b, A&A, 595, A1
- Gaia Collaboration et al. (2023c) Gaia Collaboration, Trabucchi, M., Mowlavi, N., et al. 2023c, A&A, 680, A36
- Gaia Collaboration et al. (2023d) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023d, A&A, 674, A1
- Golub & Pereyra (2003) Golub, G. & Pereyra, V. 2003, Inverse Problems, 19, R1
- Gosset et al. (2001) Gosset, E., Royer, P., Rauw, G., Manfroid, J., & Vreux, J. M. 2001, MNRAS, 327, 435
- Griffin (1969) Griffin, R. F. 1969, MNRAS, 145, 163
- Griffin (2007) Griffin, R. F. 2007, The Observatory, 127, 113
- Halbwachs et al. (2005) Halbwachs, J. L., Mayor, M., & Udry, S. 2005, A&A, 431, 1129
- Halbwachs et al. (2023) Halbwachs, J.-L., Pourbaix, D., Arenou, F., et al. 2023, A&A, 674, A9
- Heck et al. (1985) Heck, A., Manfroid, J., & Mersch, G. 1985, A&AS, 59, 63
- Holanda et al. (2018) Holanda, N., Barros, P. P., & Silva, A. N. 2018, New A, 65, 21
- Holl et al. (2023a) Holl, B., Fabricius, C., Portell, J., et al. 2023a, A&A, 674, A25
- Holl et al. (2023b) Holl, B., Sozzetti, A., Sahlmann, J., et al. 2023b, A&A, 674, A10
- Imbert (1987) Imbert, M. 1987, A&A, 175, 30
- Imbert et al. (1985) Imbert, M., Andersen, J., Ardeberg, A., et al. 1985, A&AS, 61, 259
- Imbert et al. (1989) Imbert, M., Andersen, J., Ardeberg, A., et al. 1989, A&AS, 81, 339
- Jayasinghe et al. (2023) Jayasinghe, T., Rowan, D. M., Thompson, T. A., Kochanek, C. S., & Stanek, K. Z. 2023, MNRAS, 521, 5927
- Katz et al. (2019) Katz, D., Sartoretti, P., Cropper, M., et al. 2019, A&A, 622, A205
- Katz et al. (2023) Katz, D., Sartoretti, P., Guerrier, A., et al. 2023, A&A, 674, A5
- Kawata et al. (2018) Kawata, D., Baba, J., Ciucǎ, I., et al. 2018, MNRAS, 479, L108
- Kervella et al. (2019) Kervella, P., Arenou, F., Mignard, F., & Thévenin, F. 2019, A&A, 623, A72
- Khoperskov et al. (2020) Khoperskov, S., Gerhard, O., Di Matteo, P., et al. 2020, A&A, 634, L8
- Kochanek et al. (2017) Kochanek, C. S., Shappee, B. J., Stanek, K. Z., et al. 2017, PASP, 129, 104502
- Koppelman et al. (2019) Koppelman, H. H., Helmi, A., Massari, D., Roelenga, S., & Bastian, U. 2019, A&A, 625, A5
- Kunder et al. (2017) Kunder, A., Kordopatis, G., Steinmetz, M., et al. 2017, AJ, 153, 75
- Lebzelter et al. (2023) Lebzelter, T., Mowlavi, N., Lecoeur-Taibi, I., et al. 2023, A&A, 674, A15
- Lester (2020) Lester, K. V. 2020, PhD thesis, Georgia State University, United States
- Lindegren (2019) Lindegren, L. 2019, Gaia Data Processing and Analysis Consortium (DPAC) technical note GAIA-C3-TN-LU-LL-125, http://www.cosmos.esa.int/web/gaia/public-dpac-documents
- Lindegren et al. (2018) Lindegren, L., Hernández, J., Bombrun, A., et al. 2018, A&A, 616, A2
- Lindegren et al. (2021) Lindegren, L., Klioner, S. A., Hernández, J., et al. 2021, A&A, 649, A2
- López-Corredoira et al. (2020) López-Corredoira, M., Garzón, F., Wang, H. F., et al. 2020, A&A, 634, A66
- López-Corredoira & Sylos Labini (2019) López-Corredoira, M. & Sylos Labini, F. 2019, A&A, 621, A48
- Lucy & Sweeney (1971) Lucy, L. B. & Sweeney, M. A. 1971, AJ, 76, 544
- Marchetti et al. (2019) Marchetti, T., Rossi, E. M., & Brown, A. G. A. 2019, MNRAS, 490, 157
- Mathias et al. (2004) Mathias, P., Le Contel, J. M., Chapellier, E., et al. 2004, A&A, 417, 189
- Mazeh (2008) Mazeh, T. 2008, in EAS Publications Series, Vol. 29, EAS Publications Series, ed. M. J. Goupil & J. P. Zahn, 1–65
- Monari et al. (2019) Monari, G., Famaey, B., Siebert, A., Wegg, C., & Gerhard, O. 2019, A&A, 626, A41
- Mowlavi et al. (2023) Mowlavi, N., Holl, B., Lecoeur-Taïbi, I., et al. 2023, A&A, 674, A16
- Nelson & Robb (2017) Nelson, R. H. & Robb, R. M. 2017, Information Bulletin on Variable Stars, 6203, 1
- Nitschai et al. (2020) Nitschai, M. S., Cappellari, M., & Neumayer, N. 2020, MNRAS, 494, 6001
- Perryman (2009) Perryman, M. 2009, Astronomical Applications of Astrometry: Ten Years of Exploitation of the Hipparcos Satellite Data (Cambridge University Press)
- Petrosky et al. (2021) Petrosky, E., Hwang, H.-C., Zakamska, N. L., Chandra, V., & Hill, M. J. 2021, MNRAS, 503, 3975
- Pourbaix et al. (2022) Pourbaix, D., Arenou, F., Gavras, P., et al. 2022, Gaia DR3 documentation Chapter 7: Non-single stars, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium. Online at ¡A href=“https://gea.esac.esa.int/archive/documentation/GDR3/index.html”¿ https://gea.esac.esa.int/archive/documentation/GDR3/index.html¡/A¿, id. 7
- Pourbaix et al. (2004) Pourbaix, D., Tokovinin, A. A., Batten, A. H., et al. 2004, A&A, 424, 727
- Press et al. (2007) Press, W. H., Teukolsky, S. A., Vetterling, W. T., & Flannery, B. P. 2007, Numerical recipes: the art of scientific computing, 3rd Edition (Cambridge University Press)
- Prevot et al. (1985) Prevot, L., Andersen, J., Ardeberg, A., et al. 1985, A&AS, 62, 23
- Price-Whelan et al. (2017) Price-Whelan, A. M., Hogg, D. W., Foreman-Mackey, D., & Rix, H.-W. 2017, ApJ, 837, 20
- Price-Whelan et al. (2020a) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020a, ApJ, 895, 2
- Price-Whelan et al. (2020b) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020b, ApJ, 895, 2
- Ramos et al. (2018) Ramos, P., Antoja, T., & Figueras, F. 2018, A&A, 619, A72
- Recio-Blanco et al. (2023) Recio-Blanco, A., de Laverny, P., Palicio, P. A., et al. 2023, A&A, 674, A29
- Riello et al. (2018) Riello, M., De Angeli, F., Evans, D. W., et al. 2018, A&A, 616, A3
- Riello et al. (2021) Riello, M., De Angeli, F., Evans, D. W., et al. 2021, A&A, 649, A3
- Ripepi et al. (2023) Ripepi, V., Clementini, G., Molinaro, R., et al. 2023, A&A, 674, A17
- Rowan et al. (2022) Rowan, D. M., Jayasinghe, T., Stanek, K. Z., et al. 2022, MNRAS, 517, 2190
- Sartoretti et al. (2022) Sartoretti, P., Blomme, R., David, M., & Seabroke, G. 2022, Gaia DR3 documentation Chapter 6: Spectroscopy, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium. Online at ¡A href=“https://gea.esac.esa.int/archive/documentation/GDR3/index.html”¿ https://gea.esac.esa.int/archive/documentation/GDR3/index.html¡/A¿, id. 6
- Sartoretti et al. (2018) Sartoretti, P., Katz, D., Cropper, M., et al. 2018, A&A, 616, A6
- Sartoretti et al. (2023) Sartoretti, P., Marchal, O., Babusiaux, C., et al. 2023, A&A, 674, A6
- Schlesinger (1910) Schlesinger, F. 1910, Publications of the Allegheny Observatory of the University of Pittsburgh, 1, 33
- Seabroke et al. (2021) Seabroke, G. M., Fabricius, C., Teyssier, D., et al. 2021, A&A, 653, A160
- Sharma (2017) Sharma, S. 2017, ARA&A, 55, 213
- Siopis et al. (2024) Siopis et al. 2024, A&A, in prep.
- Soubiran et al. (2018) Soubiran, C., Jasniewicz, G., Chemin, L., et al. 2018, A&A, 616, A7
- Sterne (1941) Sterne, T. E. 1941, Proceedings of the National Academy of Science, 27, 175
- Tatum (1968) Tatum, J. B. 1968, MNRAS, 141, 43
- Trick et al. (2019) Trick, W. H., Coronado, J., & Rix, H.-W. 2019, MNRAS, 484, 3291
- Wilson & Hilferty (1931) Wilson, E. B. & Hilferty, M. M. 1931, Proceedings of the National Academy of Science (USA), 17, 684
- Zechmeister & Kürster (2009) Zechmeister, M. & Kürster, M. 2009, A&A, 496, 577
Appendix A Historical developments
Once the period evaluated, the general model to fit (Eq. 1) is still non-linear in some of the parameters (mainly the eccentricity ). This implies that the of the fit could present a few or even numerous minima in the parameter space, and to find the global one is a true challenge. In particular, the derivation of the eccentricity is a notorious pitfall. The search for the global minimum might be performed empirically by guessing its approximate position and then by descending the valley by classical numerical methods (such as the Levenberg-Marquardt algorithm). However, in a pipeline as the one developed for Gaia, identifying the location of the global minimum is uncertain and a specific type of search is necessary. This is particularly frustrating because a well-trained expert eye is able to estimate a rough value for the eccentricity with great efficiency. In this context, we explored the possibilities to adopt alternative approaches. We acted along two main axes. On one hand, we tried to invent a method based on pattern recognition in the phase diagram of the RVs folded with the candidate periods. Such an approach would allow having a first guess for the eccentricity, which could be used for further classical derivation. In order to apply pattern recognition techniques, it is mandatory to define an interpolation function in the phase diagram in order to obtain a robust description of the shape of the RV curve. Several tools have been explored such that the nearest-neighbour interpolation, the linear interpolation, the interpolation by cosines, the local Akima periodic interpolation, the periodic interpolation by cubic splines, the periodic interpolation by Catmull-Rom splines. We also explored the possibilities of performing classical adjustment by least-square methods. In particular, we considered classical polynomials, Fourier series, Bézier curves (using De Casteljan algorithm) and finally, cubic B-splines in their open uniform version. Once the interpolation curve is chosen and the shape is defined, we tested various pattern recognition methods such as algorithms for regression based on the nearest neighbourhood, algorithms for regression based on trees and in particular on extremely randomised trees and, finally, algorithms for regression based on neural networks. Full details of this exploration and detailed references are given in the work of Delchambre (2010).
On the other hand, we developed an approach based on genetic algorithms. These algorithms have the advantage to circumvent any preliminary guess of the solution. This is at the expense of being obliged to explore (even cleverly) the space of the parameters using the generation of random numbers (Monte Carlo methods). This work is described in details in Delchambre (2010). The use of genetic algorithms is a trade-off between the probability to effectively detect the global minimum and the consumption of computer time. The code using genetic algorithms were written, delivered and tested. The detection of the global minimum was possible but was implying a prohibitively large amount of computer time despite strong efforts made to accelerate it. The large amount of objects to treat in the Gaia processing suggests using much more efficient algorithms. The genetic algorithm code was not activated for the operational runs.
Other methods are now available to derive orbital solutions. They became widespread in the domain of exoplanets, where the orbits are particularly incompletely covered and complex. The majority of these methods are based on Monte Carlo simulations. Very often, they use the Markov Chain Monte Carlo (MCMC) method (see e.g. Sharma 2017) often utilised in conjunction with Bayesian approaches. These algorithms necessitate a dense coverage of the full space of priors, at the expense of the computer time. Comparing the works of Price-Whelan et al. (2017) and Price-Whelan et al. (2020b) pinpoints the strong impact that could have the choice of priors on the resulting posteriors. The Bayesian methods have admittedly the big advantage to potentially provide marginal posterior probabilities for a series of models and thus estimate how much the observational data could favour a particular model. According to Ford & Gregory (2007), the challenge of estimating marginal posterior probabilities is still a crucial problem and obvious estimators are suffering from poor convergence properties. In addition, for the preliminary DR3 catalogue, the directives were that we should output one and only one model per treated object, thus mitigating the interest of using Bayesian methods. In addition, these Monte Carlo methods are computer time-consuming and were not acceptable for the DR3 processing.
We finally adopted a method based on fully analytical and classical numerical developments.
Appendix B Practical details on the processing
The Data Processing Centre at CNES (DPCC181818https://www.cosmos.esa.int/web/gaia/dpac) is located at the Centre National d’Etudes Spatiales in Toulouse (France). It ran the spectroscopic processing chain (CU6) and the object processing chain (CU4). Globally, it ran the whole set of NSS codes and, in particular, the spectroscopic orbital solution chain described in the present paper. Our NSS-SB1 processing chain delivers the results to the combiner. The Main Data Base (MDB) is fed with solutions from this combiner. If the combiner finds no other solutions to combine with, the spectroscopic results are directly transmitted to the MDB. If a combination is possible, only the output of the combination is saved.
In the framework of the Gaia endeavour, the DPCC is hosting a cluster of 250 computers (6000 cores, 40 TBytes of RAM and 7 PBytes of HDFS disk storage). The resources are managed by an Hadoop system. The codes are integrated into a framework SAGA (System of Accomodation of Gaia Algorithms) developed by Thales. The spectroscopic processing was executed over 2500 cores and occupied 17 TBytes of RAM. Some 3106 hours of CPU were necessary, equivalent in real time to 120 days. Concerning the NSS chain, the full treatment (Ingestion, Processing, and Post-Processing) was executed on the Hadoop cluster over 1400 cores and necessitated 16105 days of CPU corresponding to an effective duration of 17 days. All the codes were written in Java 8.0. The version of the code tagged and executed for the operational run was 20.10.
Appendix C The separable variable technique applied to the Keplerian model fit
C.1 A reminder of the Levenberg-Marquardt method
Eq. 10 links the dependent variables (the RVs of the model) to the independent ones (the observing times). Hereafter, we represent these two variables by the column arrays and respectively. represents the Keplerian function, such that
being the array of 6 parameters where the superscript denotes the transpose operation. The Newton-Raphson algorithm is based on the Taylor first order development
where , and denotes the iteration number. is the Jacobian matrix which elements are
The objective function to be minimised is the quadratic sum of residuals
with being the inverse of the data variance-covariance matrix. The is given at the iteration by
where contains the observed RVs. The optimal parameter step is obtained by solving
leading in the case of Newton-Raphson to the algebraic system
The matrix is a positive semi-definite matrix called the hessian matrix (it is not a positive definite matrix because it might be singular).
The Levenberg-Marquardt algorithm consists in rescaling the diagonal elements of by the damping factor . The system (Eq. 27) becomes
is the damping factor, chosen to be large when approaches the singularity.
C.2 The separable variable technique
The parameters are linear because they satisfy the property
Hereafter, subscripts and refer to linear and non-linear respectively. is the Jacobian matrix whose columns are the derivatives of with respect to .
While in both Newton-Raphson and the (commonly used) Levenberg-Marquardt (LM) algorithms, linear and non-linear parameters are treated the same way, the separable variable technique has the advantage that can be obtained for every non-linear parameter by a linear least-square regression, such that 191919The inversion of the normal equations is known to be numerically unstable and should be avoided in favour of orthogonal decomposition approaches.
with
Therefore, Eq. 22 becomes a function of only
where the step of the non-linear parameters
is .
The Jacobian matrix encompasses both the derivatives of
with respect to and
is a 33 matrix containing the derivatives of the linear parameters with respect to the non-linear ones . It is obtained by computing the derivative of Eq. 30 with respect to
Using Eq. 31 and Eq. 32, we deduce the derivatives
Therefore, the system (Eq. 28) becomes
with being the new hessian matrix of the system. We have then split the 66 system (Eq. 28) into two smaller systems, each one of 33, in Eq. 30 and Eq. 39 respectively.
C.3 Analytical expressions of the involved matrices
To perform the above computations, we need the expressions of the Jacobian matrices and . The former, relative to the linear system of Eq. 29 is given by
Its derivatives with respect to the non-linear parameters are
Very similarly, the Jacobian matrix is given by
The derivatives of the true anomalies with respect to the three non-linear parameters are
C.4 Computation steps
The parameter step in Eq. 39 is computed as follows:
- 1.
- 2.
- 3.
- 4.
- 5.
- 6.
- 7.
Appendix D Pie charts illustrating the tallies
Appendix E A few examples of resulting orbital solutions
Appendix F Comparison of orbital parameters with respect to literature values
Figures illustrating the deviations in
(expressed in units) with respect to external
catalogues are shown in Sect. 9.4
as a function of the reference period. Here we extend such
comparisons to other orbital parameters (, , and )
and show the dependence with other relevant quantities
(including the from our catalogue, not the reference one).
Appendix G Processing steps and outcome of analysis for SB9 systems
As for Griffin’s catalogue, the small number of SB9 systems eventually retained for validation (Sect. 9.4.2 and Figs. 64- 67) arises from the use of only a subset of the reference data set, along with the numerous selection criteria applied to the DR3 data described in Tables 2, 3, 4 and 18. Figure 72 provides a summary of the various cuts that were applied, along with the number of SB9 sources remaining at each stage and the final outcome of the analysis of spectroscopic binaries by CU4/NSS.
Step 1: selection applied to the cross-matches with SB9
We start off with the 3323 cross-matches out of the 4021 binaries in the SB9. Among them, close to 30% are visual/multiple systems, and are rejected straight away from the input catalogue (see beginning of Sect. 9.4.1). Furthermore, a few objects have duplicated entries in the SB9. -
Step 2: selection based on SB types in SB9
We did not consider further the SB2 solutions reported in SB9, which at this stage constitute 25% of the reference sample. As for the previous step, of course, rejection for validation purposes of a given system does not preclude it entering the DR3 subcatalogue. Among the 1750 SB1 objects from SB9, 1476 (=1750-226-45-3) do not have spectroscopic deterministic solutions (Sect. 9.4.2). -
Step 3: CU6 data not feeding the present pipeline
For more than half of the remaining systems, the CU6 data are either not transmitted to our pipeline for analysis or will not lead to any solutions. As discussed in Sects. 2 and 6.3.1, it is mainly because there is either no compelling evidence for variability in the RV time-series, there are not enough valid RV measurements, or the source lies outside the or ranges. We note that 20% of the entries in the whole SB9 contain an OBA component. -
Step 4: internal selection within our filtering scheme
Finally, only a relatively small fraction of the remaining sources have either a proposed SB1, SB1C, or TrendSB1 solution as a result of the numerous quality cuts described in Sect. 7.1 (see also Tables 2, 3, 4 and 18). As an example, Fig. 73 gives full details about the internal filtering applied to the TrendSB1 and SB1 solution types. In addition, we recall that the cross-matches with the combined and StochasticSB1 DR3 solutions are not counted in Sect. 9.4.2. The same holds for the purely astrometric orbital solutions or those for EBs only based on photometry.
Appendix H Parameters delivered to the catalogue for classes SB1, SB1C and TrendSB1.
Table 15 lists the parameters computed in the present work concerning the SB1 class. Table 16 deals with the SB1C class and Table 17 with the TrendSB1 one. Table 18 summarises the selection criteria applied on the way to the SB-subcatalogue.
| Parameter | Symbol | Table entry | Type | Units | Remarks |
| Solution type | SB1 | nss_solution_id | string | - | |
| Orbital period | period | double | d | ||
| Uncertainty | period_error | float | d | ||
| Systemic velocity | center_of_mass_velocity | double | km s-1 | ||
| Uncertainty | center_of_mass_velocity_error | float | km s-1 | ||
| Semi-amplitude | semi_amplitude_primary | double | km s-1 | ||
| Uncertainty | semi_amplitude_primary_error | float | km s-1 | ||
| Eccentricity | eccentricity | double | none | ||
| Uncertainty | eccentricity_error | float | none | ||
| Argument of periastron | periastron_argument | double | degree | radian (this work) | |
| Uncertainty | periastron_argument_error | float | degree | ||
| Periastron epoch | t_periastron | double | Barycentric JD | ||
| Uncertainty | t_periastron_error | float | Barycentric JD | ||
| Total number of points | rv_n_obs_primary | integer | none | ||
| Total number of valid points | rv_n_good_obs_primary | integer | none | ||
| Objective function | obj_function | double | none | ||
| Goodness of fit | goodness_of_fit | float | none | ||
| Efficiency | - | efficiency | float | none | |
| Significance | significance | float | none | ||
| Flags | - | flags | long | - | see appendix I |
| Confidence on period | - | conf_spectro_period | float | - | |
| Concerned parameters | - | bit_index | long | - | |
| Vector of correlations | - | corr_vec (length 15) | float | - |
| Parameter | Symbol | Table entry | Type | Units | Remarks |
| Solution type | SB1C | nss_solution_id | string | - | |
| Orbital period | period | double | d | ||
| Uncertainty | period_error | float | d | ||
| Systemic velocity | center_of_mass_velocity | double | km s-1 | ||
| Uncertainty | center_of_mass_velocity_error | float | km s-1 | ||
| Semi-amplitude | semi_amplitude_primary | double | km s-1 | ||
| Uncertainty | semi_amplitude_primary_error | float | km s-1 | ||
| Maximum velocity epoch | t_periastron | double | Barycentric JD | entry misleading | |
| Uncertainty | t_periastron_error | float | Barycentric JD | entry misleading | |
| Total number of points | rv_n_obs_primary | integer | none | ||
| Total number of valid points | rv_n_good_obs_primary | integer | none | ||
| Objective function | obj_function | double | none | ||
| Goodness of fit | goodness_of_fit | float | none | ||
| Efficiency | - | efficiency | float | none | |
| Significance | significance | float | none | ||
| Flags | - | flags | long | - | see appendix I |
| Confidence on period | - | conf_spectro_period | float | - | |
| Concerned parameters | - | bit_index | long | - | |
| Vector of correlations | - | corr_vec (length 6) | float | - |
| Parameter | Symbol | Table entry | Type | Units | Remarks |
| Solution type | First DegreeTrendSB1 | nss_solution_id | string | - | |
| SecondDegreeTrendSB1 | |||||
| Mean velocity | mean_velocity | double | km s-1 | ||
| Uncertainty | mean_velocity_error | float | km s-1 | ||
| First order derivative | first_deriv_velocity | double | km s-1 d-1 | ||
| Uncertainty | first_deriv_velocity_error | float | km s-1 d-1 | ||
| Second order derivative | second_deriv_velocity | double | km s-1 d-2 | ||
| Uncertainty | second_deriv_velocity_error | float | km s-1 d-2 | ||
| Total number of points | rv_n_obs_primary | integer | none | ||
| Total nbr of valid points | rv_n_good_obs_primary | integer | none | ||
| Objective function | obj_function | float | none | ||
| Goodness of fit | goodness_of_fit | float | none | ||
| Flags | - | flags | long | - | appendix I |
| Concerned parameters | - | bit_index | long | - | or 15 |
| Vector of correlations | - | corr_vec (length 1 or 3) | float | - |
| Stage of application | Selection criterion | Description in |
| Transfer from CU6 products | RV time-series available | Katz et al. (2023) |
| Normal stars, no spectral peculiarities | Katz et al. (2023) | |
| Object not classified as SB2 | Damerdji et al. (2024) | |
| Selection of | rv_renormalized_gof larger than 4 | Sect. 2 |
| objects/time-series at entry | in [3875 - 8125] K | Sect. 2 |
| mag in [5.5 - 12] | Sect. 2 | |
| rv_chisq_pvalue less than or equal to 0.01 | Sect. 2 | |
| RV outlier rejected from time-series | Sect. 6.3.1 | |
| larger than or equal to 10 | Sect. 6.3.1 | |
| Internal filtering on solutions | 1000 km s-1 | Sect. 7.1 |
| less than or equal to 3 | Sect. 7.1 | |
| Rejection of solutions with bad flags | Sect. 7.1 | |
| less than or equal to 250 km s-1 | Sect. 7.1 | |
| Efficiency greater than or equal to 0.1 | Sect. 7.1 | |
| less than or equal to 0.3 | Sect. 7.1 | |
| conf_spectro_period | Sect. 7.1 | |
| Significance larger than or equal to 5 | Sect. 7.1 | |
| less than 2 (solely for SB1) | Sect. 7.1 | |
| Loss by combination | — | Sect. 7.2 |
| Post-filtering | 380 objects rejected as intrinsic variables (Cepheids and RR Lyrae) | Sect. 10.1.1 |
| 1475 objects rejected having more than 10 % double-line transits | Sect. 10.1.2 | |
| 164 objects rejected suffering from a scan-angle effect | Sect. 10.1.3 |
Appendix I Various flags linked to the processing
Table 19 details the various flags that could be activated during the spectroscopic orbital solution processing, along with the relevant explanations in the framework of the decision tree.
| Flag bit nr. | Value | Logical Variable |
| 8 | 256 | BAD_UNCHECKED_NUMBER_OF_TRANSITS |
| Explanation: The number of transits is not sufficient to process the star (before removing bad transits) | ||
| 9 | 512 | NO_MORE_VARIABLE_AFTER_FILTERING |
| Explanation: The RV curve is no more variable after velocity filtering (at the threshold 0.9) | ||
| 10 | 1024 | BAD_CHECKED_NUMBER_OF_TRANSITS |
| Explanation: The number of transits is not sufficient to process the star (after removing bad transits) | ||
| 11 | 2048 | SB2_REDIRECTED_TO_SB1_CHAIN_NOT_ENOUGH_COUPLE_MEASURES |
| Explanation: SB2 is redirected to the SB1 chain because there are not enough couples of measures | ||
| 12 | 4096 | SB2_REDIRECTED_TO_SB1_CHAIN_PERIODS_NOT_COHERENT |
| Explanation: SB2 is redirected to the SB1 chain because the periods found by the SB1 and SB2 chains are not coherent | ||
| 13 | 8192 | NO_SIGNIFICANT_PERIOD_CAN_BE_FOUND |
| Explanation: No significant period can be derived (no period from photometric variability analysis and periodogram peaks below the cut-off threshold) | ||
| 14 | 16384 | REFINED_SOLUTION_DOES_NOT_CONVERGE |
| Explanation: The refined orbital solution does not converge (after 1000 iterations) | ||
| 15 | 32768 | REFINED_SOLUTION_SINGULAR_VARIANCE_COVARIANCE_MATRIX |
| Explanation: The variance-covariance matrix can not be obtained (singular) for the refined solution | ||
| 16 | 65536 | CIRCULAR_SOLUTION_SINGULAR_VARIANCE_COVARIANCE_MATRIX |
| Explanation: The variance-covariance matrix can not be obtained (singular) for the circular solution | ||
| 17 | 131072 | TREND_SOLUTION_SINGULAR_VARIANCE_COVARIANCE_MATRIX |
| Explanation: The variance-covariance matrix can not be obtained (singular) for the trend solution | ||
| 18 | 262144 | REFINED_SOLUTION_NEGATIVE_DIAGONAL_OF_VARIANCE_COVARIANCE_MATRIX |
| Explanation: The diagonal of the variance-covariance matrix is negative for the refined solution | ||
| 19 | 524288 | CIRCULAR_SOLUTION_NEGATIVE_DIAGONAL_OF_VARIANCE_COVARIANCE_MATRIX |
| Explanation: The diagonal of the variance-covariance matrix is negative for the circular solution | ||
| 20 | 1048576 | TREND_SOLUTION_NEGATIVE_DIAGONAL_OF_VARIANCE_COVARIANCE_MATRIX |
| Explanation: The diagonal of the variance-covariance matrix is negative for the trend solution | ||
| 21 | 2097152 | CIRCULAR_SOLUTION_DOES_NOT_CONVERGE |
| Explanation: The Lucy refined orbital solution diverges (after 1000 iterations) | ||
| 22 | 4194304 | LUCY_TEST_APPLIED |
| Explanation: The Lucy test has been applied | ||
| 23 | 8388608 | TREND_SOLUTION_NOT_APPLIED |
| Explanation: The trend analysis has not been applied (case of unsorted SB2) | ||
| 24 | 16777216 | SOLUTION_OUTSIDE_E_LOGP_ENVELOP |
| Explanation: The orbital solution is above the - envelop | ||
| 25 | 33554442 | PERIOD_FOUND_IN_CU7_PERIODICITY |
| Explanation: The period is equal to a period issued from the CU7 (within the quadratic sum of their uncertainties) | ||
| 26 | 67108864 | FORTUITOUS_SB2 |
| Explanation: The RV1 and RV2 seem to be uncorrelated (contrary to the case of any binary system) | ||
Appendix J Examples of retrieving a gold sample
For an SB1 gold sample, we suggest limiting the efficiency to values larger than or equal to 0.2. To secure the determination of the periods and the subsequent fits, we advise adopting larger or equal to 20. The most important selection is related to the significance that must remain larger or equal to 40. This condition eliminates dubious solutions, notably with small periods and large eccentricity. Solutions with periods larger than 800 d will not be secure, and are known over 1000 d to present biases and a small efficiency. Periods between 2 and 10 d can be kept thanks to the choice on the significance. An alternative possibility would be to relax the condition on the significance to 25-30 and to discard periods below 10 d. Below 2 d, we advise being cautious, although some positive validation took place. The following ADQL command will deliver some 15314 objects:
SELECT * FROM gaiadr3.nss_two_body_orbit
WHERE nss_solution_type = ’SB1’
AND efficiency >= 0.2
AND rv_n_good_obs_primary >= 20
AND significance >=40
AND period <= 800
AND period >= 2 .The recovery rate of the period for this gold sample reaches 93 %.
Since objects characterised by a ruwe above 1.4 have thus a dubious position and consequently possible problems with the RVs, we recommend an additional selection on the astrometric ruwe. This selection implies the use of two tables and restricts the sample to 9578 objects:
SELECT * FROM gaiadr3.nss_two_body_orbit
INNER JOIN gaiadr3.gaia_source_lite USING (source_id)
WHERE nss_solution_type = ’SB1’
AND efficiency >= 0.2
AND rv_n_good_obs_primary >= 20
AND significance >=40
AND period <= 800
AND period >= 2
AND ruwe < 1.4 .
Finally, we also present a gold selection on the linear trends that delivers 12650 objects:
SELECT * FROM gaiadr3.nss_non_linear_spectro
WHERE nss_solution_type = ’FirstDegreeTrendSB1’
AND rv_n_obs_good_primary >= 20 .

### v2

@@
 Context.
-The Gaia satellite constitutes one of the cornerstone missions of ESA.
-This is an astrometric space experiment that is measuring positions, proper motions as well as parallaxes for a huge number of stars.
-It is also performing photometric and spectrophotometric observations.
-Gaia also operates a medium-dispersion spectrometer, the RVS, that provides spectra and thus radial velocity (RV) time-series.
+The Gaia satellite constitutes one of ESA’s cornerstone missions.
+Being primarily an astrometric space experiment measuring positions, proper motions, and parallaxes for a huge number of stars, it also performs photometric and spectrophotometric observations.
+Gaia operates a medium-dispersion spectrometer, known as Radial Velocity Spectrometer (RVS), which provides spectra and radial velocity (RV) time series.
 Aims.
-The paper is centred on the analysis of the RV time-series.
-We try to fit orbital models and trend models and restrict ourselves to the objects of spectral types F-G-K brighter than magnitude 12 presenting only one single spectrum (SB1).
+The paper is focussed on the analysis of the RV time series.
+We fit orbital and trend models, restricting our study to objects of spectral types F-G-K that are brighter than a magnitude of 12, presenting only one single spectrum (SB1).
 Methods.
-Suitable time-series are processed and analysed object per object, providing orbital or trend solutions.
-The results of the various fits are further filtered internally on the basis of several quality measures in order to discard spurious solutions.
-The objects only having a spectroscopic solution are classified in one of the three classes SB1 (eccentric model), SB1C (circular model) or TrendSB1 (mere trend model).
+Suitable time series were processed and analysed on an object-per-object basis, providing orbital or trend solutions.
+The results of the various fits were further filtered internally on the basis of several quality measures to discard spurious solutions.
+The objects with solely a spectroscopic solution were classified in one of the three classes: SB1 (eccentric model), SB1C (circular model), or TrendSB1 (mere trend model).
 Results.
-We detail the methods used and describe the derived parameters and results.
-After a description of the models considered, and of the related quality tests of the fit, we detail the internal filtering that is intended to reject bad solutions.
+We detail the methods used in this work and describe the derived parameters and results.
+After a description of the models considered and the related quality tests of the fit, we detail the internal filtering process aimed at rejecting bad solutions.
 We also present a full validation of the pipeline.
-A description of the current content of the catalogue is provided.
+A description of the current content of the catalogue is also provided.
 Conclusions.
-We present the SB1, SB1C and the TrendSB1 spectroscopic solutions contained in the SB-subcatalogue which is part of the DR3 catalogue.
-We delivered some 181327 orbital solutions in the class SB1, 202 in the class SB1C and 56808 in the associated class TrendSB1.
-This is a first release and the delivered SB-subcatalogue could be tuned and refined.
-However the majority of the entries are correct; this data set constitutes by far the largest set of spectroscopic orbital solutions ever constituted.
+We present the SB1, SB1C, and TrendSB1 spectroscopic solutions contained in the SB subcatalogue, part of the DR3 catalogue.
+We deliver some 181 327 orbital solutions in class SB1, 202 in class SB1C, and 56 808 in the associated class TrendSB1.
+This is a first release and the delivered SB subcatalogue could be further tuned and refined.
+However, the majority of the entries are correct.
+Thus, this data set constitutes by far the largest set of spectroscopic orbital solutions to be computed.
 Key Words.:
 Techniques: spectroscopy; Stars: Techniques: radial velocities; Stars: binaries; Catalogues: Survey; Gaia1 Introduction
-The Gaia satellite constitutes one of the cornerstone missions of the European Space Agency.
-This is an astrometric space experiment intended to represent a new breakthrough compared to the previous pioneering one named Hipparcos (Perryman 2009).
-The Gaia satellite has two telescopes, looking at 1065 of each other, and three associated scientific instruments, continuously scanning the celestial sphere thanks to its six-hour spinning on itself (around an axis perpendicular to the two telescope lines of sight), to the precession of the spinning axis around the direction of a fictitious, nominal Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2) with a 63-day period and to the one-year orbit of the satellite around the Sun accompanying the Earth in its revolution (Gaia Collaboration et al.
+The Gaia satellite constitutes one of the cornerstone missions of the European Space Agency (ESA).
+This astrometric space experiment is intended to represent a new breakthrough, compared to the previous pioneering one, Hipparcos (Perryman 2009).
+The Gaia satellite has two telescopes, observing at 1065 of each other, and three associated scientific instruments that are continuously scanning the celestial sphere thanks to its six-hour rotation (around an axis perpendicular to the two telescope’s lines of sight), as well as the precession of the spinning axis around the direction of a fictitious, nominal Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2), with a 63-day period, and thanks to the one-year orbit of the satellite around the Sun, while it accompanies the Earth in its revolution (Gaia Collaboration et al.
 2016b).
-Multiple visits are necessary to reach a convenient S/N ratio, but also to gather information on binarity (or higher order multiplicity) and variability.
+Multiple visits are necessary to reach a convenient signal-to-noise (S/N) ratio, but also to gather information on binarity (or higher order multiplicity) and variability.
 The Gaia satellite is measuring the positions and transverse proper motions for some 1.8 billion stars, as well as precise parallaxes and thus distances for a very large number of objects.
 The detectors are working in a Time Delay Integration mode (TDI).
-The satellite was launched on December 19th, 2013 and is scanning the sky since the end of the commissioning period in July 2014.
+The satellite was launched on 19 December 2013, and has been scanning the sky since the end of the commissioning period in July 2014.
 The detailed description of the project and of the payload can be found in Gaia Collaboration et al.
-(2016b) that was published along with the first Data Release (DR1).
+(2016b), published alongside the first Data Release (DR1).
 The description of DR1 and of the related data products is provided in Gaia Collaboration et al.
-(2016a) whereas the DR2 was detailed in Gaia Collaboration et al.
+(2016a) whereas DR2 was detailed in Gaia Collaboration et al.
 (2018a).
@@
 2023d).
-Besides astrometry performed with the Astrometric Field (AF) instrument (Lindegren et al.
-2018, 2021), the satellite also produces photometry from the same instrument as well as spectrophotometry provided by another channel: the BP/RP spectrophotometer purveyor of very low resolution spectra as well as of photometric colours (Riello et al.
+In addition to astrometry performed with the Astrometric Field (AF) instrument (Lindegren et al.
+2018, 2021), the satellite also produces photometry from the same instrument as well as spectrophotometry provided by another channel: the BP/RP spectrophotometer purveyor of very low-resolution spectra as well as of photometric colours (Riello et al.
 2018, 2021; Evans et al.
 2018).
-The two above-mentioned channels are supposed to observe stellar objects down to about magnitude 20th in band.
-Besides astrometry and spectrophotometry, the satellite also incorporates a third channel: a medium resolution spectrometer to measure the radial velocities (RVs) of stellar objects, named the Radial Velocity Spectrometer (RVS).
-This channel and the related instrument are described in great details in Cropper et al.
-(2018) which also includes a nice review of the genesis of the RVS instrument.
+The two above-mentioned channels are supposed to observe stellar objects down to a magnitude of about 20 in the band.
+In addition to astrometry and spectrophotometry, the satellite also incorporates a third channel: a medium-resolution spectrometer to measure the radial velocities (RVs) of stellar objects, known as the Radial Velocity Spectrometer (RVS).
+This channel and the related instrument are described in great detail in Cropper et al.
+(2018), which also includes a comprehensive review of the genesis of the RVS instrument.
 The RVS brings the third component of the space velocity (to join with the transverse proper motion on the sky).
-It also provides for the brightest stars information on the nature of the object (, , chemical abundances, chromospheric activity, and so on).
-Stars brighter than magnitude = 2.76 produce saturated pixels and are not considered (for the definition of the magnitudes, see Sartoretti et al.
+It also provides, for the brightest stars, information on the nature of the object (, , chemical abundances, chromospheric activity, etc.).
+Stars brighter than a magnitude of = 2.76 produce saturated pixels and, thus, are not considered here (for the definition of the magnitudes, see Sartoretti et al.
 2023).
 The scanning law for RVS is intended to typically produce 40 transits over the nominal five-year mission.
-The focal plane of the RVS is covered by 12 CCDs (4 rows of 3 in a row).
+The focal plane of the RVS is covered by 12 CCDs (four rows of three in a row).
 The RVS instrument acquires, in TDI mode, spectra in the range 845-872 nm, including the well-known Ca ii triplet (in late-type stars).
-With a resolving power of about (see section 8.2 and 11 of Cropper et al.
+With a resolving power of about (see Sects.
+8.2 and 11 of Cropper et al.
 2018), it obtains three spectra per transit, each corresponding to a maximum exposure time of 4.42 s (they are analysed individually and concomitantly).
@@
 (2019).
-The stars exhibiting a double-line spectrum (also named composite spectrum) were rejected from the process (as described in Sartoretti et al.
+The stars exhibiting a double-line spectrum (also known as the composite spectrum) were rejected from the process (as described in Sartoretti et al.
 2018; Damerdji et al.
@@
 The first modules are working per transit and per trending epochs.
-The flowchart is illustrated in section 2.3 of Sartoretti et al.
+The flowchart is illustrated in Sect.
+2.3 of Sartoretti et al.
 (2018).
-The Single Transit Analysis (STA) is the workflow that actually performs the measurements of the RV per transit.
+The single transit analysis (STA) refers to the workflow that actually performs the measurements of the RV per transit.
 Three submodules are measuring the RVs of all the selected convenient stars (single spectrum ones).
@@
 2018).
-The interest to make use of several methods is explained in the early investigation of the problem, whose conclusions are detailed in David et al.
-(2014) (see also section 7 of Sartoretti et al.
+The incentive in making use of several methods is explained in the early investigation of the problem, with conclusions detailed in David et al.
+(2014) (see also Sect.
+7 of Sartoretti et al.
 2018).
@@
 Thereafter, the multiple transit analysis (MTA) workflow performs some statistical analysis tests based on the RVs from all objects at all transits.
-It computes per star a probability of constancy that is used to spot out variable objects.
+It computes (per star) a probability of constancy that is used to spot out variable objects.
 It also computes the median of the RVs over all transits for all the bright objects (Sartoretti et al.
@@
 2023; Holl et al.
-2023b), the photometric data (Siopis et al.
+2023b), photometric data (Siopis et al.
 2024; Pourbaix et al.
-2022, see section 7.6 by Siopis & Sadowski) and the spectroscopic data (the present paper, and Damerdji et al.
+2022, see Sect.
+7.6 by Siopis & Sadowski), and the spectroscopic data (the present paper, and Damerdji et al.
 2024).
 The detailed description of the CU4 spectroscopic processing chain for SB1 (spectroscopic binaries with only one spectrum visible; hereafter labelled NSS-SB1) orbital solutions or trend ones is the subject of the present paper whereas Damerdji et al.
-(2024) describes the SB2 channel.
-These DR3 products are preliminary results; this is the first time that orbital solutions issued from Gaia are delivered.
+(2024) describe the SB2 channel.
+These DR3 products are preliminary results; this is the first time that orbital solutions issued from Gaia have been delivered.
 The results of the three channels can be further processed by combining their independent outputs.
@@
 2023a; Pourbaix et al.
-2022, see section 7.7 by Gavras & Arenou).
-The scientific conception of the SB1 chain described here was elaborated by Yassine Damerdji, Eric Gosset and Thierry Morel; most of the codes were benchmarked in MATLAB and finally written in Java by YD with some help from TM; the validation tests (including the critical building of the ground-based data catalogues used for comparison) were mainly performed by TM.
-The whole work was made in the highly favourable and boosting context of the NSS team.
-Section 2 describes the nature of the data that are entering the process, whereas Section 3 describes the simple mathematical formalism used and lists the models that are considered.
+2022, see Sect.
+7.7 by Gavras & Arenou).
+The scientific conception of the SB1 chain described here was elaborated on by Yassine Damerdji, Eric Gosset and Thierry Morel; most of the codes were benchmarked in MATLAB and finally written in Java by YD with some help from TM; the validation tests (including the critical building of the ground-based data catalogues used for comparison) were mainly performed by TM.
+The whole work was carried out in the highly favourable and boosting context of the NSS team.
+Section 2 describes the nature of the data that are entering the process, whereas Sect.
+3 describes the simple mathematical formalism used and lists the models that are considered.
 Section 4 explicits the importance of the period search and the related difficulties.
-Sect.
-5 briefly summarises the main trials we did to finally arrive at the method described in Sect.
-6 that presents the chain as adopted for the DR3 (NSS-SB1 main processing) whereas Sect.
-7 introduces the notion of post-processing.
+Section 5 briefly summarises the main trials we did to arrive at the method described in Sect.
+6, where we present the chain as adopted for the DR3 (NSS-SB1 main processing).
+In Sect.
+7, we introduce the notion of post-processing.
 Section 8 illustrates the results.
@@
 10 explain the details of the post-filtering and provide additional information in the light of the conclusions from the previous section.
-The catalogue description, the caveats and the precautions necessary in using the results of the NSS-SB1 processing are given in Section 11.
-Section 12 concludes the paper.
-The content of the present paper is more detailed and contains material that supplants the information given in the Gaia archives (see sections 7.4 and 7.5 in Pourbaix et al.
-2022)111https://gea.esac.esa.int/archive/documentation/GDR3/ at the epoch of the release.
+The catalogue description, the caveats and the precautions necessary in using the results of the NSS-SB1 processing are given in Sect.
+11.
+Section 12 presents our conclusions.
+The contents of the present paper are more detailed and contain material that supplants the information given in the Gaia archives (see Sects.
+7.4 and 7.5 in Pourbaix et al.
+2022)111https://gea.esac.esa.int/archive/documentation/GDR3/ at the time of the release.
+The main core of the paper is accompanied by appendices.
+Appendices D to J are only available on Zenodo (see https://zenodo.org/records/13990211).
 2 Description of the input data
@@
 The RVs are measured individually on the three spectra.
-Three methods are used to derive the radial velocities (Cropper et al.
+Three methods are used to derive the radial velocities (Sartoretti et al.
 2018), each of them providing three velocities (one per CCD) and a combined RV value (over the three CCDs).
@@
 2023).
-Finally, if none of these sources provides the necessary parameters, the latter are estimated as previously done for DR2 (DetermineAP: section 6.5 of Sartoretti et al.
+Finally, if none of these sources provides the necessary parameters, the latter were estimated as previously done for DR2 (DetermineAP: Sect.
+6.5 of Sartoretti et al.
 2018).
@@
 The final adopted template is the one that is the major contributor over all the transits.
-More details can be found in section 3.6.1 of Katz et al.
+More details can be found in Sect.
+3.6.1 of Katz et al.
 (2023).
@@
 Consequently, a template mismatch is expected to essentially introduce a systematic RV shift (for single-line objects).
-The measurements of the RVs are performed under the responsibility of the STA Development Unit.
-The list of good RVs corresponding to a particular star are then analysed by the MTA in order to detect and separate variable RV sets from constant RV sets.
+The measurements of the RVs were performed under the supervision of the STA Development Unit.
+The list of good RVs corresponding to a particular star are then analysed by the MTA to detect and separate variable RV sets from constant RV sets.
 For each star, the median of the RVs is computed and appears in the main catalogue.
-The stars with variable RVs are further considered by the global pipeline and are forwarded to the CU4 (Object processing: NSS).
+The stars with variable RVs were further considered by the global pipeline and forwarded to the CU4 (Object processing: NSS).
 For DR3, the adopted cutting threshold corresponds to a probability of variability of 0.99 (rv_chisq_pvalue .
-The composite spectra and SB2 objects are separately treated already at the level of the CU6 processing (Damerdji et al.
+The composite spectra and SB2 objects were already treated separately at the level of the CU6 processing (Damerdji et al.
 2024) for the determination of both RVs and are then dispatched to CU4/NSS in order to search for the corresponding orbital solutions (Damerdji et al.
@@
 The non-single star (NSS) processing is intended to perform the reduction of the corresponding data from the point of view of multiple star studies.
-In the present paper, we will consider the sole spectroscopic channel for the SB1 objects.
-The input data are made of a list of RVs (and their 1-uncertainties) that are corrected to make reference to the barycentre of the solar system.
-It should be pointed out that for DR3 the rms error on the barycentric correction ( km s-1) is not propagated to the uncertainty on the individual RVs.
+In the present paper, we consider the sole spectroscopic channel for the SB1 objects.
+The input data were drawn from a list of RVs (and their 1-uncertainties) that were corrected to make reference to the barycentre of the solar system.
+It should be pointed out that for DR3 the rms error on the barycentric correction ( km s-1) was not propagated to the uncertainty on the individual RVs.
 The RVs are accompanied by a time of observations expressed in barycentric Julian days (BJD).
-The number of points of the time-series corresponds to the number of transits for which the RV has been properly measured with success.
-The data could present several weaknesses such that outliers and wrong RVs due e.g.
-to template mismatch errors (for the latter case, most probably the same systematic for all values within a time-series).
+The number of points of the time series corresponds to the number of transits for which the RV has been properly measured with success.
+The data could present several weaknesses such that outliers and wrong RVs due, for instance, to template mismatch errors (for the latter case, most likely the same systematic for all values within a time series).
 These problems and additional selection rules are further documented in Katz et al.
@@
 The threshold has been chosen at mas, equivalent to about 29 km s-1.
-Other filtering applied to the input CU6 data is described in section 4 of Katz et al.
+Other filtering applied to the input CU6 data is described in Sect.
+4 of Katz et al.
 (2023).
@@
 2023).
-The correction has not been applied for the catalogue, nor for the transmission of the RVs to the NSS processing.
-The effect is quite small and should only induce per object a global shift of all the RVs with as a consequence a small bias on the systemic velocities (see Sect.
+The correction was not applied for the catalogue, nor for the transmission of the RVs to the NSS processing.
+The effect is quite small and should only induce a global shift of all the RVs per object along with, as a consequence, a small bias on the systemic velocities (see Sect.
 3).
 The current astrometric solution for DR3 has been released as part of the EDR3 and contains, among others, positions for a large number of objects.
-The astrometric solution (AGIS) is based on a treatment assuming the objects are single, unresolved stars and thus having an image characterised by the instrumental line-spread function (for the one dimensional observations in the astrometric channel) or point-spread function (for the two-dimensional observations), see the description in Lindegren et al.
+The astrometric solution (AGIS) is based on a treatment assuming the objects are single, unresolved stars and thus having an image characterised by the instrumental line-spread function (for the 1D observations in the astrometric channel) or point-spread function (for the 2D observations), as described in Lindegren et al.
 (2021).
@@
 Lindegren et al.
-(2021) also associate to these objects an excess noise (impacting the position) that is delivered in the catalogue under the name astrometric_excess_noise and which corresponds to the additional dispersion necessary to accept the concerned object in the global solution.
+(2021) also associated an excess noise (impacting the position) to these objects, which is delivered in the catalogue under the name astrometric_excess_noise and which corresponds to the additional dispersion necessary to accept the concerned object in the global solution.
 Objects with a ruwe larger than 1.4 are certainly problematic, but smaller values already have an impact on the RVS calibration.
 Since the RVS is a slitless spectrometer, the calibration in wavelength of the spectra is requesting an epoch position of the object that is accurate enough.
-Therefore, for the objects that are deviating from the single star behaviour, we could expect that the wavelength calibration is not correct, and thus the measured transit RVs are somewhat biased and this creates artefacts in the time-series.
+Therefore, for the objects that are deviating from the single star behaviour, we could expect that the wavelength calibration is not correct, and thus the measured transit RVs are somewhat biased and this creates artefacts in the time series.
 No correction of this effect is offered for DR3.
@@
 This term is derived as being (km s-1) = 0.146 astrometric_excess_noise (mas) where the factor 0.146 is related to the dispersion of the spectrometer (transformation of mas into km s-1).
-Typically, the values of astrometric_excess_noise are below 10 mas, and thus is below 1.46 km s-1.
-Some selections are also made on the objects at the moment of their entry in the NSS-SB1 processing.
-In the case of DR3, the kind of stars is restricted to the range of effective temperatures 3875 K to 8125 K (mainly ((M))-K-G-F-(A) stars).
-No emission-line object has been selected, and in principle only stars without spectral peculiarities are entering the game.
-The objects with a rv_renormalised_gof larger than 4 were selected (Katz et al.
+Typically, the values of astrometric_excess_noise are below 10 mas and, thus, is below 1.46 km s-1.
+Some selections were also made on the objects at the moment of their entry in the NSS-SB1 processing.
+In the case of DR3, these kinds of stars are restricted to the range of effective temperatures 3875 K to 8125 K (mainly ((M))-K-G-F-(A) stars).
+No emission-line object was selected and, in principle, only stars without spectral peculiarities were permitted.
+The objects with a rv_renormalised_gof value greater than 4 were selected (Katz et al.
 2023) to enter the NSS-SB1 processing because they were considered as clearly non-constant.
 The measurements of the individual (per transit) RVs are restricted to the range –1000 to +1000 km s-1.
-A global view of the selections applied beforehand is available in Table 18 (Appendix H).
+A global view of the selections applied beforehand is available in Table 18.
 This measurement of the individual RVs on individual spectra is only possible for bright objects.
@@
 9.2.1.
-A threshold on has been selected at 12 mag.
-Fainter objects are considered to produce individual spectra for which the S/N ratio forbids to derive a sound radial velocity.
+A threshold on was selected at 12 mag.
+Fainter objects are considered to produce individual spectra for which the S/N ratio does not enable the derivation of a sound radial velocity.
 This strongly reduces the amount of selected objects.
@@
 When the SB2 chain (Damerdji et al.
-2024) analyses an SB2 time-series and is not able to derive the corresponding SB2 orbital solution, the object can be redirected to the SB1 chain.
-An SB1 time-series is built with the RVs characterised by the smallest uncertainties.
+2024) analyses an SB2 time series and is not able to derive the corresponding SB2 orbital solution, the object can be redirected to the SB1 chain.
+An SB1 time series is built with the RVs characterised by the smallest uncertainties.
 The object is then considered as an SB1 and is forwarded to the present chain.
@@
 3 The underlying orbital models
-Four kinds of orbital solutions are considered here: they are labelled SB1, SB1C, TrendSB1 and StochasticSB1.
+Four kinds of orbital solutions are considered here: they are labelled SB1, SB1C, TrendSB1, and StochasticSB1.
 They are described here below.
@@
 The uncertainties on the parameters are calculated in the classical way on the basis of the diagonal elements of the inverse of the curvature matrix (i.e.
-the variance-covariance matrix, see Bevington & Robinson 2003).
+the variance-covariance matrix; see Bevington & Robinson 2003).
 Despite the fact that we use the Greek symbol for all (i.e.
@@
 which in turn is expressed as a function of the mean anomaly
-where is the time of passage at periastron.
-The is the argument of periastron measured in the sense of the orbit with its origin at the ascending node.
+Here, is the time of passage at periastron, while is the argument of periastron measured in the sense of the orbit with its origin at the ascending node.
 In this model, the eccentricity enters as a strongly non-linear parameter.
@@
 Concerning the , the model can be linearised by using the following equivalent equation
-We then compute
+We then computed
 and
@@
 However, it could be slightly shorter for some objects because it combines with the gaps in the scanning law.
-The present modelisation is restricted to polynomials of degrees one to four and presenting only one extremum on the interval.
+The present modelisation is restricted to polynomials of one to four degrees and presenting only one extremum on the interval.
 The degree of the polynomial is chosen on the basis of an F-test for nested models.
@@
 The outputs of the NSS spectroscopic pipeline are delivered to the combiner that operates downstream.
-The task of the combiner is to inspect the output of the astrometric, photometric and spectroscopic NSS channels in order to detect possible solutions corresponding to the same object and that could be combined.
+The task of the combiner is to inspect the output of the astrometric, photometric, and spectroscopic NSS channels in order to detect possible solutions corresponding to the same object and that could be combined.
 This will be described in Sect.
@@
 4 Determination of the periodicity
-The first step of the analysis is the search for a periodic behaviour of the RVs.
+The first step in the analysis is the search for a periodic behaviour of the RVs.
 The analysis is based on a periodogram (actually a frequencygram), which has the nature of the Fourier power spectrum.
 Since the time sampling of the Gaia satellite is particularly complex, the classical Fourier formalism is not applicable.
-To well understand the limitation of the method, it is necessary to have some knowledge of the impact of the sampling and of a few characteristic timescales.
+To aptly understand the limitation of the method, it is necessary to have some knowledge of the impact of the sampling and of a few characteristic timescales.
 4.1 Effect of the Gaia sampling
-In order to explain the effect of the Gaia sampling, we have to start from a single, idealistic, case and progressively add some complexity.
+To explain the effect of the Gaia sampling, we have to start from a single, idealistic, case and progressively add certain degrees of complexity.
 We start by considering the spinning of the satellite on itself that makes that a star is measured every 0.25 d (this is our first basic idealistic case since this simple view is never occurring in practice).
@@
 Therefore, 0.25 d is the typical, smallest regular and periodic timescale in the sampling.
-The 0.25 d step should generate in the time-series some aliasing with a typical 2 d-1 Nyquist frequency.
+The 0.25 d step should generate in the time series some aliasing with a typical 2 d-1 Nyquist frequency.
 This means that a progenitor frequency will be duplicated at , , , and so on (in d-1 units).
@@
 It has been introduced by Deeming (1975, 1976).
-In the case of an equidistant sampling, the spectral window is the well-known Dirac comb (or Shah function, see Bloomfield 1976).
+In the case of an equidistant sampling, the spectral window is the well-known Dirac comb (or Shah function; see Bloomfield 1976).
 The general spectral window represents an extension of this Dirac comb to the general case of odd sampling.
 The property that the generalised Fourier Spectrum of the sampled data is the result of the convolution of the underlying Fourier Spectrum of the signal by the spectral window remains true for almost any kind of sampling (Deeming 1975, 1976).
-If the satellite was limited to a rotation on itself, the semi-amplitude spectral window corresponding to the combined two fields of view would have the aspect given in the upper panel of Fig.
+If the satellite was limited to a rotation on itself, the semi-amplitude spectral window corresponding to the combined two fields of view would resemble the plot in the upper panel of Fig.
 1.
@@
 2016b).
-Therefore, the sky sampling is markedly non-uniform (see section 5.2 of Gaia Collaboration et al.
+Therefore, the sky sampling is markedly non-uniform (see Sect.
+5.2 of Gaia Collaboration et al.
 2016b).
@@
 1.
-We produce typical time-series of Gaia RVS with the GOST tool222https://gaia.esac.esa.int/gost/.
+We produce typical time series of Gaia RVS with the GOST tool222https://gaia.esac.esa.int/gost/.
 Fig.
-1 second panel represents a typical spectral window of a time-series containing 26 observations, the third panel 19 and the bottom one 14 observations (against the observations for the first panel).
+1 second panel represents a typical spectral window of a time series containing 26 observations, the third panel 19 and the bottom one 14 observations (against the observations for the first panel).
 This small number of observations induces an aliasing but the aliasing typical of the Gaia sampling is not leaking power towards a very few other frequencies but towards numerous frequencies.
@@
 Typically, pseudo-aliasing can occur at any .
-If existing, the value of is most probably located between and let say 3 d-1.
+If existing, the value of is most probably located below let say 3 d-1.
 The structure due to the motion considered in the top panel of Fig.
@@
 Looking at the third panel of Fig.
-1, one can notice marked aliasing at and at .
+1, we notice a marked aliasing at and at .
 Therefore, in a few cases, noise could conspire to shift the largest peak position from the progenitor peak by this amount, thus producing pseudo-aliases of the progenitor that could dominate.
@@
 On purpose, we discarded the first month of scanning which was characterised by the ecliptic pole scanning law (EPSL).
-For the latter, the precession of the satellite was stopped and the scanning law is more relevant to the scanning due to the spinning of the satellite except for a slow rotation due to the orbital motion around the Sun.
-We thus neglect the EPSL to concentrate on the effect of the NSL.
-For each simulation, we derive the corresponding amplitude spectral window (for the RVS instrument).
+For the latter, the precession of the satellite was stopped and the scanning law clearly becomes more relevant to the scanning due to the spinning of the satellite, except for a slow rotation due to the orbital motion around the Sun.
+We thus neglected the EPSL to instead concentrate on the effect of the NSL.
+For each simulation, we derived the corresponding amplitude spectral window (for the RVS instrument).
 Figure 2 illustrates (in red) the median of the amplitude spectral windows.
 Despite the fact that we restricted our simulations to the NSL, the aliasing due to the spinning of the satellite is clearly visible.
-However, even the peaks at d-1 (and at d-1) are rather low compared to the peak at d-1, at most 30% (in amplitude).
+However, even the peaks at d-1 (and at d-1) are rather low, compared to the peak at d-1, at most 30% (in amplitude).
 We also plot in Fig.
@@
 Although a few positions on the sky could still exhibit the effect of the spinning aliasing.
-Of course, if, over the 34 months, the EPSL scanning is dominating the definition of the times in the analysed time-series, the effect will increase exhibiting a stronger spinning aliasing.
+Of course, if, over the 34 months, the EPSL scanning is dominating the definition of the times in the analysed time series, the effect will increase exhibiting a stronger spinning aliasing.
 We must insist here on the fact that no signal is analysed in the present section; the dispersion of the amplitude spectral windows is originating from the randomness of the selection of the sky positions and from the variety of nature of the aliasing as a function of the latter.
-For any time-series, the observing time could be used to pre-calculate the particular spectral window independently of the signal.
+For any time series, the observing time could be used to pre-calculate the particular spectral window independently of the signal.
 4.2 Search for the periodicity
-When a periodicity is present in a time-series, the periodogram exhibits a peak at the corresponding frequency (the progenitor), but other peaks are also present at the location of the pseudo-aliases of the progenitor.
+When a periodicity is present in a time series, the periodogram exhibits a peak at the corresponding frequency (the progenitor), but other peaks are also present at the location of the pseudo-aliases of the progenitor.
 Contrary to the case of even sampling, the actual sampling generates pseudo-aliases that are in principle lower than the progenitor peak.
 However, the combination of the aliasing with background noise or other phenomena (blending) induces the fact that the choice of the true progenitor among the aliased peaks is never fully secure.
-The total span of time of the DR3 time-series is at most of 34 months and thus the largest periodicity accessible from the point of view of the Fourier corresponds roughly to a frequency of 1/1000 d-1.
+The total span of time of the DR3 time series is at most of 34 months and thus the largest periodicity accessible from the point of view of the Fourier corresponds roughly to a frequency of 1/1000 d-1.
 This value could actually be larger taking into account the gaps generated by the actual sampling law.
-The constraints should thus be expressed as a function of the actual length of the time-series and thus as a function of .
+The constraints should thus be expressed as a function of the actual length of the time series and thus as a function of .
 It has been decided that it would be adequate to limit the range of investigation for periodicities to since beyond the possible orbit is certainly not well determined.
@@
 6.3 for more details).
-The adopted frequency step is actually where is the actual total time-span of the individual time-series.
-This represents an oversampling of a factor one hundred that gives directly access to the amplitude of the signal as being the height of the peak.
-From the previous section discussion, we decided that the adopted largest frequency to be explored is 4 d-1.
-This choice is rather atypical since it is chosen to correspond to the spinning period.
-If the star indeed varies in a periodic way, the largest peak in the periodogram is an estimate of the possible period but definitively not a proof of its existence.
-One should consider that a periodicity is proven if a minimum of three identical cycles are observed.
+The adopted frequency step is actually where is the actual total time-span of the individual time series.
+This represents an oversampling of a factor of 100, which gives direct access to the amplitude of the signal as the height of the peak.
+Based on the discussion detailed in the previous section, we decided that the adopted largest frequency to be explored is 4 d-1.
+This choice is rather atypical since it was chosen to correspond to the spin period.
+If the star indeed varies in a periodic way, the largest peak in the periodogram is an estimate of the possible period, but does not stand as definitive proof of its existence.
+We should consider that a periodicity is proven if a minimum of three identical cycles are observed.
 Although the long periods are indicative, they are associated to large errors due to the small number of cycles included in the computation.
-The method used is the one of Heck-Manfroid-Mersch (hereafter HMM: Heck et al.
+The method used here is that of Heck-Manfroid-Mersch (hereafter HMM: Heck et al.
 1985; Gosset et al.
@@
 This is further described in the next section.
-We keep in a list up to one hundred candidate frequencies.
+We keep in a list up to 100 candidate frequencies.
 This is more than the amount of independent ones in the classical Fourier.
@@
 on the basis of a photometric curve).
-In DR3 in particular, this possibility is only used for the subset of solutions of the type EclipsingSpectro combined with SB1 solutions (see section 7.6 by Siopis & Sadowski in Pourbaix et al.
+In DR3 in particular, this possibility is only used for the subset of solutions of the type EclipsingSpectro combined with SB1 solutions (see Sect.
+7.6 by Siopis & Sadowski in Pourbaix et al.
 2022; Siopis et al.
@@
 The actual statistic is thus
-If the significance level is very close to zero, the peak could not be considered as generated by the noise and thus a high suspicion is permitted that the outstanding peak is due to a deterministic periodic signal.
-In the case of even sampling of the time-series, such a significance could be easily computed.
+If the significance level is very close to zero, the peak could not be considered as generated by the noise; thus, there is considerable suspicion that the outstanding peak is due to a deterministic periodic signal.
+In the case of even sampling of the time series, such a significance could be easily computed.
 However, in the case of irregular sampling and particularly in the case of the Gaia sampling, the situation is much more complex.
@@
 In principle, this should be performed individually for each object characterised by their own sampling, but this is not affordable for such a large number of sources.
-The first dependency of the statistical distribution is on the number of data points in the time-series.
+The first dependency of the statistical distribution is on the number of data points in the time series.
 The secondary dependency concerns the actual distribution in time of the data points.
@@
 At fixed, each realisation of the sampling provides a particular ECDF function.
-The chosen length 10000 of the ECDF allows us to assign false alarm probabilities with a precision better than 0.5 % at 95 % of confidence.
+The choice of 10 000 in the framework of the ECDF allows us to assign false alarm probabilities with a precision better than 0.5 % at 95 % confidence.
 The look-up table contains the set of median ECDFs for every given amount of valid transits.
@@
 In addition, the signal-to-noise contrast increases with .
-When calculating the periodogram of an input time-series with the length , its corresponding adopted ECDF (on the basis of the value) is read from the look-up table.
+When calculating the periodogram of an input time series with the length , its corresponding adopted ECDF (on the basis of the value) is read from the look-up table.
 The table has a comprehensive coverage of the number of transits, without gaps from = 10 to 72, but there are some entries for large values that are not present.
 In this case, the algorithm opts for the closest lowest input.
-5 The calculation of the orbital solutions
+5 Calculating the orbital solutions
 Our main task is to derive simple orbital motion based on Keplerian models.
@@
 The Keplerian fitting, even at fixed period, is time-consuming if it is to be performed over a vast set of periods: as for example in a genuine Keplerian periodogram (see e.g.
-Cumming 2004; Zechmeister & Kürster 2009, their sections 2.3 and 5, respectively).
+Cumming 2004; Zechmeister & Kürster 2009, their Sects.
+2.3 and 5, respectively).
 Therefore, we decided to strew subtleties over our procedure in order to render it much more robust.
@@
 However, the detection of a periodicity does not imply that the motion is Keplerian.
-Therefore, after the establishment of a list of candidate periods, it remains to fit the Keplerian model for each of them and to deduce the orbital solution and corresponding parameters that better fit the data.
-We first perform a Keplerian fit that is linearised at the maximum (but not entirely) and represents a preliminary fit, not definitive but robust.
-The actual least-square fit is performed with the true anomaly as the independent variable.
-The Fourier periodograms have the reputation to be inefficient at addressing the case of eccentric RV curves (for , see e.g.
+Therefore, after the establishment of a list of candidate periods, the Keplerian model needs to be fit for each of them to deduce the orbital solution and corresponding parameters that are a better fit to the data.
+We first performed a Keplerian fit that is linearised at the maximum (but not entirely) and represents a preliminary fit, not definitive but robust.
+The actual least-square fit was performed with the true anomaly as the independent variable.
+The Fourier periodograms are known to be inefficient at addressing the case of eccentric RV curves (for , see e.g.
 Endl et al.
 2002; Cumming 2004).
-We tend to think that this is due to the fact that, for eccentric orbits by opposition to the circular case, not all the data points contain the same amount of information.
+We tend to think that this is due to the fact that for eccentric orbits by opposition to the circular case, not all the data points contain the same amount of information.
 In fact, for very eccentric orbits, the data points related to observations around the maximum separation are more determining, in particular for the semi-amplitude.
@@
 After several tests, the adopted method appeared to be faster, more stable, trustworthy and robust.
-6 The main NSS-SB1 processing chain
+6 Main NSS-SB1 processing chain
 6.1 Global structure of the chain
@@
 These are InputChecker, PeriodFinder, SB1Refiner, SB1CRefiner and TrendSolution.
-The package InputChecker reads the input time-series in the NSS data model and copy them into arrays of primitives that are shared with the entire pipeline.
+The package InputChecker reads the input time series in the NSS data model and copies them into arrays of primitives that are shared with the entire pipeline.
 This reduces the number of modifications required after every data model update.
@@
 After these computations, a decision is to be taken to adopt the most adapted periodic orbital solution.
-The fifth package (TrendSolution) deals with the RV time-series exhibiting a trend as a function of time (see Sect.
+The fifth package (TrendSolution) deals with the RV time series exhibiting a trend as a function of time (see Sect.
 3.3).
-This means time-series that can be fitted by polynomials.
+This means time series that can be fitted by polynomials.
 These are acceleration models.
-After the execution of all these packages, a choice is made to select the solution that presents the best fit: SB1, SB1C, TrendSB1 or StochasticSB1.
+After the execution of all these packages, a choice is made to select the solution that presents the best fit: SB1, SB1C, TrendSB1, or StochasticSB1.
 The five packages are fully described in the respective five Sects.
-6.3.1, 6.3.2, 6.3.3, 6.3.4 and 6.3.5.
+6.3.1, 6.3.2, 6.3.3, 6.3.4, and 6.3.5.
 The flow chart of the full NSS-SB1 processing is described in Fig.
-4 along with the relations between the five packages mentioned above.
-It is interesting to notice that after the upper violet lozenge that contains a test on (that will be defined in Sect.
+4, along with the relations between the five packages mentioned above.
+It is interesting to notice that after the upper violet lozenge that contains a test on (defined in Sect.
 6.3.1), there is a bifurcation towards two sub-chains that receive the data.
@@
 From the purely logical point of view, they are working in parallel although from the practical point of view, they are executed in sequence.
-During the processing of the full NSS-SB1 chain, some flags are raised to retain more information about the path followed by the time-series, as well as to point out a few characteristics that could be useful for further treatment.
+During the processing of the full NSS-SB1 chain, some flags are raised to retain more information about the path followed by the time series, as well as to point out a few characteristics that could be useful for further treatment.
 In Sects.
-6.3.1, 6.3.2, 6.3.3, 6.3.4 and 6.3.5, the flags are identified by their number and a full description of their nature is given in Sect.
+6.3.1, 6.3.2, 6.3.3, 6.3.4, and 6.3.5, the flags are identified by their number and a full description of their nature is given in Sect.
 6.5.
@@
 Indeed, when eccentricity approaches zero, and parameters become progressively redundant (Sterne 1941), and thus the eccentric model becomes overstated, too powerful and thus prone to overfitting.
-The boundary between the two regimes, circular and eccentric, can not be set by a fixed value of the eccentricity since it is governed by the signal-to-noise ratio of the input time-series, and thus depends on many parameters such as the sampling, the amplitude of the signal and the involved noise (Lucy & Sweeney 1971).
+The boundary between the two regimes, circular and eccentric, cannot be set by a fixed value of the eccentricity since it is governed by the signal-to-noise ratio of the input time series, and thus depends on many parameters such as the sampling, the amplitude of the signal and the involved noise (Lucy & Sweeney 1971).
 Eigenvalue decompositions of matrices with rank smaller than 5 are performed analytically to improve robustness and reduce the processing time.
@@
 6.3.1 Ingestion of the data
-All stars that present RV variations according to the following criteria rv_chisq_pvalue0.01 and rv_normalised_gof4 are processed by the pipeline starting with the InputChecker described here.
+All stars that present RV variations according to the following criteria rv_chisq_pvalue0.01 and rv_renormalised_gof4 are processed by the pipeline starting with the InputChecker described here.
 Other details about these input data were already provided in Sect.
 2.
-NSS input data are requested from the Main Data Base (MDB) and copied into StarObject Java objects during data ingestion.
-A StarObject contains the single star astrometric solution (the reference epoch, the mean position, the proper motions, the parallax, ruwe and astrometric_excess_noise), the atmospheric parameters as measured by CU8 (Coordination Unit Astrophysical Parameters, Andrae et al.
+In the ingestion process, the NSS input data are requested from the Main Data Base (MDB) and copied into StarObject Java objects during data ingestion.
+A StarObject contains the single star astrometric solution (the reference epoch, the mean position, the proper motions, the parallax, ruwe, and astrometric_excess_noise), the atmospheric parameters as measured by CU8 (Coordination Unit Astrophysical Parameters, Andrae et al.
 (2023); Recio-Blanco et al.
-(2023)), and the light curve classification along with the orbital period if the corresponding source is suspected to be an eclipsing binary by CU7 (Coordination Unit Variability Analysis, Mowlavi et al.
+(2023)), as well as the light curve classification and the orbital period if the corresponding source is suspected to be an eclipsing binary by CU7 (Coordination Unit Variability Analysis, Mowlavi et al.
 (2023)).
@@
 Observation dates and error bars must be positive numbers, while RVs must fall within the range of km s-1.
-The length of the time-series is expected to be larger than or equal to 10 and if it is not the case a flag is raised (flag nr.8, see Sect.
+The length of the time series is expected to be larger than or equal to 10 and if it is not the case a flag is raised (flag nr.8; see Sect.
 6.5) and the star is rejected from further processing.
@@
 2024).
-The resulting time-series after cleaning contains data points distributed over a genuine individual time span (the difference between the two dates of observations at the extreme of the interval).
+The resulting time series after cleaning contains data points distributed over a genuine individual time span (the difference between the two dates of observations at the extreme of the interval).
 Finally, we perform a few modifications.
@@
 The reference epoch J2016.0 (BJD 2 457 389.0) of the astrometric solutions, is subtracted from the observation dates for reducing the correlations between the derived solution parameters.
-A cleaned time-series of RVs, still containing the required number of 10 observations, is shared with the whole software for the rest of processing, whilst, in the opposite case, it is rejected from the pipeline, and the corresponding flag is raised (flag nr.10, see Sect.
+A cleaned time series of RVs, still containing the required number of 10 observations, is shared with the whole software for the rest of processing, whilst, in the opposite case, it is rejected from the pipeline, and the corresponding flag is raised (flag nr.10; see Sect.
 6.5 for flag descriptions).
@@
 The package PeriodFinder is first establishing a list of candidate frequencies.
-The time-series of length is analysed thanks to an HMM periodogram, as explained in Sect.
+The time series of length is analysed thanks to an HMM periodogram, as explained in Sect.
 4.2.
@@
 From these and , the Campbell parameters and are easily extracted.
-In Equation (10), , , and appear in a linear way.
+In Eq.
+(10), , , and appear in a linear way.
 They can be optimised through an ordinary linear least-square regression and this for every set of trial .
@@
 The smallest value identifies the best fit and the best set of parameters, including the period.
-The algorithm is efficient at finding the best minimum, but the fit is performed in two steps which is clearly an approximation concerning the cross-covariance terms between families of parameters (linear versus non-linear).
+The algorithm is efficient at finding the best minimum, but the fit is performed in two steps, which is clearly an approximation concerning the cross-covariance terms between families of parameters (linear versus non-linear).
 Therefore, we need to refine the approximate model.
@@
 The module described here is the SB1CRefiner.
-The best approached period obtained in the PeriodFinder package is refined and fitted along with the linear parameters , and to fit a circular model to data (see equations in Sect.
+The best approached period obtained in the PeriodFinder package is refined and fitted along with the linear parameters , and to fit a circular model to data (see the equations in Sect.
 3.2).
@@
 We stop iterating at the 4th order, or when the higher-degree model exhibits more than one extremum over the time baseline.
-This last criterion has been adopted in order to avoid any possible overlaps with the orbital models with periods between and (see below).
+This last criterion has been adopted to avoid any possible overlaps with the orbital models with periods between and (see below).
 Most of the objects entering the trend category are probably long-term periodic curves but with a timescale much longer than the present time span of the mission ().
@@
 This problem has been corrected for DR4.
-The coefficients of the polynomial are included in the database but the work is performed in an Hipparcos-like formalism (see section 2.3.3, ESA, 1997).
+The coefficients of the polynomial are included in the database but the work is performed in an Hipparcos-like formalism (see Sect.
+2.3.3, ESA, 1997).
 The coefficients saved in the database are related to the derivatives of the velocity curve and are expressed with respect to the DR3 reference epoch (J2016.0, equivalent to BJD 2 457 389.0).
 We thus decided to express the trends as a function of the various derivatives. The simple mathematical model is
-where is the time (in days) relative to the reference epoch. We thus fit to the time-series the polynomial
-with potentially the fixed to zero. In the case of the first degree polynomial ( = 0), we adopt the following rigorous formulae
+where is the time (in days) relative to the reference epoch. We thus fit to the time series the polynomial,
+with the value potentially fixed to zero. In the case of the first degree polynomial ( = 0), we adopt the following rigorous formulae,
 The and parameters are not correlated if .
 This is approximated by subtracting to the BJD dates the reference epoch (J2016.0) as a proxy of the middle of the current mission.
-Practically, not all actual time-series are centred on this value since some transits might be missing (e.g.
+Practically, not all actual time series are centred on this value since some transits might be missing (e.g.
 more numerous before the mid-epoch or alternatively after it).
 This is in any case still an approximation since it is implicitly assumed that the distribution of the sample times are homogeneously apportioned.
-This is certainly not true, particularly if one includes EPSL data.
+This is certainly not true, particularly if we include the EPSL data.
 In the case of the second degree polynomial ( fitted), we note that the fitted polynomial can be expressed as
-where is the total duration of the mission treated in the framework of DR3. We thus adopt the approximate formulae
+where is the total duration of the mission treated in the framework of DR3. We thus adopt the approximate formulae,
 The term shifting by has been introduced such that the contribution by the second derivative is tending to be neutral over the interval .
@@
 Actually, the procedure described above has been applied when orbital solutions presented a period lower than ; it concerns the regime where the orbital motion is expected, when selected, to be convincingly robust.
-For longer periods, the folding with the period does not permit to test the perfect overlap homogeneously at all phases.
+For longer periods, the folding with the period does not permit us to test the perfect overlap homogeneously at all phases.
 We found it preferable to adopt an alternative procedure.
@@
 All the flags involved in the NSS spectroscopic processing are enumerated in Appendix I.
-As an example, the SB1Refiner processes the time-series of an object and produces a singular matrix.
+As an example, the SB1Refiner processes the time series of an object and produces a singular matrix.
 The flag nr.15 is thus raised, giving to the global flag an additive value of 32 768.
@@
 Although the majority of the flags are important for the decision tree of the algorithm, some of them are purely indicative.
-This is the case of the flag nr.24 that draws the attention on the fact that the attributed eccentric orbit appears, in the eccentricity versus log diagram, outside the zone where the circularisation of the orbit could have partly taken place.
+This is the case of the flag nr.24 that draws the attention to the fact that the attributed eccentric orbit appears, in the eccentricity versus log diagram, outside the zone where the circularisation of the orbit could have partly taken place.
 The borderline used is inspired from the one defined by Halbwachs et al.
@@
 2024).
-7 The post-processing
+7 Post-processing
 The post-processing described here further cleans and distributes the results of the NSS-SB1 main processing chain.
@@
 The operational runs took place in summer 2021.
-In order to maintain the quality level of the various solutions, we defined an internal filtering of these solutions that will allow keeping the very best and most secure ones.
+In order to maintain the quality level of the various solutions, we defined an internal filtering of these solutions that would allow us to retain the very best and most secure ones.
 The aim is also to spot out and reject solutions that are problematic.
@@
 The corresponding selection was of course considered.
-Since the internal filtering is supposed to protect the results from spurious solutions and that, in the present purely spectroscopic SB-subcatalogue, no confirmation comes from other channels, we decided to be highly cautious for this step.
+Since the internal filtering is supposed to protect the results from spurious solutions and that, in the present purely spectroscopic SB subcatalogue, no confirmation comes from other channels, we decided to be highly cautious for this step.
 For the SB1 solutions, we discarded objects having at least one of the three flags nr.14, nr.15, nr.18 raised.
@@
 We reject solutions leading to .
-The value of the gap for remaining solutions is not propagated to the SB-subcatalogue but will be accessible in DR4 with the release of the epoch RVs.
+The value of the gap for remaining solutions is not propagated to the SB subcatalogue but will be accessible in DR4 with the release of the epoch RVs.
 The threshold value itself has been optimised on the basis of the various validation runs (see Sect.
@@
 Solutions in-between are kept but a flag (nr.13, i.e.
-8192, see Appendix I) is raised if they are below the defined threshold.
+8192; see Appendix I) is raised if they are below the defined threshold.
 Another natural statistic is the spectroscopic significance defined as .
@@
 A high significance is necessary to adopt a solution.
-However, one should not forget that is itself an estimator and that statistical fluctuations could give a value of that is exceedingly small, boosting the significance to very large values.
+However, we should not forget that is itself an estimator and that statistical fluctuations could give a value of that is exceedingly small, boosting the significance to very large values.
 The opposite case is also true.
@@
 We decided to filter out on the flags nr.15 and nr.18, although there could be some unfortunate interactions with the SB1 channel.
-We rejected solutions presenting a gap (in time) in the time-series larger than 0.3.
+We rejected solutions presenting a gap (in time) in the time series larger than 0.3.
 7.2 Combined solutions
 After the spectroscopic processing described here above, the solutions are delivered to the combiner.
-The combiner and combinations are described elsewhere (see section 7.7 by Gavras & Arenou in Pourbaix et al.
+The combiner and combinations are described elsewhere (see Sect.
+7.7 by Gavras & Arenou in Pourbaix et al.
 2022) and are not extensively addressed here.
@@
 Either the NSS combines objects from the spectroscopic channel with those issued from the photometric channel and present in the class EclipsingBinary, i.e.
-presenting eclipses in photometry (see section 7.6 by Siopis & Sadowski in Pourbaix et al.
+presenting eclipses in photometry (see Sect.
+7.6 by Siopis & Sadowski in Pourbaix et al.
 2022; Siopis et al.
@@
 If the tentative combination is successful, a global fit over the photometric and the spectroscopic data is performed.
-In such a case, the new (improved) parameters supersede the old one and are introduced in the catalogue in the new class EclipsingSpectro (see section 7.7 by Gavras & Arenou in Pourbaix et al.
+In such a case, the new (improved) parameters supersede the old one and are introduced in the catalogue in the new class EclipsingSpectro (see Sect.
+7.7 by Gavras & Arenou in Pourbaix et al.
 2022).
@@
 2023a) can be combined with objects having a solution issued from the spectroscopic pipeline.
-The relevant combined solutions and the related objects are then redirected towards the new class labelled AstroSpectroSB1 (see section 7.7 by Gavras & Arenou in Pourbaix et al.
+The relevant combined solutions and the related objects are then redirected towards the new class labelled AstroSpectroSB1 (see Sect.
+7.7 by Gavras & Arenou in Pourbaix et al.
 2022).
 New fits are here also performed and the old values of the parameters are also superseded.
-We recall that, in this case, the fit in DR3 is done in the framework of the formalism of Thieles-Innes (Halbwachs et al.
+We recall that in this case, the fit in DR3 is done in the framework of the formalism of Thieles-Innes (Halbwachs et al.
 2023; Gaia Collaboration et al.
@@
 Even after the combination process, it might be that an object appears in two classes.
-In the framework of DR3, these are the sole allowed/activated possibilities.
-The criteria to consider a combination opportunity are described in details in section 7.7 of Gaia archive documentation444https://gea.esac.esa.int/archive/documentation/GDR3/ (Pourbaix et al.
+In the framework of DR3, these are the sole allowed and/or activated possibilities.
+The criteria to consider a combination opportunity are described in details in Sect.
+7.7 of Gaia archive documentation444https://gea.esac.esa.int/archive/documentation/GDR3/ (Pourbaix et al.
 2022).
@@
 The classes EclipsingSpectro and AstroSpectroSB1 are not further considered in the present paper.
-In order to avoid losing interesting objects and in a purely prototypical context, the combination was performed before the internal filtering.
+To avoid losing any interesting objects (and in a purely prototypical context), the combination was performed before the internal filtering.
 It could have been more secure to perform it after the internal filtering.
@@
 8 Results
-DR3 is the first release providing orbital solutions for the RV time-series. As such, the present results must be considered preliminary even if the performances are already good.
+DR3 is the first release providing orbital solutions for the RV time series. As such, the present results must be considered preliminary even if the performances are already good.
 8.1 Tallies
-Some 33,812,183 objects have validated (non NULL) RV values in DR3.
-Among them some 995315 objects obey the input criteria.
-We recall them: RVs always between –1000 and +1000 km, brighter than 12, effective temperature between 3875 and 8125 K, and rv_normalised_gof above 4.
+Some 33 812 183 objects have validated (non NULL) RV values in DR3.
+Among them, some 995 315 objects obey the input criteria.
+We recall them: RVs always between –1000 and +1000 km, brighter than 12, effective temperature between 3875 and 8125 K, and rv_renormalised_gof above 4.
 Only the objects suspected of variability enter the orbital solution chain.
 Thus, an additional cut-off of rv_chisq_pvalue at 0.01 (kept below 0.01) further decreases the sample.
-A little more than 866 kilo-objects (866403) entered the present main NSS-SB1 processing.
+A little more than 866 kilo-objects (866 403) entered the present main NSS-SB1 processing.
 If we consider the SB1+SB2 processing, the total amount is 876 644.
@@
 Table 5 deals with the very few objects coming from the SB2 chain.
-Due to the filtering on flags nr.11 and nr.12, these objects do not pervade the final SB-subcatalogue.
+Due to the filtering on flags nr.11 and nr.12, these objects do not pervade the final SB subcatalogue.
 No class is associated to these objects.
 We have now to estimate the number of solutions from our pipeline that were removed because they were combined and are now appearing in the EclipsingSpectro and AstroSpectroSB1 classes.
-Since we are only interested here on the disappearance of some solutions from our output due to the combination process, we can formally consider the tallies as if the combination were made after the internal filtering.
+Since we are only interested here in the disappearance of some solutions from our output due to the combination process, we can formally consider the tallies as if the combination were made after the internal filtering.
 The results of the combination mechanism for the 198 885 SB1-type solutions are detailed in Table 5.
-The column Single corresponds to the solutions appearing in one class only whereas the column Duplicated refers to the objects appearing in two classes.
+The column marked ’Single’ corresponds to the solutions appearing in one class only, whereas the column marked ’Duplicated’ refers to the objects appearing in two classes.
 All the 198 885 objects are classified, except 2201 that will not be published due to a rejection during the process.
@@
 Some 94 other objects were rejected for other reasons.
-The majority of these 94 unpublished objects is corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB-subcatalogue.
+The majority of these 94 unpublished objects are corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB subcatalogue.
 The Orbital and Others classes are given for the sake of completeness but they are not further considered either in the present work nor, for some, in DR3.
@@
 (2023a), some 155 objects are in this class.
-Over these 155 objects, 1 was originally attributed a solution of type SB2C, one SB1C, 2 SB2 and 151 SB1.
+Over these 155 objects, 1 was originally attributed a solution of type SB2C, 1 of type SB1C, 2 of type SB2, and 151 of type SB1.
 Among these 151 SB1 solutions coming out of our main processing, 79 were internally filtered out.
@@
 The details are explicited in Table 7.
-The majority of the 21 unpublished objects is corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB-subcatalogue.
+The majority of the 21 unpublished objects are corresponding to tentative combinations that finally failed and the objects were not re-injected in our SB subcatalogue.
 In this table, we also counted the TrendSB1 objects that presented or not a significant period at the level of the main processing (flag nr.13 raised means no period found).
@@
 However, it should be made clear that it is partly constrained by the fact that the spectroscopic solutions are further used by the combiner, and thus we should be careful not to eliminate good solutions that could be useful downstream in the process.
-In the present section, we restrict ourselves to the effect of the applied post-filtering on the countings of the objects among the various classes.
+In the present section we restrict ourselves to the effect of the applied post-filtering on the countings of the objects among the various classes.
 Table 6 details the effect on the number of objects considered in the important classes.
@@
 8.2.1 Good results
-In the present section, we introduce a few examples of the solution fitted to selected RV time-series.
+Here, we introduce a few examples of the solution fitted to selected RV time series.
 The selection is somewhat arbitrary and is essentially motivated by the illustration of the efficiency of the chain.
-All these illustrating good solutions are part of the DR3 spectroscopic orbit catalogue (SB-subcatalogue).
+All these illustrating good solutions are part of the DR3 spectroscopic orbit catalogue (SB subcatalogue).
 Figure 54 illustrates various orbital solutions that are part of the SB1 class.
@@
 At mid to large period, we also present a few examples of highly eccentric orbits, up to .
-It is a matter of fact that a good fit to a very eccentric orbit necessitates a large number of data points in the time-series, particularly near the periastron.
+It is a matter of fact that a good fit to a very eccentric orbit necessitates a large number of data points in the time series, particularly near the periastron.
 An homogeneous distribution in phase is another necessity.
-Since the vast majority of the analysed time-series covers a span of time around 800-1000 d, the solutions with short periods are extremely convincing since the large number of cycles folded in the phase diagram do not mitigate the goodness of fit and this can be associated to a high coherency of the periodicity.
+Since the vast majority of the analysed time series covers a span of time around 800-1000 d, the solutions with short periods are extremely convincing since the large number of cycles folded in the phase diagram do not mitigate the goodness of fit and this can be associated to a high coherency of the periodicity.
 This conclusion holds when more than three cycles are covered (i.e.
@@
 For example, someone interested in objects in circular orbits, i.e.
-to study only genuine circular orbit should certainly consider both SB1 and SB1C classes.
+to study only genuine circular orbit, should certainly consider both SB1 and SB1C classes.
 If a user wants to transform an SB1-type solutions into an SB1C one, they can certainly use the period, the centre-of-mass velocity and the semi-amplitude as given by the SB1-type solution.
@@
 Additional clues are necessary to converge to a sound classification.
-The binary nature of SB2 objects is much more convincing (although problems could also exist, see Damerdji et al.
+The binary nature of SB2 objects is much more convincing (although problems could also exist; see Damerdji et al.
 2024).
@@
 7.
-A general finding is the fact that most of the concerned time-series are characterised by a rather limited number of data points.
+A general finding is the fact that most of the concerned time series are characterised by a rather limited number of data points.
 They also correspond to objects characterised by .
@@
 This does not seem to be coherent with our results (d).
-The solution of this discrepancy will certainly be automatically found in future releases thanks to the important increase of the number of data points in the time-series under analysis.
+The solution of this discrepancy will certainly be automatically found in future releases thanks to the important increase of the number of data points in the time series under analysis.
 This conclusion is supported by the simulations reported in Sect.
@@
 The solutions shown in Figs.
-54, 57 and 58 are just illustrative and much more objects with solutions of similar quality are present in the SB-subcatalogue.
+54, 57, and 58 are just illustrative and much more objects with solutions of similar quality are present in the SB subcatalogue.
 This subcatalogue also contains solutions of slightly less good quality as well as solutions of intermediate quality.
-In the Appendix E, we provide a few examples of SB1-type solutions selected from the SB-subcatalogue at random but with around 0 (Fig.
+In Appendix E, we provide a few examples of SB1-type solutions selected from the SB subcatalogue at random but with around 0 (Fig.
 59).
@@
 We present here a typical case of badly handled data.
-Figure 9 shows the time-series of a star observed by Gaia RVS ( d).
+Figure 9 shows the time series of a star observed by Gaia RVS ( d).
 The data are unpublished and the solution is not part of the DR3 catalogue.
@@
 Figure 10 illustrates the fitted orbital model published by Griffin (2007).
-The ground observations at phase 0.98 (almost at larger separation, see figure 5 of Griffin 2007) suggest that the lines of the two stars are almost deblended.
+The ground observations at phase 0.98 (almost at larger separation; see Fig.
+5 of Griffin 2007) suggest that the lines of the two stars are almost deblended.
 | Parameter | Griffin | Gaia |
 | (d) | 1175.11.3 | 977.9104.3 |
-| (km s-1)∗*∗∗*∗The systemic velocity is expressed in the Griffin RV reference system (see also Sect. 9.4.1) | –9.560.06 | –9.940.10 |
+| (km s-1)∗*∗∗*∗The systemic velocity is expressed in the Griffin RV reference system (see also Sect. 9.4.1). | –9.560.06 | –9.940.10 |
 | (km s-1) | 16.380.11 | 1.188 |
@@
 10.
-The red dots represent the RVs derived by the single-line channel (they are put in the Griffin reference system by adding 1 km s-1, see Sect.
+The red dots represent the RVs derived by the single-line channel (they are put in the Griffin reference system by adding 1 km s-1; see Sect.
 9.4.1 below).
@@
 There is a tendency for the position of the blended lines to moderately follow the behaviour of the primary variation.
-According to (Griffin 2007), the lines of the secondary are slightly broader and the lines of the primary slightly deeper.
+According to Griffin (2007), the lines of the secondary are slightly broader and the lines of the primary slightly deeper.
 This thus implies that the primary lines are moderately dominating the blend.
@@
 This could thus be termed a spectrocentre effect888It should however be pointed out that this effect is mainly dependent on the characteristics of both sets of lines (with a different contribution from line to line) and thus on the Astrophysical Parameters characterising the two stars..
-It is interesting to notice that, in the case of HD 223323, the most affected parameter is essentially the semi-amplitude.
-On the opposite, it is clear from Table 7 that the other parameters are not strongly perturbed.
+It is interesting to notice that in the case of HD 223323, the most affected parameter is essentially the semi-amplitude.
+On the contrary, it is clear from Table 7 that the other parameters are not strongly perturbed.
 The eccentricity in particular is well determined, as well as the argument of periastron.
@@
 An illustrative case is the one of Gaia DR3 702159502868472832.
-Its RV time-series suggests a significant period of 1.0844000.000058 d with a probability of 0.997 and a resulting folded RV curve that is given in Fig.
+Its RV time series suggests a significant period of 1.0844000.000058 d with a probability of 0.997 and a resulting folded RV curve that is given in Fig.
 11.
-With a significance of 15.41 and an , nothing allows us to doubt of the nature of this object.
+With a significance of 15.41 and an , there is no solid reason to doubt the nature of this object.
 However, besides the Gaia RV curve and the fit, we plot also in the figure the few folded RVs that can be found in the literature from the APOGEE survey DR17 (Abdurro’uf et al.
@@
 Again the RV variations are convincing as well as the fit.
-The object is one of the very few ones that have been observed at several epochs in the framework of the LAMOST survey.
+The object is one of very few that have been observed at several epochs in the framework of the LAMOST survey.
 All the corresponding RVs are contained between the two green lines, further underlining the incompatibility of the ground-based data with the Gaia ones.
 The fact that two ground-based surveys disagree with the results from Gaia underlines the fact that some instrumental effects are most probably present in the Gaia data.
-It is particularly intriguing that the RV time-series exhibit what could appear as a convincing periodic variability characterised by timescale not known to be problematic.
-Clearly, the problem is linked to the production of the RV time-series rather than to the behaviour of the algorithm of orbital parameter determination, except perhaps for the period.
+It is particularly intriguing that the RV time series exhibit what could appear as a convincing periodic variability characterised by timescale not known to be problematic.
+Clearly, the problem is linked to the production of the RV time series rather than to the behaviour of the algorithm of orbital parameter determination, except perhaps for the period.
 9 Validation of the results
@@
 The aim is evidently to test and fine-tune the pipeline itself in order to address its efficiency but also to discern the unavoidable biases.
-The understanding of the pipeline behaviour is directly related to the quality of the SB-subcatalogue and looking at its characteristics from all angles is a necessity.
+The understanding of the pipeline behaviour is directly related to the quality of the SB subcatalogue and looking at its characteristics from all angles is a necessity.
 This validation process can be made on the basis of the internal consistency of the solutions but also on the basis of a comparison between the parameters delivered from the Gaia data with those obtained from the literature.
@@
 Both sets of RV curves are generated with a random phase.
-We considered a number of data points of 10, 17, 30 and 80.
+We considered a number of data points of 10, 17, 30, and 80.
 The times of observations are mimicking the Gaia time series and are drawn in the same manner as in Sect.
 4.4.
-The adopted signal-to-noise ratio is considered for values 5, 10 and 50.
-Fig.
-13 reveals the recovery rate (in percent) for various true periods.
+The adopted S/N ratio is considered for values 5, 10, and 50.
+Figure 13 reveals the recovery rate (in percent) for various true periods.
 The period is assumed to be recovered if its value does not differ from the input period by more than .
-For equal to or larger than 10, the recovery rate is mostly better than 80 percent except when the number of data points is too low, see e.g.
+For equal to or larger than 10, the recovery rate is mostly better than 80 percent except when the number of data points is too low (see e.g.
 .
@@
 The basic reason for this choice is that the principal motivation is to first scrutinise the behaviour of the pipeline.
-The post-filtering is based on information from other channels and could thus be more complex to interpret, although it is necessary in order to constitute the DR3 SB-subcatalogue.
+The post-filtering is based on information from other channels and could thus be more complex to interpret, although it is necessary in order to constitute the DR3 SB subcatalogue.
 In any case, for DR3, the differences between the analysed data set and the data set after post-filtering that contains 181 327 objects, are almost negligible from a statistical point of view.
@@
 This cut-off below 10 transits should not discard too many objects.
-Typically, we have slightly over 20 data points in our various time-series.
+Typically, we have slightly over 20 data points in our various time series.
 The largest value is 228 (not shown here) but these objects are exceptional and are corresponding to the EPSL observations.
@@
 We immediately notice that no object is present between and 5.5.
-This is an artefact of the selection based on the rv_renormalized_gof at the level of the input data.
+This is an artefact of the selection based on the rv_renormalised_gof at the level of the input data.
 This quantity is defined when and thus below this threshold, the value for these objects contains NaN even if these bright objects should be kept.
@@
 2023) for objects fainter than (corresponding for a solar-like star to ).
-Therefore, the difficulties to measure RVs on noisy spectra, the selection of variable objects in noisy time-series and the subsequent fit of orbital solutions conspire to induce this incompleteness.
+Therefore, the difficulties to measure RVs on noisy spectra, the selection of variable objects in noisy time series and the subsequent fit of orbital solutions conspire to induce this incompleteness.
 The observed distribution of the quality-of-fitting statistic deduced from the chi-square, the goodness of fit , is given for the filtered data set of solutions (see Fig.
@@
 A more sophisticated study, beyond the scope of the present paper, is necessary.
-However, one should not be misled by the normalisation of the comparison Gaussian curves.
-Actually, if we normalised with respect to the area, there would appear a deficit (roughly 25%) of objects with solutions characterised by an broadly located around zero (essentially between -1.
+However, we should not be misled by the normalisation of the comparison Gaussian curves.
+Actually, if we normalised with respect to the area, there would appear a deficit (roughly 25%) of objects with solutions characterised by an broadly located around zero (essentially between –1.
 and +1.).
@@
 This is dominantly an artefact of the decision we took to enforce an SB1 solution in this region of periods against the possibility to consider trends.
-Indeed, let us consider a time-series of a noisy linear trend.
+Indeed, let us consider a time series of a noisy linear trend.
 The enforced SB1 solution will deliver a period roughly about (only one ‘cycle’ but this depends on the exact sampling).
-The trend folded in phase will very often mimic an eccentric curve with a jump between the last data point of the time-series and the first one.
+The trend folded in phase will very often mimic an eccentric curve with a jump between the last data point of the time series and the first one.
 Therefore, this jump, after folding, will be associated to the passage at periastron and, consequently, the returned will be attributed a value around .
@@
 20.
-One clearly sees that the excess of solutions is effectively due to the range of periods enforced to an SB1 solution (prominent zone in yellow in the bottom right panel, not present on the upper left and right ones).
+We can clearly see that the excess of solutions is effectively due to the range of periods enforced to an SB1 solution (prominent zone in yellow in the bottom right panel, not present on the upper left and right ones).
 It is also clear that the effect is absent for quasi-circular orbits (which are fully correct) and is associated to larger eccentricities.
-In addition, one clearly sees that the solutions at the basis of the excess are confined to values of around and (very same zone in yellow).
+In addition, we can clearly see that the solutions at the basis of the excess are confined to values of around and (very same zone in yellow).
 This indicates that the trends enforced into an SB1 solution generate a RV curve that is predominantly of the saw-tooth type, further confirming our conclusions.
@@
 The critical point when fitting the orbital SB1 models is the determination of the period (see above).
-Fig.
-25 displays the histogram of the period (in log).
+Figure 25 displays the histogram of the period (in log).
 Several features are visible.
@@
 On the other hand, the deficit of solutions around the period of 182.6 days is due to the yearly Earth orbit that the Gaia satellite follows by staying at L2.
-The satellite avoiding the direction of the Sun, this produces regular gaps with a timescale of half a year.
+The satellite’s avoidance of the direction of the Sun produces regular gaps with a timescale of half a year.
 These gaps presented by the sampling are associated to large gaps in phase leading to the removal of the solutions because they correspond to badly determined periods and orbits.
-Fig.
-27 is showing a diagram of the value of the largest gap in the solution as a function of the period.
+Figure 27 is showing a diagram of the value of the largest gap in the solution as a function of the period.
 Exceptionally, here we use the data set of 367 328 objects constituting the set of SB1 solutions before the application of the internal filtering.
@@
 25, is an artefact of the application of an ad hoc selection criterion (see the paragraph on the internal filtering of the period in Sect.
-7.1) that prevents the proliferation of the small periodicities.
-This effect should be further investigated for the next Data Releases.
+7.1) that prevents the proliferation of small periodicities.
+This effect should be further investigated in the next data releases.
 In Fig.
@@
 The maximum accepted value for being 250 km s-1, the maximum significance value corresponds to a typical uncertainty of about 0.3 km s-1 which is coherent with expectation.
-The distribution shows no inflection locus that could help to set a selection threshold.
+The distribution shows no inflection locus that could be useful in setting a selection threshold.
 In this context, we decided to introduce a threshold at 5 but not more restrictive ones.
@@
 We leave the possibility to the catalogue users to define a threshold that could be appropriate for their own work.
-Fig.
-29 exhibits the distribution of the standard deviations on the eccentricity.
+Figure 29 exhibits the distribution of the standard deviations on the eccentricity.
 The typical uncertainty on the eccentricity is of 0.03-0.10 .
@@
 31 represents the plot of the derived SB1 mass function as a function of the adopted period.
-One clearly sees that the majority of the data points are between 0.001 and 0.5 M☉.
+We can clearly see that the majority of the data points are between 0.001 and 0.5 M☉.
 Objects outside this range are very few, in particular on the side of the high mass function.
@@
 Data points at an unexpected location could be due to problems in the derivation of the period or of the eccentricity.
-In particular, one should be very careful while searching for abnormal mass function values or massive black holes.
+In particular, we should be very careful while searching for abnormal mass function values or massive black holes.
 In addition, the overdensities tracing vertical lines at days and at one year (much broader) are due to objects with a varying apparent position inducing systematic effects on the RVs associated to the variation of the scan-angle.
@@
 This statement needs further confirmation.
-Fig.
-33 illustrates the distribution of the first derivative normalised by its uncertainty.
+Figure 33 illustrates the distribution of the first derivative normalised by its uncertainty.
 The distribution is very similar for the two classes and presents an almost perfect symmetry, as expected.
@@
 2023) have also an SB1 solution.
-They should be eventually excluded from our SB-subcatalogue but they also offer the unique opportunity to have an independent test of the spectroscopic periods by comparing them to the photometric ones.
+They should be eventually excluded from our SB subcatalogue but they also offer the unique opportunity to have an independent test of the spectroscopic periods by comparing them to the photometric ones.
 Indeed, we compared our spectroscopic periods to those determined from DR3 photometric data for the 147 RR Lyrae and the 333 Cepheids.
@@
 The breakdown of DR3 spectroscopic solution types is as follows for Griffin’s SB1 catalogue:
-SB1: 83 systems
-SB1C: 0 system
-TrendSB1: 10 systems
-SB2: 1 system
-SB2C: 0 system
+SB1: 83 systems,
+SB1C: 0 system,
+TrendSB1: 10 systems,
+SB2: 1 system,
+SB2C: 0 system,
 Without any DR3 spectroscopic deterministic solutions: 320 systems.
@@
 The most recent orbit was selected when a given system had multiple entries in the catalogue.
-As described above, visual/multiple systems were also excluded.
+As described above, visual and/or multiple systems were also excluded.
 Six SB2 orbits were replaced by those determined by Lester (2020).
 This validation catalogue contains the orbits of 1750 SB1’s and 647 SB2’s.
-The breakdown of DR3 spectroscopic solution types is as follows for SB1’s in the SB9 catalogue:
-SB1: 226 systems
-SB1C: 0 system
-TrendSB1: 45 systems
-SB2: 3 systems
-SB2C: 0 system
+The breakdown of DR3 spectroscopic solution types is as follows for SB1s in the SB9 catalogue:
+SB1: 226 systems,
+SB1C: 0 system,
+TrendSB1: 45 systems,
+SB2: 3 systems,
+SB2C: 0 system,
 Without any DR3 spectroscopic deterministic solutions: 1476 systems.
-All but one of the 45 systems with a trend solution have a very long period. Eighteen and two systems among the SB2 validation catalogue have received a SB1 and TrendSB1 solution, respectively.
+All (but one) of the 45 systems with a trend solution have very long periods. Furthermore, 18 and 2 systems in the SB2 validation catalogue have received a SB1 and a TrendSB1 solution, respectively.
 Figures 64, 65, 66, and 67 show the comparison between the orbital parameters. The deviations in expressed in units are shown in Fig. 39, as a function of the reference period.
@@
 9.4.3 Comparison against APOGEE
-Orbital parameters for 20,000 SB1’s have been released as part of APO Galactic Evolution Experiment (APOGEE)131313http://www.sdss3.org/surveys/apogee.php DR16 (Price-Whelan et al.
-2020a, hereafter PW20).
+Orbital parameters for 20 000 SB1’s have been released as part of APO Galactic Evolution Experiment (APOGEE)141414http://www.sdss3.org/surveys/apogee.php DR16 (Price-Whelan et al.
+2020, hereafter PW20).
 Our validation sample is their ’Gold Sample’ that contains 1032 systems with high-quality, unimodal posterior samplings.
-The breakdown of DR3 spectroscopic solution types is as follows for the SB1’s in PW20:
-SB1: 126 systems
-SB1C: 0 system
-TrendSB1: 4 systems
-SB2: 3 systems
-SB2C: 0 system
+The breakdown of DR3 spectroscopic solution types is as follows for the SB1s in PW20:
+SB1: 126 systems,
+SB1C: 0 system,
+TrendSB1: 4 systems,
+SB2: 3 systems,
+SB2C: 0 system,
 Without any DR3 spectroscopic deterministic solutions: 899 systems.
@@
 9.4.4 Comparison against WISE
-A total of 56,000 short-period ( 10 d) binary systems observed by the Wide-field Infrared Survey Explorer (WISE)141414https://www.jpl.nasa.gov/missions/wide-field-infrared-survey-explorer-wise have been discussed by Petrosky et al.
+A total of 56 000 short-period ( 10 d) binary systems observed by the Wide-field Infrared Survey Explorer (WISE)151515https://www.jpl.nasa.gov/missions/wide-field-infrared-survey-explorer-wise have been discussed by Petrosky et al.
 (2021).
@@
 The infrared photometric measurements are made through the W1 filter centred at about 3.4 m.
-A subset of 49,465 systems can be unambiguously associated to a DR3 source151515http://cdn.gea.esac.esa.int/Gaia/gedr3/cross_match/.
+A subset of 49 465 systems can be unambiguously associated to a DR3 source161616http://cdn.gea.esac.esa.int/Gaia/gedr3/cross_match/.
 This sample allows us to test the reliability of our orbital periods for very close binaries, although other orbital parameters are not available.
-The breakdown of DR3 spectroscopic solution types is as follows for the SB1’s in Petrosky et al. (2021):
-SB1: 704 systems
-SB1C: 3 systems
-TrendSB1: 12 systems
-SB2: 69 systems
-SB2C: 23 systems
-Without any DR3 spectroscopic deterministic solutions: 48,654 systems.
+The breakdown of DR3 spectroscopic solution types is as follows for the SB1s in Petrosky et al. (2021):
+SB1: 704 systems,
+SB1C: 3 systems,
+TrendSB1: 12 systems,
+SB2: 69 systems,
+SB2C: 23 systems,
+Without any DR3 spectroscopic deterministic solutions: 48 654 systems.
 We find a close correspondence in most cases with either the main WISE period or its second harmonic (Fig.
@@
 9.4.5 Comparison against ASAS-SN
-The physical parameters of 35,464 detached eclipsing binaries (DEBs) observed in the framework of the All-Sky Automated Survey for Supernovae (ASAS-SN; Kochanek et al.
-2017)161616https://www.astronomy.ohio-state.edu/asassn/ have been determined by Rowan et al.
+The physical parameters of 35 464 detached eclipsing binaries (DEBs) observed in the framework of the All-Sky Automated Survey for Supernovae (ASAS-SN; Kochanek et al.
+2017)171717https://www.astronomy.ohio-state.edu/asassn/ have been determined by Rowan et al.
 (2022).
-A total of 35,307 systems have a Gaia DR3 ID.
+A total of 35 307 systems have a Gaia DR3 ID.
 The data are based on the analysis of - and -band light curves.
@@
 The breakdown of DR3 spectroscopic solution types is as follows for the binaries in Rowan et al. (2022):
-SB1: 626 systems
-SB1C: 3 systems
-TrendSB1: 6 systems
-SB2: 92 systems
-SB2C: 35 systems
-Without any DR3 spectroscopic deterministic solutions: 34,545 systems.
+SB1: 626 systems,
+SB1C: 3 systems,
+TrendSB1: 6 systems,
+SB2: 92 systems,
+SB2C: 35 systems,
+Without any DR3 spectroscopic deterministic solutions: 34 545 systems.
 Most objects have 13 mag and are fainter than our magnitude cut-off.
@@
 (2022).
-One can also note the excellent success rate for the pulsation periods of the RR Lyrae stars that are below one day (Sect.
+We can also note the excellent success rate for the pulsation periods of the RR Lyrae stars that are below one day (Sect.
 9.3).
@@
 However, there may be more complex dependencies lurking in the data that cannot be easily and reliably quantified given the quite limited number of cross-matches with the validation catalogues (a few hundreds systems at most).
-In this respect, the fact that only a relatively small fraction of the SB1’s in the validation catalogues have a spectroscopic deterministic DR3 solution is the consequence of the various filtering steps, both at CU4 and CU6 levels, described previously.
+In this respect, the fact that only a relatively small fraction of the SB1s in the validation catalogues have a spectroscopic deterministic DR3 solution is the consequence of the various filtering steps, both at CU4 and CU6 levels, described previously.
 In addition, we note that for the SB9 and Griffin’s sample, only a subset of the reference catalogue was used.
@@
 10.1.1 Intrinsic variables
-Since the NSS-SB1 processing is applied to time-series of variable RVs, it is expected to also treat some intrinsic variables that are exhibiting well-defined RV variations.
+Since the NSS-SB1 processing is applied to time series of variable RVs, it is expected to also treat some intrinsic variables that are exhibiting well-defined RV variations.
 This could be the case for example for Cepheids (actually the bright ones).
-The only firewall preventing the processing to deliver spurious orbital solutions for intrinsic variables is that the processing only considers an harmonic content of the RV curve that is Keplerian (or nearly so).
+The only firewall preventing the processing from delivering spurious orbital solutions for intrinsic variables is the fact that the processing only considers an harmonic content of the RV curve that is Keplerian (or nearly so).
 Unfortunately, Cepheids RV curves, for example, are renowned for their harmonic content being quite similar to the Keplerian one.
@@
 As part of DR3, Ripepi et al.
-(2023) published a complete analysis of the Gaia photometric curves for a little more than 15,000 Cepheids (of all classes), 799 having time-series of RVs.
+(2023) published a complete analysis of the Gaia photometric curves for a little more than 15 000 Cepheids (of all classes), 799 having time series of RVs.
 This list was cross-correlated with our SB1/SB1C list.
@@
 After an eye inspection of photometric and RV curves, five of them were found doubtful Cepheids and have been rejected.
-Finally, 333 spurious SB1 solutions due to the Cepheid nature of the objects have been blacklisted from the SB-subcatalogue.
+Finally, 333 spurious SB1 solutions due to the Cepheid nature of the objects have been blacklisted from the SB subcatalogue.
 Along these lines, we also considered the case of RR Lyrae stars (of all classes too).
 Clementini et al.
-(2023) gathered a list of a little more than 270,000 such objects, among which some 1096 have time-series of RVs.
+(2023) gathered a list of a little more than 270 000 such objects, among which some 1096 have time series of RVs.
 The cross-match with our SB1 list yields 152 objects in common.
@@
 In summary, we rejected from our SB1 list some 480 sources as being intrinsic variables.
-Although a few pulsating stars (Cepheids, RR Lyrae) are rejected from our SB-subcatalogue, they are nevertheless studied as pulsators elsewhere (see e.g.
+Although a few pulsating stars (Cepheids, RR Lyrae) are rejected from our SB subcatalogue, they are nevertheless studied as pulsators elsewhere (see e.g.
 Eyer et al.
@@
 In addition, from the later provisional analysis, it became clear that this preliminary list most probably also contains a lot of other objects and in particular ellipsoidal variables.
-We will illustrate here our concerns by further studying the case of Gaia DR3 5829160851462523008.
+We illustrate here our concerns by further studying the case of Gaia DR3 5829160851462523008.
 This object was not known to vary before Gaia.
 It appears in the catalogue of LPV candidates with a derived photometric period of 90.5 d.
-This object also appears in the SB-subcatalogue, produced in the present work, as a binary with a period of 178.886 d.
+This object also appears in the SB subcatalogue, produced in the present work, as a binary with a period of 178.886 d.
 Figure 47 exhibits both the RV curve and the photometric light curve of this object with the phase computed according to the RV period.
@@
 2023c) where some 20 % of the LPV candidates are expected to be ellipsoidal variables.
-We encourage the user of the SB-subcatalogue to filter out true intrinsic variables thanks to these subsequent results.
+We encourage the user of the SB subcatalogue to filter out true intrinsic variables thanks to these subsequent results.
 Yet another source of contaminants are main-sequence pulsators.
@@
 2004).
-The problem of the rejection of intrinsic variables is certainly a difficult one and much progress is still to be made on this topic.
-As mentioned above, the measurement of RVs by cross-correlating the spectra with template prevents to identify the origin of the RV variations.
+The problem of the rejection of intrinsic variables is certainly a difficult one, and much progress is still to be made on this topic.
+As mentioned above, the measurement of RVs by cross-correlating the spectra with template prevents us from identifying the origin of the RV variations.
 10.1.2 Double-line transits
@@
 They are thus associated to large values of the ruwe parameter.
-This point was already mentioned in Sect.
+This point is also mentioned in Sect.
 2.
@@
 In order to alleviate the problem, the objects in the class SB1 were first selected on the basis of an astrometric ruwe larger than 1.4 and a solution leading to a period within 3% of the precession period (62.97 d).
-In addition, we selected objects obeying the condition
-where is the mean of the uncertainties on the individual epoch RVs over the full time-series.
-Fig.
-49 illustrates the distribution of the astrometric_excess_noise as a function of the parameter for the selected sample.
-One clearly sees that all the values are lower than three times the astrometric_excess_noise suggesting that these orbital solutions are spurious and are due to a scan-angle effect.
+In addition, we selected objects obeying the condition:
+where is the mean of the uncertainties on the individual epoch RVs over the full time series.
+Figure 49 illustrates the distribution of the astrometric_excess_noise as a function of the parameter for the selected sample.
+We can clearly see that all the values are lower than three times the astrometric_excess_noise suggesting that these orbital solutions are spurious and are due to a scan-angle effect.
 The semi-amplitude is too small relatively to the error induced by the astrometric solution.
@@
 These objects have not been rejected.
-We encourage the catalogue user to be very cautious when using the present SB-subcatalogue despite the fact that persisting problems should only represent a minority.
+We encourage the catalogue user to be very cautious when using the present SB subcatalogue despite the fact that persisting problems should only represent a minority.
 All the remaining and identified problems should be corrected in the future DR4.
@@
 Here below, we propose an explanation for this discrepancy by investigating in details the case of Gaia DR3 3376949338201650112.
-This example has the advantage to be enlightening and will allow us to correct the pipeline for future releases.
+This example has the advantage of being enlightening and will allow us to correct the pipeline for future releases.
 The astrometric solution for this object is characterised by a ruwe of 25.35, pinpointing to a problem with the fit.
-Referring to the IPD parameters (Image Parameter Determination, see e.g.
+Referring to the IPD parameters (image parameter determination; see e.g.
 Holl et al.
@@
 50 where it becomes clear that the measured RVs are function of the scan angle.
-The parameter ipd_gof_harmonic_phase adopts a value of 103.4 ° that suggests that the line joining the two stars of the pair has a position angle on the sky of 13.4 °, after the application of a 90 ° correction as explained in section 3.1.2 of Holl et al.
+The parameter ipd_gof_harmonic_phase adopts a value of 103.4 ° that suggests that the line joining the two stars of the pair has a position angle on the sky of 13.4 °, after the application of a 90 ° correction as explained in Sect.
+3.1.2 of Holl et al.
 (2023a).
-The semi-amplitude of the fake orbital solution is about 16.6 km s-1 which translates according to equation 2 of Holl et al.
+The semi-amplitude of the fake orbital solution is about 16.6 km s-1 which translates according to Eq.
+2 of Holl et al.
 (2023a) into a separation of the pair of 114 mas.
@@
 However, in the present case, the period is 0.657582 d which does not appear to be compatible and could shed some doubt on our interpretation.
-The spectral window associated to the DR3 RV time-series corresponding to this object is given in Fig.
+The spectral window associated to the DR3 RV time series corresponding to this object is given in Fig.
 51.
-Besides the main peak at d-1 as usual, it exhibits an important peak at d-1 that is exceeding a value of 0.8; this is an indication that isolated aliasing could be important for this particular time-series.
-An aliasing with d-1 explains the ambiguity between a period of 0.657582 d and one of one year.
-The combination of the aliasing phenomenon and of the scan-angle effect explains the existence of such fake SB1 objects in our SB-subcatalogue: such objects were not post-filtered, since the post-filtering was based on the 62.97 d or 1 y timescales.
+Besides the main peak at d-1 as usual, it exhibits an important peak at d-1 that is exceeding a value of 0.8; this is an indication that isolated aliasing could be important for this particular time series.
+An aliasing with d-1 explains the ambiguity between a period of 0.657582 d and that of 1 year.
+The combination of the aliasing phenomenon and of the scan-angle effect explains the existence of such fake SB1 objects in our SB subcatalogue: such objects were not post-filtered, since the post-filtering was based on the 62.97 d or 1 y timescales.
 We conclude that it would be more appropriate to include a statistical test inspired by graphs like the one shown in Fig.
@@
 10.3 The case of astrometric binaries
-In Sect.
-10.1.3 and in Sect.
-10.2, we discussed the effect of astrometric close pairs on the determination of RVs.
+In Sects.
+10.1.3 and 10.2, we discussed the effect of astrometric close pairs on the determination of RVs.
 A close pair fixed on the sky could often generate a scan-angle effect and lead to corrupted RV values but also could generate fake orbital motions.
-An additional degree of complexity is reached if one considers astrometric binaries.
+An additional degree of complexity is reached if we consider astrometric binaries.
 Due to their binarity, these close pairs are not fixed on the sky and the main axis of the pair is rotating due to the orbital motion.
@@
 In this latter case, the correction presents a timescale equal to the phenomenon we are studying.
-The correcting scheme is explained in section 3.3.1 of Holl et al.
+The correcting scheme is explained in Sect.
+3.3.1 of Holl et al.
 (2023a) and is also illustrated there with the (somewhat extreme) example of Gaia DR3 6631710607341412096.
@@
 For DR4, a preliminary run of the NSS astrometric channel has been included in order to compute a correction before the CU6 processing performing the RV measurements.
-11 The SB-subcatalogue
+11 The SB subcatalogue
 11.1 Description of the output
@@
 The first term of the vector gives the element (0,1) of the matrix, the element gives the element (0,2) of the matrix and element the (1,2) one.
-Elements , , correspond respectively to the matrix elements (0,3), (1,3) and (2,3) and so on.
+Elements , , correspond respectively to the matrix elements (0,3), (1,3), (2,3), and so on.
 This is thus a column-major storage scheme.
@@
 We recall that the StochasticSB1 solutions are not delivered in DR3.
-11.2 Weaknesses, caveats and recommendations
+11.2 Weaknesses, caveats, and recommendations
 The present section gathers the caveats dispersed all over the paper and delivers a summary of the characteristics of the catalogue that could be beneficial to the general user.
@@
 In addition, no such step was made for other intrinsic variables in DR3 (notably Long-Period Variables).
-Another case of trickery is the one of true SB2 stars, where the two lines never fully deblend.
+Another tricky case is that of true SB2 stars, where the two lines never fully deblend.
 They are appearing most probably in the SB1 channel.
@@
 Close astrometric pairs could thus mimic SB1 RV variations (see the work by Holl et al.
-2023a), and are polluting the present sample (even if some post-filtering took place, see also Sect.
+2023a), and are polluting the present sample (even if some post-filtering took place; see also Sect.
 10.1.3).
@@
 This problem will be corrected for DR4.
-Besides, it should be noted that all decisions are taken in a statistical way; this means that the output is statistically correct but that some incorrect individual fits can survive near the borders where the decisions are taken (e.g.
+In addition, it should be noted that all decisions are taken in a statistical way; this means that the output is statistically correct but that some incorrect individual fits can survive near the borders where the decisions are taken (e.g.
 between an orbital fit and a trend solution for example).
@@
 (2022).
-To some extent, other information can be considered as quality factors (such as flags but, mainly, the number of points in the time-series ).
-They could also be included in the selection process.
-The global weak point of the algorithms could be the determination of the period even if here above it is finally shown that the success rate is at the very least 80% and is more often located in the range 85-95%.
+To some extent, other information can be considered as quality factors (such as flags but, mainly, the number of points in the time series ), which could also be included in the selection process.
+The global weak point of the algorithms could be the determination of the period even if here above it is shown that the success rate is at the very least 80% and is more often located in the range 85-95%.
 There is evidently at a very low number of transits, a difficulty to correctly determine the period that forced us to neglect data with less than 10.
@@
 9.1, we conclude that a value of smaller than 17 is somewhat problematic (although it depends on the signal-to-noise ratio) for the period recovery.
-We also showed from the validation (see Sect.
+We also show, from the validation (see Sect.
 9.2), that an overfitting phenomenon characterises part of the solutions associated to an in the range 10-15.
-One should further add that, at small , the periodogram has a larger propensity to produce isolated aliasing peaks (generating ambiguity due to aliasing).
+We should further add that, at small , the periodogram has a larger propensity to produce isolated aliasing peaks (generating ambiguity due to aliasing).
 While interpreting the simulations, it appears that it is unclear if the determining parameter is the or the significance.
@@
 If the user of the catalogue only considers the sole results of the spectroscopy fits, it is advisable to restrict the sample to some kind of gold sample.
-This can be done by performing additional selections in the catalogue: for example, a larger number of transits, a higher significance, periods in a dedicated range (between ten days and , or allowing for going down to smaller periods but at the expense of a selection on the basis of a larger significance), period with small uncertainties, solutions with negligible phase gaps (although the gap values are not provided in the DR3 catalogue), etc…
+This can be done by performing additional selections in the catalogue: for example, a larger number of transits, a higher significance, periods in a dedicated range (between ten days and , or allowing it to go down to smaller periods but at the expense of a selection on the basis of a larger significance), periods with small uncertainties, solutions with negligible phase gaps (although the gap values are not provided in the DR3 catalogue), and so on.
 In Appendix J, we present indicative but simple examples of extraction of a gold sample from the catalogue.
 12 Conclusions
-In the framework of the Gaia Data Release 3, we present here the first version of the Gaia catalogue of spectroscopic orbital solutions (SB-subcatalogue).
-We present 181327 objects with a solution of class SB1 and 202 objects with a solution of class SB1C.
-We also present objects exhibiting trends of variations of their RVs with time.
-The present catalogue constitutes by far the largest collection of objects with a spectroscopic orbital solution, as can be evaluated by comparison with the notorious SB9 compilation171717https://sb9.astro.ulb.ac.be/.
-The improvement in number of concerned objects is more than one (almost two) order of magnitude.
-The present SB-subcatalogue has also the major advantage of being homogeneous from the point of view of both data acquisition (properties) and data processing, surpassing the characteristics of existing compilation.
-The only exception is the typical number of measurements that constitutes the individual time-series, which remains limited (at least for the present DR3 catalogue).
+Within the framework of the Gaia Data Release 3, we present the first version of the Gaia catalogue of spectroscopic orbital solutions (SB subcatalogue).
+We present 181 327 objects that have been found to be characterised by class SB1 and 202 objects by class SB1C.
+We also present objects exhibiting trends of variations of their RVs over time.
+The present catalogue constitutes the largest collection of objects with a spectroscopic orbital solution by far, based on a comparison with the notorious SB9 compilation181818https://sb9.astro.ulb.ac.be/.
+The improvement in terms of number of objects is more than one order (nearly two) of magnitude.
+The present SB subcatalogue has also the major advantage of being homogeneous from the point of view of both data acquisition (properties) and data processing, surpassing the characteristics of existing compilation.
+The only exception is the typical number of measurements that constitutes the individual time series, which remains limited (at least for the present DR3 catalogue).
 The situation will undoubtedly improve for future data releases.
-After a description of the data entering the NSS-SB1 processing, we describe in details the logic and structure of the related code.
-The general philosophy behind it is explicited along with the mathematical formalism used.
-We also describe, after the main processing, the filtering of the results that we performed to discard bad solutions and improve the catalogue.
-Examples illustrating the quality of the output are given as well as a few examples of detected problematic cases.
+After a description of the data entering the NSS-SB1 processing, we describe in detail the logic and structure of the related code.
+We reveal the general philosophy behind, along with the mathematical formalism used.
+We also describe, following the main processing details, the filtering of the results performed to discard bad solutions and improve the catalogue.
+We give a number of examples illustrating the quality of the output, as well as a few examples of detected problematic cases.
 A full validation of the output is performed both on the basis of internal consistency and on the basis of external existing catalogues from the literature.
-The characteristics of the catalogue are further analysed.
-The way the SB-subcatalogue is included in the Gaia Data Release 3 is also described.
-A section is included that scrutinises the weaknesses and provides caveats and recommendations for the scientific use of the catalogue.
+We further analyse the characteristics of the catalogue and also describe the way the SB subcatalogue has been included in the Gaia Data Release 3.
+We additionally scrutinises the weaknesses of the approach and provide the necessary caveats and recommendations for scientific applications of the catalogue.
 Data availability
-The epoch RVs constituting the time-series analysed here have not been published in the context of the Gaia Data Release 3.
-The orbital solutions and trends we derived are available since June 13, 2022, in the ESA Gaia DR3 archive accessible at https://gea.esac.esa.int/archive.
+The epoch RVs constituting the time series analysed here have not been published in the context of the Gaia Data Release 3.
+The orbital solutions and trends we derived are available since 13 June 2022, in the ESA Gaia DR3 archive accessible at https://gea.esac.esa.int/archive.
 The same results can also be found at CDS Strasbourg via the link https://cdsarc.cds.unistra.fr/viz-bin/Cat?I/357.
+Appendices D to J are only available on Zenodo (see https://zenodo.org/records/13990211).
 Acknowledgements.
@@
 2022, Gaia DR3 documentation Chapter 7: Non-single stars, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium.
-Online at ¡A href=“https://gea.esac.esa.int/archive/documentation/GDR3/index.html”¿ https://gea.esac.esa.int/archive/documentation/GDR3/index.html¡/A¿, id.
-7
+Online at https://gea.esac.esa.int/archive/documentation/GDR3/index.html, 7
 - Pourbaix et al. (2004) Pourbaix, D., Tokovinin, A. A., Batten, A. H., et al. 2004, A&A, 424, 727
@@
 - Price-Whelan et al. (2017) Price-Whelan, A. M., Hogg, D. W., Foreman-Mackey, D., & Rix, H.-W. 2017, ApJ, 837, 20
-- Price-Whelan et al. (2020a) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020a, ApJ, 895, 2
-- Price-Whelan et al. (2020b) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020b, ApJ, 895, 2
+- Price-Whelan et al. (2020) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020, ApJ, 895, 2
 - Ramos et al. (2018) Ramos, P., Antoja, T., & Figueras, F. 2018, A&A, 619, A72
@@
 2022, Gaia DR3 documentation Chapter 6: Spectroscopy, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium.
-Online at ¡A href=“https://gea.esac.esa.int/archive/documentation/GDR3/index.html”¿ https://gea.esac.esa.int/archive/documentation/GDR3/index.html¡/A¿, id.
-6
+Online at https://gea.esac.esa.int/archive/documentation/GDR3/index.html, 6
 - Sartoretti et al. (2018) Sartoretti, P., Katz, D., Cropper, M., et al. 2018, A&A, 616, A6
@@
 Appendix A Historical developments
-Once the period evaluated, the general model to fit (Eq.
-1) is still non-linear in some of the parameters (mainly the eccentricity ).
+Once the period has been evaluated, the general model to fit (Eq.
+1) remains non-linear in some of the parameters (mainly the eccentricity, ).
 This implies that the of the fit could present a few or even numerous minima in the parameter space, and to find the global one is a true challenge.
@@
 We acted along two main axes.
-On one hand, we tried to invent a method based on pattern recognition in the phase diagram of the RVs folded with the candidate periods.
-Such an approach would allow having a first guess for the eccentricity, which could be used for further classical derivation.
-In order to apply pattern recognition techniques, it is mandatory to define an interpolation function in the phase diagram in order to obtain a robust description of the shape of the RV curve.
-Several tools have been explored such that the nearest-neighbour interpolation, the linear interpolation, the interpolation by cosines, the local Akima periodic interpolation, the periodic interpolation by cubic splines, the periodic interpolation by Catmull-Rom splines.
+On the one hand, we tried to invent a method based on pattern recognition in the phase diagram of the RVs folded with the candidate periods.
+Such an approach would allow us to take a first guess for the eccentricity, which could then be used for further classical derivation.
+In order to apply pattern recognition techniques, it is mandatory to define an interpolation function in the phase diagram to obtain a robust description of the shape of the RV curve.
+Several tools have been explored such that the nearest-neighbour interpolation, the linear interpolation, the interpolation by cosines, the local Akima periodic interpolation, the periodic interpolation by cubic splines, and the periodic interpolation by Catmull-Rom splines.
 We also explored the possibilities of performing classical adjustment by least-square methods.
@@
 The use of genetic algorithms is a trade-off between the probability to effectively detect the global minimum and the consumption of computer time.
-The code using genetic algorithms were written, delivered and tested.
+The codes using genetic algorithms were written, delivered and tested.
 The detection of the global minimum was possible but was implying a prohibitively large amount of computer time despite strong efforts made to accelerate it.
@@
 (2017) and Price-Whelan et al.
-(2020b) pinpoints the strong impact that could have the choice of priors on the resulting posteriors.
+(2020) pinpoints the strong impact that could have the choice of priors on the resulting posteriors.
 The Bayesian methods have admittedly the big advantage to potentially provide marginal posterior probabilities for a series of models and thus estimate how much the observational data could favour a particular model.
@@
 In addition, these Monte Carlo methods are computer time-consuming and were not acceptable for the DR3 processing.
-We finally adopted a method based on fully analytical and classical numerical developments.
+Finally, we adopted a method based on fully analytical and classical numerical developments.
 Appendix B Practical details on the processing
-The Data Processing Centre at CNES (DPCC181818https://www.cosmos.esa.int/web/gaia/dpac) is located at the Centre National d’Etudes Spatiales in Toulouse (France).
-It ran the spectroscopic processing chain (CU6) and the object processing chain (CU4).
+The Data Processing Centre at CNES (DPCC191919https://www.cosmos.esa.int/web/gaia/dpac), located at the Centre National d’Etudes Spatiales in Toulouse (France), ran the spectroscopic processing chain (CU6) and the object processing chain (CU4).
 Globally, it ran the whole set of NSS codes and, in particular, the spectroscopic orbital solution chain described in the present paper.
@@
 The version of the code tagged and executed for the operational run was 20.10.
-Appendix C The separable variable technique applied to the Keplerian model fit
+Appendix C Separable variable technique applied to the Keplerian model fit
 C.1 A reminder of the Levenberg-Marquardt method
@@
 represents the Keplerian function, such that
-being the array of 6 parameters where the superscript denotes the transpose operation. The Newton-Raphson algorithm is based on the Taylor first order development
+with being the array of 6 parameters where the superscript denotes the transpose operation. The Newton-Raphson algorithm is based on the Taylor first order development
 where , and denotes the iteration number. is the Jacobian matrix which elements are
-The objective function to be minimised is the quadratic sum of residuals
+The objective function to be minimised is the quadratic sum of residuals as
 with being the inverse of the data variance-covariance matrix. The is given at the iteration by
 where contains the observed RVs. The optimal parameter step is obtained by solving
-leading in the case of Newton-Raphson to the algebraic system
+In the case of Newton-Raphson, this leads to the following algebraic system:
 The matrix is a positive semi-definite matrix called the hessian matrix (it is not a positive definite matrix because it might be singular).
 The Levenberg-Marquardt algorithm consists in rescaling the diagonal elements of by the damping factor . The system (Eq. 27) becomes
-is the damping factor, chosen to be large when approaches the singularity.
-C.2 The separable variable technique
-The parameters are linear because they satisfy the property
-Hereafter, subscripts and refer to linear and non-linear respectively. is the Jacobian matrix whose columns are the derivatives of with respect to .
-While in both Newton-Raphson and the (commonly used) Levenberg-Marquardt (LM) algorithms, linear and non-linear parameters are treated the same way, the separable variable technique has the advantage that can be obtained for every non-linear parameter by a linear least-square regression, such that 191919The inversion of the normal equations is known to be numerically unstable and should be avoided in favour of orthogonal decomposition approaches.
+where is the damping factor, chosen to be large when approaches the singularity.
+C.2 Separable variable technique
+The parameters are linear because they satisfy the following property:
+Hereafter, subscripts and refer to linear and non-linear respectively. Then, is the Jacobian matrix whose columns are the derivatives of with respect to .
+While in both Newton-Raphson and the (commonly used) Levenberg-Marquardt (LM) algorithms, the linear and non-linear parameters are treated the same way, the separable variable technique has the advantage that can be obtained for every non-linear parameter by a linear least-square regression, such that 202020The inversion of the normal equations is known to be numerically unstable and should be avoided in favour of orthogonal decomposition approaches.
 with
-Therefore, Eq. 22 becomes a function of only
+Therefore, Eq. 22 becomes a function of only, such that
 where the step of the non-linear parameters
@@
 is a 33 matrix containing the derivatives of the linear parameters with respect to the non-linear ones . It is obtained by computing the derivative of Eq. 30 with respect to
-Using Eq. 31 and Eq. 32, we deduce the derivatives
+Using Eq. 31 and Eq. 32, we deduce the derivatives as
 Therefore, the system (Eq. 28) becomes
-with being the new hessian matrix of the system. We have then split the 66 system (Eq. 28) into two smaller systems, each one of 33, in Eq. 30 and Eq. 39 respectively.
+with being the new hessian matrix of the system. We have then split the 66 system (Eq. 28) into two smaller systems, each one of 33, in Eq. 30 and Eq. 39, respectively.
 C.3 Analytical expressions of the involved matrices
@@
 As discussed in Sects.
-2 and 6.3.1, it is mainly because there is either no compelling evidence for variability in the RV time-series, there are not enough valid RV measurements, or the source lies outside the or ranges.
+2 and 6.3.1, it is mainly because there is either no compelling evidence for variability in the RV time series, there are not enough valid RV measurements, or the source lies outside the or ranges.
 We note that 20% of the entries in the whole SB9 contain an OBA component.
@@
 The same holds for the purely astrometric orbital solutions or those for EBs only based on photometry.
-Appendix H Parameters delivered to the catalogue for classes SB1, SB1C and TrendSB1.
+Appendix H Parameters delivered to the catalogue for classes SB1, SB1C and TrendSB1
 Table 15 lists the parameters computed in the present work concerning the SB1 class.
@@
 | Stage of application | Selection criterion | Description in |
-| Transfer from CU6 products | RV time-series available | Katz et al. (2023) |
+| Transfer from CU6 products | RV time series available | Katz et al. (2023) |
 | Normal stars, no spectral peculiarities | Katz et al. (2023) | |
@@
 | Selection of | rv_renormalized_gof larger than 4 | Sect. 2 |
-| objects/time-series at entry | in [3875 - 8125] K | Sect. 2 |
+| objects/time series at entry | in [3875 - 8125] K | Sect. 2 |
 | mag in [5.5 - 12] | Sect. 2 | |
 | rv_chisq_pvalue less than or equal to 0.01 | Sect. 2 | |
-| RV outlier rejected from time-series | Sect. 6.3.1 | |
+| RV outlier rejected from time series | Sect. 6.3.1 | |
 | larger than or equal to 10 | Sect. 6.3.1 | |